### v2

# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

Abstract
Deep neural networks (DNNs) have become ubiquitous in machine learning, but their energy consumption remains a notable issue. Lowering the supply voltage is an effective strategy for reducing energy consumption. However, aggressively scaling down the supply voltage can lead to accuracy degradation due to random bit flips in static random access memory (SRAM) where model parameters are stored. To address this challenge, we introduce NeuralFuse, a novel add-on module that addresses the accuracy-energy tradeoff in low-voltage regimes by learning input transformations to generate error-resistant data representations. NeuralFuse protects DNN accuracy in both nominal and low-voltage scenarios. Moreover, NeuralFuse is easy to implement and can be readily applied to DNNs with limited access, such as non-configurable hardware or remote access to cloud-based APIs. Experimental results demonstrate that, at a 1% bit error rate, NeuralFuse can reduce SRAM memory access energy by up to 24% while recovering accuracy by up to 57%. To the best of our knowledge, this is the first model-agnostic approach (i.e., no model retraining) to address low-voltage-induced bit errors. The source code is available at https://github.com/IBM/NeuralFuse.
1 Introduction
Energy-efficient computing is a primary consideration for the deployment of Deep Neural Networks (DNNs), particularly on edge devices and on-chip AI systems. Increasing energy efficiency and lowering the carbon footprint of DNN computation involves iterative efforts from both chip designers and algorithm developers. Processors with specialized hardware accelerators for AI computing are now ubiquitous, capable of providing orders of magnitude more performance and energy efficiency for AI computation. In addition to reduced precision/quantization and architectural optimizations, low voltage operation is a powerful knob that impacts power consumption. There is ample evidence in computer engineering literature that studies the effects of undervolting and low-voltage operation on accelerator memories that store weights and activations during computation. Aggressively scaling down the SRAM (Static Random Access Memory) supply voltage below the rated value leads to an exponential increase in bit failures, but saves power on account of the quadratic dependence of dynamic power on voltage (Chandramoorthy et al., 2019; Ganapathy et al., 2017). Such memory bit flips in the stored weight and activation values can cause catastrophic accuracy loss.
A recent spate of works advocates low voltage operation of DNN accelerators using numerous techniques to preserve accuracy ranging from hardware-based error mitigation techniques (Chandramoorthy et al., 2019; Reagen et al., 2016) to error-aware robust training of DNN models (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021). On-chip error mitigation methods have significant performance and power overheads. On the other hand, some have also proposed to generate models that are more robust to bit errors via a specific learning algorithm (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021), thereby eliminating the need for on-chip error mitigation. However, error-aware training to find the optimal set of robust parameters for each model is time and energy-intensive and may not be possible in all access-limited settings.
In this paper, we propose a novel model-agnostic approach: NeuralFuse. NeuralFuse is a machine learning module that allows for mitigating bit errors caused by very low voltage operation, through a trainable input transformation parameterized by a relatively small DNN, to enhance the robustness of the original input and provide accurate inference. The pipeline of NeuralFuse is illustrated in Figure 1. NeuralFuse accepts the scenarios under access-limited neural networks (e.g., non-configurable hardware or remote access to cloud-based APIs) to protect the deployed models from making wrong predictions under low power. Specifically, we consider two practical access-limited scenarios: (a) Relaxed Access, where the model details are unknown but backpropagation through the black-box models is possible, and (b) Restricted Access, where models are unknown, and backpropagation is disallowed. For relaxed access, we train NeuralFuse via backpropagation, and for restricted access, we train NeuralFuse on a white-box surrogate model and transfer it to the restricted access models. To the best of our knowledge, this is the first study that leverages a learning-based method to address random bit errors for recovering accuracy in low-voltage and access-limited settings.
We summarize our main contributions as follows:
We propose NeuralFuse, a novel learning-based input transformation framework to enhance the accuracy of DNNs subject to random bit errors caused by undervolting. NeuralFuse is model-agnostic because it operates in a plug-and-play manner at the data input and does not require re-training the deployed DNN model.
We consider two practical access-limited scenarios for neural network inference: Relaxed Access and Restricted Access. In the former setting, we use gradient-based methods to train the NeuralFuse module. In the latter setting, we use a white-box surrogate model to train NeuralFuse and show its high transferability to other types of DNN architectures.
We conduct extensive experiments on various combinations of DNN models (ResNet18, ResNet50, VGG11, VGG16, and VGG19), datasets (CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10), and NeuralFuse implementations with different architectures and sizes. The results show that NeuralFuse can consistently increase the perturbed accuracy (accuracy evaluated under random bit errors in weights) by up to , while simultaneously saving the overall SRAM memory access energy by up to , based on the realistic characterization of bit cell failures for a given memory array in a low-voltage regime inducing a of bit error rate.
We demonstrate NeuralFuse’s advantages in various scenarios. The experimental results show that NeuralFuse has high transferability (adaptability to unseen base models), versatility (capable of recovering low-precision quantization loss, as well as bit errors), and competitiveness (achieving state-of-the-art performance).
2 Related Work and Background
Software-based Energy Saving Strategies.
Recent studies have proposed reducing energy consumption from a software perspective. For instance, quantization techniques have been proposed to reduce the precision of storing model weights and decrease total memory storage (Gong et al., 2014; Rastegari et al., 2016; Wu et al., 2016). On the other hand, Yang et al. (2017) proposed energy-aware pruning on each layer and finetunes the weights to maximize the final accuracy. Yang et al. have also proposed several ways to reduce the energy consumption of DNN. For example, they proposed ECC, a DNN compression framework, that compresses a DNN model to meet the given energy constraint (Yang et al., 2019a); they have also proposed to compress DNN models via joint pruning and quantization (Yang et al., 2020). Besides, it is also feasible to treat energy constraint as an optimization problem during DNN training to reduce energy consumption and maximize training accuracy (Yang et al., 2019b). However, these methods focus on changing either the model architectures or model weights to reduce energy consumption, which is orthogonal to our approach with NeuralFuse, which serves as an add-on module for any given model.
Hardware-based Energy Saving Strategies.
Existing works have studied improving energy efficiency by designing specific hardware. Several works have studied the undervolting of DNN accelerators and proposed methods to maintain accuracy in the presence of bit errors. For instance, Reagen et al. (2016) proposed an SRAM fault mitigation technique that rounds the faulty weights into zeros to avoid the degradation of the prediction accuracy. Srinivasan et al. (2016) proposed to store the sensitive MSBs (most significant bits) in robust SRAM cells to preserve accuracy. Chandramoorthy et al. (2019) proposed dynamic supply voltage boosting to a higher voltage to improve the resilience of the memory access operation. On the other hand, Stutz et al. (2021) considered a learning-based approach that tries to find models that are robust to bit errors. The paper discusses several techniques to improve the robustness, such as quantization methods, weight clipping, random bit error training, and adversarial bit error training. The authors observe that the combination of quantization, weight clipping, and adversarial bit error training achieves excellent performance in their experiments. However, they also admitted that the training process for that is sensitive to hyperparameter settings, and hence, it might come with a challenging training procedure.
We argue that the methods mentioned above are not easy to implement or not suitable for real-world scenarios in access-limited settings. For example, the weights of DNN models packed on embedded systems may not be configurable or updatable. Therefore, model retraining (e.g., Stutz et al. (2021)) is not a viable option in this scenario. Moreover, the training of DNNs is already a tedious and time-consuming task. Adding error-aware training during training may further increase the training complexity and introduce challenges in hyperparameter search, as identified in previous literature. Özdenizci & Legenstein (2022) also note that this error-aware training was also found ineffective for large DNNs with millions of bits. Our proposed NerualFuse spares the need for model retraining by attaching a trainable input transformation function parameterized by a relatively small DNN as an add-on module to any DNN model as is.
SRAM Bit Errors in DNNs. Low-voltage-induced memory bit cell failures can cause bit flips from 0 to 1 and vice versa. In practice, SRAM memory bit errors increase exponentially when the supply voltage is scaled below , which is the minimum voltage required to avoid bit errors. This phenomenon has been well studied in previous literature, such as the work by Chandramoorthy et al. (2019) and Ganapathy et al. (2017). As shown in Figure 2, for an SRAM array of size 51264 bits and 14nm technology node, the number of bit errors increases as the voltage scales down. The corresponding dynamic energy per read access of SRAM is shown on the right, measured at each voltage at a constant frequency. In this example, accessing the SRAM at 0.83 leads to a bit error rate, and at the same time, the dynamic energy per access is reduced by approximately . This can lead to inaccurate inferences in DNNs, particularly when bit flips occur at the MSBs. However, improving the robustness to bit errors allows us to lower the voltage and exploit the resulting energy savings.
It has been observed that bit cell failures for a given memory array are randomly distributed and independent of each other. That is, the spatial distribution of bit flips could be assumed random, as it is generally different between arrays, even within a chip and between different chips. In this paper, we follow the methodology in Chandramoorthy et al. (2019) and model bit errors in a memory array of a given size by generating a random distribution of bit errors with equal likelihood of 0-to-1 and 1-to-0 bit flip. Specifically, we assume that the model weights are quantized to 8-bit precision (i.e., from 32-bit floats to 8-bit integers), and randomly distributed bit errors are injected into the quantized 2’s complement representation of weights to generate perturbed models. Please refer to Section 4.1 for more implementation details.
3 NeuralFuse: Framework and Algorithms
3.1 Error-Resistant Input Transformation
As illustrated in Figure 1, to overcome the drawback of performance degradation in low-voltage regimes, we propose a novel trainable input transformation module parametrized by a relatively small DNN, NeuralFuse, to mitigate the accuracy-energy tradeoff for model inference. The rationale is to use a specially designed loss function and training scheme to derive the NeuralFuse and apply it to the input data such that the transformed inputs will become robust to low-voltage-induced bit errors.
Consider the input sampled from the data distribution and a model with random bit errors on its weights (which is called a perturbed model). When there are no bit errors (i.e., the normal-voltage settings), the perturbed model reduces to a nominal deterministic model denoted by . NeuralFuse aims to ensure the perturbed model can make correct inferences on the transformed inputs as well as retain consistent results of in regular (normal-voltage) settings. In order to adapt to different data characteristics, NeuralFuse () is designed to be input-aware, which can be formally defined as:
where is a “generator” (i.e., an input transformation function) that can generate a perturbation based on the input . The NeuralFuse transformed data will be passed to the deployed model ( or ) for final inference. Without loss of generality, we assume the transformed input lies within a scaled input range , where is the (flattened) dimension of .
3.2 Training Objective and Optimizer
To train the generator , which should ensure the correctness of both the perturbed model and the clean model , we parameterize by a neural network and design the following training objective function:
where is the set of trainable parameters for , is the ground-truth label of , denotes the likelihood of computed by a model given a transformed input , is the distribution of the perturbed models inherited from the clean model under a random bit error rate, and is a hyperparameter that balances the importance between the nominal and perturbed models.
The training objective function can be readily converted to a loss function (loss) that evaluates the cross-entropy between the ground-truth label and the prediction . That is, the total loss function becomes
To optimize the loss function entailing the evaluation of the loss term on randomly perturbed models, our training process is inspired by the EOT (Expectation Over Transformation) attacks (Athalye et al., 2018). EOT attacks aim to find a robust adversarial example against a variety of image transformations. Based on the idea, we propose a new optimizer for solving (3), which we call Expectation Over Perturbed Models (EOPM). EOPM-trained generators can generate error-resistant input transformations and mitigate the inherent bit errors. However, it is computationally impossible to enumerate all possible perturbed models with random bit errors, and the number of realizations for perturbed models is limited by the memory constraint of GPUs used for training. In practice, we only take perturbed models for each iteration to calculate the empirical average loss, i.e.,
where is the number of simulated perturbed models under random bit errors to calculate the loss. Therefore, the gradient used to update the generator can be calculated as follows:
In our implementation, we find that can already deliver stable performance, and there is little gain in using a larger value. The ablation study for different can be found in Appendix E.
3.3 Training Algorithm
Algorithm 1 in Appendix A summarizes the training steps of NeuralFuse. We split the training data into mini-batches for training the generator in each epoch. For each mini-batch, we first feed these data into to get the transformed inputs. Also, we simulate perturbed models using a random bit error rate, denoted by , from . Then, the transformed inputs are fed into these perturbed models and the clean model to calculate their losses and gradients. Finally, NeuralFuse parameters are updated based on the gradient obtained by EOPM.
4 Experiments
In this section, we present the experimental setup and results of NeuralFuse on different datasets and various architectures. In addition, we also provide the visualization results, detailed analysis, and ablation studies to better understand the properties of NeuralFuse.
4.1 Experiment Setups
Datasets.
We evaluate NeuralFuse on four different datasets: CIFAR-10 (Krizhevsky & Hinton, 2009), CIFAR-100 (Krizhevsky & Hinton, 2009), GTSRB (Stallkamp et al., 2012), and ImageNet-10 (Deng et al., 2009). CIFAR-10 consists of ten classes, with 50,000 training images and 10,000 for testing. Similarly, CIFAR-100 consists of 100 classes, with 500 training images and 100 testing images in each class. GTSRB (German Traffic Sign Recognition Benchmark) is a dataset that contains 43 classes with 39,209 training images and 12,630 testing images. Similar to CIFAR-10 and CIFAR-100, we resize GTSRB into 32323 in our experiment. For ImageNet-10, we chose the same ten categories as Huang et al. (2022), and there are 13,000 training images and 500 test images, which are cropped into 2242243. Due to the space limit, the CIFAR-100 results are given in Appendix F and G.
Base Models.
We select several common architectures for our base models: ResNet18, ResNet50 (He et al., 2016), VGG11, VGG16, and VGG19 (Simonyan & Zisserman, 2015). In order to meet the setting of deploying models on a chip, all of our based models are trained by quantization-aware training following Stutz et al. (2021).
NeuralFuse Generators.
The architecture of the NeuralFuse generator () is based on the Encoder-Decoder structure. We design and compare three types of generators, namely Convolution-based, Deconvolution-based, and UNet-based. We also consider large(L)/small(S) network sizes for each type. Both Convolution-based and Deconvolution-based variants follow a similar architecture for ease of comparison. Also, the generators were trained with quantization-aware training. More details are given in Appendix B.
Convolution-based (Conv). Conv uses Convolution with MaxPool layers for the encoder and Convolution with UpSample layers for the decoder. This architecture has shown its efficiency and effectiveness for generating input-aware backdoor triggers (Nguyen & Tran, 2020).
Deconvolution-based (DeConv). DeConv uses Convolution with MaxPool layers for the encoder and Deconvolution layers for the decoder. This modification is expected to enhance its performance and reduce energy consumption.
UNet-based (UNet). UNet uses Convolution with MaxPool layers for the encoder and Deconvolution layers for the decoder. UNet is known for its robust performance in image segmentation (Ronneberger et al., 2015), and thus, we incorporated it as one of our implementations.
Energy Consumption Calculation.
In this paper, we focus on the resilience to low-voltage-induced bit errors in model weights, and our reported energy consumption in Figure 1 is based on the product of the total number of SRAM memory accesses in a systolic-array-based CNN accelerator and the dynamic energy per read access at a given voltage. Research by Chen et al. (2016) also shows that the energy consumption of the SRAM part (buffer and array) accounts for a large part of the total system energy consumption. We assume that there is no bit error on NeuralFuse, as we consider it as an add-on data preprocessing module that could also be performed by the general-purpose core. Therefore, when implemented in the accelerator, NeuralFuse computation is performed at a higher error-free voltage with a dynamically scaled voltage supply to the memories. We report the reduction in overall weight memory energy consumption (i.e., NeuralFuse + Base Model under bit error rate) with respect to the unprotected base model in the regular voltage mode (i.e., bit error rate and without NeuralFuse).
To obtain the number of memory accesses, we used the SCALE-SIM simulator (Samajdar et al., 2020), and our chosen configuration simulates an output-stationary dataflow and a 3232 systolic array with 256KB of weight memory. We obtained the dynamic energy per read access of the SRAM at the minimum voltage () and at the voltage corresponding to a bit error rate () from Cadence ADE Spectre simulations, both at the same clock frequency.
Relaxed and Restricted Access Settings.
Our experiments consider two scenarios (relaxed/restricted access). For relaxed access, the information of the base model is not entirely transparent but it allows obtaining gradients from the black-box model through backpropagation. Therefore, this setting allows direct training of NeuralFuse with the base model using EOPM. On the other hand, for restricted access, only the inference function is allowed for the base model. Therefore, we train NeuralFuse by using a white-box surrogate base model and then transfer the generator to the access-restricted model.
Computing Resources.
Our experiments are performed using 8 Nvidia Tesla V100 GPUs and are implemented with PyTorch. NeuralFuse generally takes 150 epochs to converge, and its training time is similar for the same base model. On both CIFAR-10 and CIFAR-100 datasets, the average training times were 17 hours (ResNet18), 50 hours (ResNet50), 9 hours (VGG11), 13 hours (VGG16), and 15 hours (VGG19). For GTSRB, the average training times were 9 hours (ResNet18), 27 hours (ResNet50), 5 hours (VGG11), 7 hours (VGG16), and 8 hours (VGG19). For ImageNet-10, the average training times were 32 hours (ResNet18), 54 hours (ResNet50), 50 hours (VGG11), 90 hours (VGG16), and 102 hours (VGG19).
(a) CIFAR-10, Bit Error Rate
(b) GTSRB, Bit Error Rate
(c) ImageNet-10, Bit Error Rate
4.2 Performance Evaluation on Relaxed Access Setting
The experimental results of the Relaxed Access setting are shown in Figure 3. We train and test NeuralFuse with various base models (ResNet18, ResNet50, VGG11, VGG16, and VGG19). Two power settings have been considered: nominal voltage (no bit error) and low voltage (random bit errors), and the corresponding bit error rate (B.E.R.) due to low voltage is (CIFAR-10, GTSRB) and (ImageNet-10). The B.E.R. of ImageNet-10 is lower because the pre-trained models have more parameters than CIFAR-10 and GTSRB. For each experiment, we sample perturbed models (independent from training) for evaluation and report the mean and standard deviation of the test accuracy. In the following, clean accuracy (C.A.) means that the model is measured at nominal voltage, and perturbed accuracy (P.A.) means that it is measured at low voltage.
For CIFAR-10 and GTSRB, we observe that large generators like ConvL and UNetL can significantly recover the perturbed accuracy in the range of to on ResNet18, VGG11, VGG16, and VGG19. For ResNet50, the recover percentage is slightly worse than other base models, but it can attain up to recover percentage on GTSRB. On the other hand, the recover percentage based on small generators like DeConvS are worse than larger generators. This can be explained by the better ability to learn error-resistant generators for larger-sized networks (though they may consume more energy). For ImageNet-10, using larger generators can also gain better performance recovery on perturbed accuracy. This also demonstrates that NeuralFuse can still work well even with large input sizes and is applicable to different datasets.
4.3 Transferability for Restricted Access Setting
In the Restricted Access scenario, we train NeuralFuse generators on a white-box surrogate base model and transfer it to other black-box base models. The experimental results are shown in Table 1. We adopt ResNet18 and VGG19 as the white-box surrogate (source) models for training the generators under a bit error rate. For the generators, we choose ConvL and UNetL as they obtain better performance in Figure 3.
From Table 1, we can find that transferring from a larger B.E.R. like can give strong resilience to a smaller B.E.R. like or . We also find that using VGG19 as a surrogate model with UNet-based generators like UNetL can give better recovery than other combinations. On the other hand, in some cases, we observe that if we transfer between the same source and target models (but with different B.E.R. for training and testing), the performance may outperform the original relaxed-access results. For instance, when transferring VGG19 with UNetL under a B.E.R. to VGG19 or VGG11 under a B.E.R., the accuracy is v.s. for VGG19 (original), and v.s. for VGG11 (original), respectively. We conjecture that the generators trained on a larger B.E.R. can actually cover the error patterns of a smaller B.E.R., and even help improve generalization under a smaller B.E.R. These findings show great promise for recovering the accuracy of access-limited base models in low-voltage settings.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1.5%) UNetL (1.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 1% 92.6 38.9 12.4 89.8 89.0 0.5 50.1 85.8 85.2 0.5 46.3 0.5% 70.1 11.6 89.6 0.2 19.5 85.7 0.2 15.6 ResNet50 1% 92.6 26.1 9.4 89.2 36.1 18 10.0 84.4 38.9 16 12.8 0.5% 61.0 10.3 74.1 10 13.1 72.7 4.6 11.7 VGG11 1% 88.4 42.2 11.6 86.3 59.2 10 17.0 82.3 69.8 7.5 27.6 0.5% 63.6 9.3 78.9 4.9 15.3 77.0 4.0 13.4 VGG16 1% 90.3 35.7 7.9 89.4 62.2 18 26.5 84.7 68.9 14 33.2 0.5% 66.6 8.1 83.4 5.5 16.8 80.5 5.9 13.9 VGG19 1% 90.5 36.0 12.0 89.8 49.9 23 13.9 85.0 55.1 17 19.1 0.5% 64.2 12.4 81.8 8.5 17.6 78.5 6.8 14.3 VGG19 ResNet18 1% 92.6 38.9 12.4 88.9 62.6 13 23.7 85.0 72.3 11 33.4 0.5% 70.1 11.6 84.2 7.2 14.1 82.1 2.2 12.0 ResNet50 1% 92.6 26.1 9.4 88.8 37.9 18 11.8 85.2 46.7 17 20.6 0.5% 61.0 10.3 76.6 7.8 15.6 78.3 3.7 17.3 VGG11 1% 88.4 42.2 11.6 88.9 76.0 6.1 33.8 85.5 81.9 3.9 39.7 0.5% 63.6 9.3 85.9 2.6 22.3 84.8 0.5 21.2 VGG16 1% 90.3 35.7 7.9 89.0 76.5 9.0 40.8 85.9 79.2 7.5 43.5 0.5% 66.6 8.1 87.7 0.7 21.1 84.7 0.9 18.1 VGG19 1% 90.5 36.0 12.0 89.1 80.2 12 44.2 86.3 84.3 1.2 48.3 0.5% 64.2 12.4 88.8 0.4 24.6 85.9 0.3 21.7 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (): clean accuracy, P.A. (): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Base Model ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 19.0 29.1 21.2 27.5 24.0 28.9 ResNet50 25.4 29.9 26.4 29.2 27.7 29.9 VGG11 6.6 27.5 11.2 24.1 17.1 27.2 VGG16 17.1 28.9 19.7 27.0 23.0 28.7 VGG19 20.3 29.7 22.3 27.8 24.8 29.1
4.4 Energy-Accuracy Tradeoff
We report the total dynamic energy consumption as the total number of SRAM accesses times the dynamic energy of a single SRAM access. Specifically, we used SCALE-SIM to calculate the total weight memory access (T.W.M.A.), which can be found in Table 6 in Appendix C. In Table 2, we report the percentage of energy savings at voltages that would yield a bit error rate for various base model and generator combinations. The formula of the energy saving percentage (ES, ) is defined as:
where NV denotes nominal voltage, and LV denotes low voltage regime.
The results demonstrate that when utilizing ResNet18 as a base model, NeuralFuse can recover model accuracy by while saving energy. In Appendix C, we provided more results on the tradeoff between energy and accuracy of different NeuralFuse and Base Model combinations. Overall, our results show promising results that using NeuralFuse can effectively restore model accuracy when the SRAM encounters low-voltage-induced random bit errors.
Runtime and Latency.
On the other hand, runtime and its corresponding energy consumption may also affect the overall energy saving. For instance, previous research has shown that Multiply-and-accumulate (MAC) operations account for over 99% of total operations in state-of-the-art CNNs, dominating both processing runtime and energy consumption (Yang et al., 2017). Therefore, we also report the results of MACs-based energy consumption estimation in Appendix C, and latency report in Appendix D. We want to note that the additional latency overhead is an inevitable tradeoff for reducing energy consumption in our setting. Although runtime and latency are not the major focus of this paper, we envision future research designing a more lightweight version of the NeuralFuse module or using model compression techniques on NeuralFuse to reduce latency and runtime.
4.5 Model Size and Efficiency of NeuralFuse
To provide a full performance characterization of NeuralFuse, we analyze the relationship between the final recovery of each base model and generators of varying parameter counts. The efficiency ratio is defined as the recover percentage in perturbed accuracy divided by the parameter count of NeuralFuse. We compare the efficiency ratio for all NeuralFuse generators (trained on CIFAR-10) in Table 3. The results show that UNet-based generators have better efficiency than both Convolution-based and Deconvolution-based ones per million parameters.
![[Uncaptioned image]](x19.png)
![[Uncaptioned image]](x20.png)
Figure 4: Reduced-precision accuracy
4.6 NeuralFuse on Reduced-precision Quantization
Here we explore the robustness of NeuralFuse to low-precision quantization on model weights. Uniform quantization is the de-facto method for quantizing model weights (Gholami et al., 2022). However, it is possible to cause an accuracy drop due to precision loss. As a bit-error-oriented protector, we seek to understand whether NeuralFuse could also make a recovery to mitigate the model’s accuracy drop in this scope. We conducted an experiment that uniformly quantized the model weights to a lower bit precision and reported the resulting accuracy. Specifically, we apply symmetric uniform quantization on our base models with different numbers of bits to induce precision loss. Therefore, the quantized weight (integer) is defined as , where denotes the original model weight (full precision), denotes the quantization scale parameter, and is the precision (number of bits) that used to quantize the models. In addition, we considered the possibility of random bit errors (the low-voltage setup) in low-precision regimes.
We use GTSRB pre-trained ResNet18 as an example to evaluate two NeuralFuse generators (ConvL and UNetL, trained with B.E.R.), and vary the precision from 8 bits to 2 bits (integer). The result is shown in Figure 4. We find that when bits, NeuralFuse can take effect to recover the accuracy in both scenarios. When , while NeuralFuse can still handle the bit-error-free model (top panel), it exhibits a limited ability to recover the random bit-error case (bottom panel). We find the result encouraging because the observed robustness to reduced-precision inference is an emergent ability of NeuralFuse. That is, NeuralFuse was only trained with random bit errors, but it demonstrated high accuracy to unseen bit quantization errors. This experiment showcases NeuralFuse’s potential in protecting against accuracy drops caused by different sources of bit errors. More experimental results of different base models and datasets can be found in Appendix H.
4.7 Extended Analysis
We highlight some key findings from the additional results in Appendix. In Appendix E, we compare NeuralFuse to the simple baseline of learning a universal input perturbation. We find that the baseline is much worse than NeuralFuse, which validates the necessity of adopting input-aware transformation for learning error-resistant data representations in low-voltage scenarios. In Appendix G, we find that ensemble training of white-box surrogate base models can further improve the transferability of NeuralFuse in the restricted-access setting. In Appendix K and Appendix L, we present visualization results of data embeddings and transformed inputs via NeuralFuse. In Appendix J, we show that NeuralFuse can further recover the accuracy of a base model trained with adversarial weight perturbation in the low-voltage setting.
5 Conclusion
In this paper, we propose NeuralFuse, the first non-intrusive post-hoc protection module for model inference against bit errors induced by low voltage. NeuralFuse is particularly suited for practical machine deployment settings where access to the base model is limited or relaxed. The design of NeuralFuse includes a novel loss function and a new optimizer named EOPM to handle simulated randomness in perturbed models. Our comprehensive experimental results and analysis show that NeuralFuse can significantly recover test accuracy (by up to ) while simultaneously enjoying up to reduction in memory access energy. Furthermore, NeuralFuse demonstrates high transferability (to access-constrained models), and versatility (e.g., robustness to low-precision quantization). Our results show that NeuralFuse provides significant improvements in mitigating the energy-accuracy tradeoff of neural network inference in low-voltage regimes and sheds new insights on green AI technology. Our future work includes extending our study to other neural network architectures and modalities, such as transformer-based language models.
Impact Statements
We see no immediate ethical or negative societal consequences of our work, and it holds the potential for positive social impacts. Previous works have highlighted scenarios affected by energy concerns, necessitating strategies to reduce energy consumption (Yang et al., 2017, 2019a). By enhancing the accuracy of models in low-voltage regimes, NeuralFuse can function as an effective protector and an efficient energy saver. This technology should be particularly valuable for extending the battery life of devices reliant on batteries.
Appendix
The appendix provides more implementation details for our method, experimental results on more datasets and settings, ablation studies, and qualitative analysis. The appendices cover the following:
Appendix A Training Algorithm of NeuralFuse
Input: Base model ; Generator ; Training data samples ; Distribution of the perturbed models ; Number of perturbed models ; Total training iterations
Output: Optimized parameters for the Generator
Appendix B Implementation Details of NeuralFuse Generator
We consider two main goals in designing the NeuralFuse Generator: 1) efficiency (so the overall energy overhead is decreased) and 2) robustness (so that it can generate robust patterns on the input image and overcome the random bit flipping in subsequent models). Accordingly, we choose to utilize an encode-decoder architecture in implementing the generator. The design of ConvL is inspired by Nguyen & Tran (2020), in which the authors utilize a similar architecture to design an input-aware trigger generator, and have demonstrated its efficiency and effectiveness. Furthermore, we attempted to enhance it by replacing the Upsampling layer with a Deconvolution layer, leading to the creation of DeConvL. The UNetL-based NeuralFuse draws inspiration from Ronneberger et al. (2015), known for its robust performance in image segmentation, and thus, we incorporated it as one of our architectures. Lastly, ConvS, DeConvS, and UNetS are scaled-down versions of the model designed to reduce computational costs and total parameters. The architectures of Convolutional-based and Deconvolutional-based are shown in Table 4, and the architecture of UNet-based generators is in Table 5. For the abbreviation used in the table, ConvBlock means the Convolution block, Conv means a single Convolution layer, DeConvBlock means the Deconvolution block, DeConv means a single Deconvolution layer, and BN means a Batch Normalization layer. We use , , and Adam optimizer. For CIFAR-10, GTSRB, and CIFAR-100, we set batch size for each base model. For ImageNet-10, we set for ResNet18, ResNet50 and VGG11, and for both VGG16 and VGG19.
ConvL ConvS DeConvL DeConvS Layers #CHs Layers #CHs Layers #CHs Layers #CHs (ConvBlock)2, MaxPool 32 ConvBlock, Maxpool 32 (ConvBlock)2, MaxPool 32 ConvBlock, Maxpool 32 (ConvBlock)2, MaxPool 64 ConvBlock, Maxpool 64 (ConvBlock)2, MaxPool 64 ConvBlock, Maxpool 64 (ConvBlock)2, MaxPool 128 ConvBlock, Maxpool 64 (ConvBlock)2, MaxPool, 128 ConvBlock, Maxpool 64 ConvBlock, UpSample, ConvBlock 128 ConvBlock, UpSample 64 ConvBlock 128 DeConvBlock 64 ConvBlock, UpSample, ConvBlock 64 ConvBlock, UpSample 32 DeConvBlock, ConvBlock 64 DeConvBlock 32 ConvBlock, UpSample, ConvBlock 32 ConvBlock, UpSample 3 DeConvBlock, ConvBlock 32 DeConv, BN, Tanh 3 Conv, BN, Tanh 32 Conv, BN, Tanh 3 Conv, BN, Tanh 3 • [Note] #CHs: number of channels.
UNetL UNetS Layers #Channels Layers #Channels L1: (ConvBlock)2 16 L1: (ConvBlock)2 8 L2: Maxpool, (ConvBlock)2 32 L2: Maxpool, (ConvBlock)2 16 L3: Maxpool, (ConvBlock)2 64 L3: Maxpool, (ConvBlock)2 32 L4: Maxpool, (ConvBlock)2 128 L4: Maxpool, (ConvBlock)2 64 L5: DeConv 64 L5: DeConv 32 L6: Concat[L3, L5] 128 L6: Concat[L3, L5] 64 L7: (ConvBlock)2 64 L7: (ConvBlock)2 32 L8: DeConv 32 L8: DeConv 16 L9: Concat[L2, L8] 64 L9: Concat[L2, L8] 32 L10: (ConvBlock)2 32 L10: (ConvBlock)2 16 L11: DeConv 16 L11: DeConv 8 L12: Concat[L1, L11] 32 L12: Concat[L1, L11] 16 L13: (ConvBlock)2 16 L13: (ConvBlock)2 8 L14: Conv 3 L14: Conv 3
Appendix C NeuralFuse’s Energy-Accuracy Tradeoff
SCALE-SIM.
SCALE-SIM (Samajdar et al., 2020) is a systolic array-based CNN simulator that can calculate the number of memory accesses and the total time in execution cycles by giving the specific model architecture and accelerator architectural configuration as inputs. In this paper, we use SCALE-SIM to calculate the weights memory access of 5 based models (ResNet18, ResNet50, VGG11, VGG16, VGG19), and 6 generators (ConvL, ConvS, DeConvL, DeConvS, UNetL, UNetS). While SCALE-SIM supports both Convolutional and Linear layers, it does not yet support Deconvolution layers. Instead, we try to approximate the memory costs of Deconvolution layers by Convolution layers. We change the input and output from Deconvolution into the output and input of the Convolution layers. Besides, we also change the stride into 1 when we approximate it. We also add padding for the convolution layers while generating input files for SCALE-SIM. In this paper, we only consider the energy saving on weights accesses, so we only take the value “SRAM Filter Reads” from the output of SCALE-SIM as the total weights memory accesses (T.W.M.A.) for further energy calculation.
In Table 6, we report the total weight memory access (T.W.M.A.) using SCALE-SIM. We then showed the energy-accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 5.
| Base Model | ResNet18 | ResNet50 | VGG11 | VGG16 | VGG19 | - |
| T.W.M.A. | 2,755,968 | 6,182,144 | 1,334,656 | 2,366,848 | 3,104,128 | - |
| NeuralFuse | ConvL | ConvS | DeConvL | DeConvS | UNetL | UNetS |
| T.W.M.A. | 320,256 | 41,508 | 259,264 | 86,208 | 180,894 | 45,711 |
[Note] T.W.M.A.: total weight memory access.
Parameters and MACs Calculation.
In addition to T.W.M.A., the number of parameters and MACs (multiply–accumulate operations) are also common metrics in measuring the energy consumption of machine learning models. Yang et al. (2017) have shown that the energy consumption of computation and memory accesses are both proportional to MACs, allowing us to take computation energy consumption into account.
Here, we use the open-source package ptflops (Sovrasov, 2018-2023) to calculate the parameters and MAC values of all the base models and the NeuralFuse generators, in the same units as Bejnordi et al. (2020) used. The results are shown in Table 7. Note that we modified the base model architectures for ImageNet-10, as it has larger input sizes. For example, we use a larger instead of in the first Convolution layer in ResNet-based models to enhance the learning abilities. Therefore, the parameters of base models are different between different datasets. For NeuralFuse generators, we utilize the same architectures for implementation (including ImageNet-10). As a result, our proposed NeuralFuse generators are generally smaller than base models, either on total model parameters or MAC values.
Base Model ResNet18 ResNet50 VGG11 VGG16 VGG19 - Parameter CIFAR-10 11,173,962 23,520,842 9,231,114 14,728,266 20,040,522 - ImageNet-10 11,181,642 23,528,522 128,812,810 134,309,962 139,622,218 MACs CIFAR-10 557.14M 1.31G 153.5M 314.43M 399.47M - ImageNet-10 1.82G 4.12G 7.64G 15.53G 19.69G NeuralFuse ConvL ConvS DeConvL DeConvS UNetL UNetS Parameter CIFAR-10 ImageNet-10 723,273 113,187 647,785 156,777 482,771 121,195 MACs CIFAR-10 80.5M 10.34M 64.69M 22.44M 41.41M 10.58M ImageNet-10 3.94G 506.78M 3.17G 1.1G 2.03G 518.47M
MACs-based Energy Consumption.
We can then use the MAC values to further approximate the end-to-end energy consumption of the whole model. Assume that all values are stored on SRAM and that a MAC represents single memory access. The corresponding MACs-based energy saving percentage (MAC-ES, %) can be derived from Eq. 7 (c.f. Sec. 4.4), and results can be found in Table 8. We can observe that most combinations can save a large amount of energy, except that VGG11 with two larger NeuralFuse (ConvL and DeConvL) may increase the total energy. These results are consistent with the results reported in Table 2. In addition, we also showed the MACs-based energy-accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 6.
Base Model ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 16.2 28.7 19.0 26.6 23.2 28.7 ResNet50 24.5 29.8 25.7 28.9 27.4 29.8 VGG11 -21.8 23.9 -11.5 16.0 3.6 23.7 VGG16 5.0 27.3 10.0 23.5 17.4 27.2 VGG19 10.4 28.0 14.4 25.0 20.2 28.0
Although using ConvL or DeConvL along with base model VGG11 for CIFAR-10 implies an increase in energy consumption, using other smaller-scale generators, we can still save the overall energy and recover the base model’s accuracy. That said, developers can always choose smaller generators (with orders of magnitude fewer MAC operations than the original network) to restore model accuracy, further demonstrating the flexibility of choosing NeuralFuse generators of different sizes.
Appendix D Inference Latency of NeuralFuse
In Table 9, we report the latency (batch_size=1, CIFAR-10/ImageNet-10 testing dataset) of utilizing the different NeuralFuse generators with two different base models, ResNet18 and VGG19. We can see that although NeuralFuse brings a certain degree of extra latency, we deemed it an inevitable tradeoff for reducing energy consumption in our setting. Although latency is not the main focus of this paper, we envision future research designing a more lightweight version of the NeuralFuse module or using model compression techniques on NeuralFuse to reduce latency.
ResNet18 (CIFAR-10) VGG19 (CIFAR-10) ResNet18 (ImageNet-10) VGG19 (ImageNet-10) Base Model 5.84 ms 5.32 ms 6.21 ms 14.34 ms + ConvL 9.37 ms (+3.53) 8.96 ms (+3.64) 10.51 ms (+4.3) 17.66 ms (+3.32) + ConvS 7.86 ms (+2.02) 7.40 ms (+2.08) 8.28 ms (+2.07) 16.72 ms (+2.38) + DeConvL 9.18 ms (+3.34) 8.59 ms (+3.27) 10.07 ms (+3.86) 17.24 ms (+2.90) + DeConvS 7.49 ms (+1.65) 7.04 ms (+1.72) 7.79 ms (+1.58) 15.67 ms (+1.33) + UNetL 10.69 ms (+4.85) 10.06 ms (+4.74) 11.14 ms (+4.93) 18.54 ms (+4.20) + UNetS 10.63 ms (+4.79) 10.13 ms (+4.81) 11.36 ms (+5.15) 18.60 ms (+4.26)
Appendix E Ablation Studies
Study for in EOPM.
Here, we study the effect of used in EOPM (Eq. 5). In Figure 7, we report the results for ConvL and ConvS on CIFAR-10 pre-trained ResNet18, under a 1% bit error rate (B.E.R.). The results demonstrate that if we apply larger , the performance increases until convergence. Specifically, for ConvL (Figure 6(a)), larger empirically has a smaller standard deviation; this means larger gives better stability but at the cost of time-consuming training. In contrast, for the small generator ConvS (Figure 6(b)), we can find that the standard deviation is still large even trained by larger ; the reason might be that small generators are not as capable of learning representations as larger ones. Therefore, there exists a tradeoff between the stability of the generator performance and the total training time. In our implementation, choosing or is a good balance.
Tradeoff Between Clean Accuracy (C.A.) and Perturbed Accuracy (P.A.).
We conducted an experiment to study the effect of different values, which balance the ratio of clean accuracy and perturbed accuracy. In Table 10, the experimental results showed that a smaller can preserve clean accuracy, but result in poor perturbed accuracy. On the contrary, larger can deliver higher perturbed accuracy, but with more clean accuracy drop. This phenomenon has also been observed in adversarial training (Zhang et al., 2019).
| Base Model | C.A. | P.A. | ConvL | |||
| C.A. (NF) | P.A. (NF) | R.P. | ||||
| ResNet18 | 10 | 92.6 | 38.9 12.4 | 90.1 | 88.0 1.7 | 49.1 |
| 5 | 89.8 | 87.8 1.7 | 48.8 | |||
| 1 | 90.0 | 84.2 3.8 | 45.3 | |||
| 0.1 | 91.6 | 65.7 9.3 | 26.8 | |||
| 0.01 | 92.2 | 43.6 13 | 4.7 | |||
| VGG19 | 10 | 90.5 | 36.0 12.0 | 89.6 | 77.9 19 | 41.9 |
| 5 | 89.8 | 77.7 19 | 41.7 | |||
| 1 | 89.9 | 73.1 19 | 37.1 | |||
| 0.1 | 89.1 | 51.2 16 | 15.2 | |||
| 0.01 | 90.2 | 36.8 12 | 0.8 | |||
[Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Comparison to Universal Input Perturbation (UIP).
Moosavi-Dezfooli et al. (2017) has shown that there exists a universal adversarial perturbation to the input data such that the model will make wrong predictions on a majority of the perturbed images. In our NeuralFuse framework, the universal perturbation is a special case when we set for any data sample . The transformed data sample then becomes , where and UIP is a trainable universal input perturbation that has the same size as the input data. The experimental results with the universal input perturbation are shown in Table 11. We observe that its performance is much worse than our proposed NeuralFuse. This result validates the necessity of adopting input-aware transformation for learning error-resistant data representations in low-voltage scenarios.
| Base Model | B.E.R. | C.A. | P.A. | C.A. (UIP) | P.A. (UIP) | R.P. |
| ResNet18 | 1% | 92.6 | 38.9 12.4 | 91.8 | 37.9 11 | -1.0 |
| 0.5% | 70.1 11.6 | 92.5 | 70.6 11 | 0.5 | ||
| ResNet50 | 1% | 92.6 | 26.1 9.4 | 80.7 | 21.0 5.9 | -5.1 |
| 0.5% | 61.0 10.3 | 91.9 | 62.4 12 | 1.4 | ||
| VGG11 | 1% | 88.4 | 42.2 11.6 | 86.9 | 43.0 11 | 0.8 |
| 0.5% | 63.6 9.3 | 88.2 | 64.2 8.8 | 0.6 | ||
| VGG16 | 1% | 90.3 | 35.7 7.9 | 90.1 | 37.1 8.5 | 1.4 |
| 0.5% | 66.6 8.1 | 90.4 | 67.3 8.1 | 0.7 | ||
| VGG19 | 1% | 90.5 | 36.0 12.0 | 89.9 | 35.3 12 | -0.7 |
| 0.5% | 64.2 12.4 | 90.1 | 64.4 12 | 0.2 |
[Note] B.E.R.: the bit error rate of the base model, C.A. (%): clean accuracy, UIP: universal input transformation parameter, P.A.(%): perturbed accuracy, and R.P.: total recover percentage of P.A. (UIP) v.s. P.A.
Appendix F Additional Experimental Results on Relaxed Access Settings
We conducted more experiments on Relaxed Access settings to show that our NeuralFuse can protect the models under different B.E.R. The results can be found in Sec. F.1 (CIFAR-10), Sec. F.2 (GTSRB), Sec. F.3 (ImageNet-10), and Sec. F.4 (CIFAR-100). For comparison, we also visualize the experimental results in the figures below each table.
F.1 CIFAR-10
Base Model NF C.A. 1% B.E.R. 0.5% B.E.R. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. ResNet18 ConvL 92.6 89.8 87.8 1.7 48.8 90.4 87.9 2.2 17.8 ConvS 88.2 59.5 11 20.6 91.7 78.4 8.3 8.3 DeConvL 38.9 89.6 88.5 0.8 49.6 70.1 90.2 90.0 0.2 19.9 DeConvS 12.4 82.9 68.8 6.4 29.9 11.6 84.1 79.9 3.6 9.8 UNetL 86.6 84.6 0.8 45.6 89.7 86.3 2.4 16.2 UNetS 84.4 68.8 6.0 29.8 90.9 80.7 5.8 10.7 ResNet50 ConvL 92.6 85.5 53.2 22 27.1 90.3 86.5 3.2 25.5 ConvS 85.2 34.6 14 8.5 90.8 73.3 8.7 12.3 DeConvL 26.1 87.4 63.3 21 37.2 61.0 89.5 87.2 2.5 26.2 DeConvS 9.4 82.4 42.2 17 16.1 10.3 90.3 75.5 8.1 14.5 UNetL 86.2 75.5 12 49.4 89.9 83.9 3.6 22.9 UNetS 77.3 56.2 19 30.1 89.7 76.1 7.2 15.1 VGG11 ConvL 88.4 89.6 87.2 2.9 45.1 89.8 87.0 1.3 23.3 ConvS 84.9 66.3 7.5 24.1 88.2 74.5 5.7 10.9 DeConvL 42.2 89.3 87.2 2.6 45.0 63.6 89.6 86.9 1.1 23.2 DeConvS 11.6 85.6 68.2 7.1 26.0 9.3 88.3 75.7 4.6 12.1 UNetL 87.1 83.6 1.3 41.4 88.0 82.4 1.8 18.8 UNetS 85.5 72.7 4.6 30.5 88.1 75.8 4.3 12.2 VGG16 ConvL 90.3 90.1 86.0 6.2 50.3 90.2 88.5 0.9 21.9 ConvS 87.4 59.6 12 23.9 89.9 77.8 4.8 11.1 DeConvL 35.7 89.7 85.5 6.8 49.8 66.6 89.7 88.2 1.0 21.4 DeConvS 7.9 86.8 66.5 11 30.8 8.1 90.0 78.4 4.7 11.8 UNetL 87.4 83.4 4.4 47.7 89.0 86.2 1.5 19.6 UNetS 87.4 71.2 8.2 35.5 89.0 80.2 3.5 13.7 VGG19 ConvL 90.5 89.8 77.7 19 41.7 90.4 88.1 1.8 23.9 ConvS 87.3 52.7 17 16.7 89.6 74.5 9.0 10.3 DeConvL 36.0 86.3 78.4 18 42.4 64.2 90.4 88.5 1.4 24.3 DeConvS 12.0 86.5 58.2 18 22.2 12.4 89.7 75.2 8.6 11.0 UNetL 86.3 82.1 4.8 46.0 89.1 85.0 2.7 20.8 UNetS 86.3 66.4 13 30.4 89.2 77.1 7.3 12.9 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
(a) CIFAR-10, Bit Error Rate
(b) CIFAR-10, Bit Error Rate
F.2 GTSRB
Base Model NF C.A. 1% B.E.R. 0.5% B.E.R. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. ResNet18 ConvL 95.5 95.7 91.1 4.7 54.2 93.4 89.5 1.9 14.3 ConvS 94.4 68.6 12 31.7 94.8 87.7 4.2 12.4 DeConvL 36.9 95.6 91.3 4.3 54.4 75.2 95.4 93.4 1.1 18.1 DeConvS 16.0 95.7 78.1 9.1 41.2 12.7 95.8 90.1 3.3 14.9 UNetL 96.2 93.8 1.0 56.9 96.2 93.5 1.6 18.3 UNetS 95.9 85.1 6.9 48.2 95.5 91.4 2.8 16.2 ResNet50 ConvL 95.0 95.6 71.6 20 42.1 94.6 90.6 3.7 16.6 ConvS 94.8 50.5 22 21.0 95.4 84.5 8.5 10.5 DeConvL 29.5 94.9 71.6 21 42.0 74.0 94.7 91.6 2.9 17.6 DeConvS 16.9 93.0 56.4 17 26.9 13.0 94.6 87.4 5.9 13.5 UNetL 94.5 80.6 15 51.1 96.5 93.7 2.3 19.7 UNetS 94.7 64.7 22 35.2 95.9 90.6 4.8 16.7 VGG11 ConvL 91.9 94.8 85.7 7.2 50.9 93.9 92.6 0.7 27.7 ConvS 91.1 62.2 11 27.3 90.9 80.5 3.5 15.7 DeConvL 34.9 95.0 84.6 7.6 49.7 64.9 93.6 91.9 0.6 27.1 DeConvS 12.4 92.4 67.5 11 32.6 10.8 92.3 83.1 3.7 18.2 UNetL 92.2 83.2 6.0 48.3 94.8 90.6 1.7 25.7 UNetS 94.7 73.4 10 38.5 94.6 88.9 2.2 24.1 VGG16 ConvL 95.2 96.3 72.4 12 57.3 95.6 93.2 1.8 34.4 ConvS 94.1 39.8 13 24.6 94.3 82.2 6.2 23.4 DeConvL 15.1 96.4 72.0 12 56.9 58.8 95.6 93.1 2.0 34.3 DeConvS 6.8 93.8 50.9 13 35.8 8.9 95.1 84.0 5.3 25.2 UNetL 95.8 78.6 11 63.5 96.0 92.8 2.0 34.0 UNetS 94.3 63.3 14 48.1 95.4 87.8 3.6 29.0 VGG19 ConvL 95.5 96.0 88.3 7.2 51.7 95.6 93.4 2.1 24.2 ConvS 93.8 69.0 14 32.4 94.9 87.0 4.4 17.8 DeConvL 36.6 95.4 87.2 7.5 50.6 69.1 95.5 92.4 2.2 23.3 DeConvS 6.8 94.5 73.1 12 36.5 11.1 95.5 88.8 3.7 19.7 UNetL 95.4 88.2 6.7 51.7 94.9 91.7 2.5 22.6 UNetS 94.6 80.6 9.0 44.1 96.5 90.8 3.4 21.6 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
(a) GTSRB, Bit Error Rate
(b) GTSRB, Bit Error Rate
F.3 ImageNet-10
| Base Model | NF | C.A. | 0.5% B.E.R. | |||
| P.A. | C.A. (NF) | P.A. (NF) | R.P. | |||
| ResNet18 | ConvL | 92.2 | 72.3 7.0 | 94.0 | 88.0 2.0 | 15.7 |
| ConvS | 91.8 | 83.6 4.1 | 11.3 | |||
| DeConvL | 94.0 | 89.2 1.3 | 16.9 | |||
| DeConvS | 92.8 | 87.5 2.3 | 15.2 | |||
| UNetL | 94.0 | 88.1 1.4 | 15.8 | |||
| UNetS | 93.2 | 86.4 2.2 | 14.1 | |||
| ResNet50 | ConvL | 89.8 | 39.4 11 | 92.2 | 80.0 5.8 | 40.6 |
| ConvS | 91.8 | 65.0 11 | 25.6 | |||
| DeConvL | 93.0 | 79.4 5.9 | 40.0 | |||
| DeConvS | 93.2 | 70.9 9.1 | 31.5 | |||
| UNetL | 92.2 | 80.5 5.8 | 41.1 | |||
| UNetS | 92.4 | 73.6 8.9 | 34.2 | |||
| VGG11 | ConvL | 91.6 | 47.8 13 | 92.0 | 86.1 3.7 | 38.3 |
| ConvS | 89.4 | 66.4 7.1 | 18.6 | |||
| DeConvL | 91.0 | 86.0 3.0 | 38.2 | |||
| DeConvS | 89.0 | 72.5 7.8 | 24.7 | |||
| UNetL | 92.4 | 83.0 3.5 | 35.2 | |||
| UNetS | 86.2 | 73.5 6.0 | 25.7 | |||
| VGG16 | ConvL | 94.6 | 38.4 17 | 90.8 | 77.1 11 | 38.7 |
| ConvS | 90.2 | 60.2 14 | 21.8 | |||
| DeConvL | 91.2 | 77.2 11 | 38.8 | |||
| DeConvS | 90.0 | 62.3 14 | 23.9 | |||
| UNetL | 90.6 | 81.1 5.9 | 42.7 | |||
| UNetS | 86.4 | 72.3 8.8 | 33.9 | |||
| VGG19 | ConvL | 92.4 | 37.2 11 | 91.4 | 75.5 8.8 | 38.3 |
| ConvS | 88.8 | 56.5 13 | 19.3 | |||
| DeConvL | 91.0 | 75.9 8.9 | 38.7 | |||
| DeConvS | 88.8 | 64.0 11 | 26.8 | |||
| UNetL | 89.4 | 77.9 6.1 | 40.7 | |||
| UNetS | 87.6 | 65.9 10 | 28.7 | |||
[Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
F.4 CIFAR-100
As mentioned in the previous section, larger generators like ConvL, DeConvL, and UNetL have better performance than small generators. For CIFAR-100, we find that the gains of utilizing NeuralFuse are less compared to the other datasets. We believe this is because CIFAR-100 is a more challenging dataset (more classes) for the generators to learn to protect the base models. Nevertheless, NeuralFuse can still function to restore some degraded accuracy; these results also demonstrate that our NeuralFuse is applicable to different datasets. In addition, although the recover percentage is less obvious on CIFAR-100 (the more difficult dataset), we can still conclude that our NeuralFuse is applicable to different datasets.
Base Model NF C.A. 1% B.E.R. 0.5% B.E.R. 0.35% B.E.R. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. ResNet18 ConvL 73.7 54.8 11.0 7.7 6.4 65.2 39.0 7.1 18.1 69.4 42.9 6.2 11.4 ConvS 49.7 4.2 2.2 -0.4 70.0 24.5 7.6 3.6 72.1 35.1 7.3 3.7 DeConvL 4.6 55.2 11.9 8.2 7.3 20.9 66.3 38.2 6.9 17.3 31.4 69.2 42.9 5.5 11.4 DeConvS 2.9 32.7 4.0 2.2 -0.6 7.4 68.2 25.9 6.8 5 7.6 71.6 35.8 5.5 4.4 UNetL 50.6 14.5 8.9 10.0 66.2 40.1 6.4 19.2 70.3 46.3 5.5 14.9 UNetS 26.8 4.6 2.5 -0.0 67.1 28.8 6.8 7.9 70.9 38.3 6.4 6.9 ResNet50 ConvL 73.5 63.5 3.2 1.7 0.1 68.4 28.8 6.7 7.6 72.0 40.8 7.5 5.1 ConvS 65.5 3.2 1.6 0.1 71.9 23.1 6.9 1.9 73.0 37.4 8.0 1.7 DeConvL 3.0 59.6 3.2 1.7 0.2 21.3 68.1 28.6 7.0 7.4 35.7 71.7 41.7 7.7 6.1 DeConvS 1.8 61.1 3.2 1.7 0.1 7.0 70.3 25.0 6.7 3.7 8.6 72.8 38.9 7.9 3.3 UNetL 39.0 5.0 1.7 1.9 66.6 36.5 6.2 15.3 70.8 45.3 6.7 9.6 UNetS 47.7 3.4 1.8 0.3 69.1 26.1 6.6 4.8 72.6 39.6 7.8 3.9 VGG11 ConvL 64.8 58.3 19.7 11 11.5 63.1 38.8 9.3 15.0 63.9 42.4 9.0 11.1 ConvS 56.6 10.4 7.4 2.2 62.7 27.9 10 4.0 63.9 41.8 8.3 10.5 DeConvL 8.2 60.3 21.2 11 13.0 23.9 63.9 40.0 9.0 16.2 31.3 64.0 42.8 9.1 11.5 DeConvS 5.7 58.3 11.8 7.9 3.5 9.4 61.9 29.8 9.9 5.9 10 63.5 36.1 10 4.8 UNetL 51.1 22.1 8.2 13.9 61.8 37.8 9.0 13.9 63.5 40.9 9.3 9.6 UNetS 51.9 13.1 7.9 4.9 61.7 29.8 9.7 6.0 63.8 35.7 9.9 4.5 VGG16 ConvL 67.8 51.4 19.2 6.0 12.6 61.8 41.1 5.6 18.7 64.9 44.9 5.3 13.8 ConvS 44.3 6.7 2.3 0.1 63.8 27.5 6.8 5.1 66.0 36.3 6.1 5.1 DeConvL 7.0 53.1 20.8 6.2 14.2 22.4 62.8 42.1 5.5 19.8 31.1 65.0 46.6 5.2 15.5 DeConvS 3.5 23.5 4.8 1.7 -1.8 7.0 62.1 29.9 6.7 7.5 7.2 64.9 38.1 6.3 7.0 UNetL 50.2 25.3 1.7 18.7 61.7 41.3 5.0 18.9 64.8 46.8 4.6 15.7 UNetS 27.7 9.9 2.1 3.3 61.6 31.3 6.3 8.9 65.0 39.8 5.9 8.7 VGG19 ConvL 67.8 59.4 29.2 8.1 18.6 65.6 46.5 6.8 12.5 66.9 49.2 7.4 7.0 ConvS 63.7 14.4 5.1 3.8 66.6 38.3 6.8 4.2 67.7 45.3 8.5 3.2 DeConvL 10.6 60.1 29.6 8.5 19.0 34.0 65.7 46.9 7.1 12.9 42.1 67.3 49.8 7.6 7.6 DeConvS 4.3 60.9 16.1 6.0 5.6 9.6 66.5 39.0 3.7 5.0 9.4 67.7 45.7 8.4 3.6 UNetL 58.7 30.2 8.2 19.6 65.5 46.9 6.5 12.9 67.4 50.0 7.5 7.9 UNetS 59.1 18.0 6.2 7.4 66.3 40.1 8.0 6.1 67.5 46.6 8.4 4.5 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
(a) CIFAR-100, Bit Error Rate
(b) CIFAR-100, Bit Error Rate
(c) CIFAR-100, Bit Error Rate
Appendix G Additional Experimental Results on Restricted Access Settings (Transferability)
We conduct more experiments with Restricted Access settings to show that our NeuralFuse can be transferred to protect various black-box models. The experimental results are shown in Sec. G.1 (CIFAR-10), Sec. G.2 (GTSRB), and Sec. G.3 (CIFAR-100).
We find that using VGG19 as a white-box surrogate model has better transferability than ResNet18 for all datasets. In addition, we can observe that some NeuralFuse generators have downward applicability if base models have a similar architecture. In other words, if we try to transfer a generator trained on a large B.E.R. (e.g., 1%) to a model with a small B.E.R. (e.g., 0.5%), the performance will be better than that of a generator trained with the original B.E.R. (e.g., 0.5%). For example, in Table 16, we could find that if we use VGG19 as a source model to train the generator ConvL (1%), the generator could deliver better performance (in terms of P.A. (NF)) when applied to similar base models (e.g., VGG11, VGG16, or VGG19) under a 0.5% B.E.R., compared to using itself as a source model (shown in Table 12). We conjecture that this is because the generators trained on a larger B.E.R. can also cover the error patterns of a smaller B.E.R., and thus they have better generalizability across smaller B.E.Rs.
To further improve the transferability to cross-architecture target models, we also conduct an experiment in Sec. G.4 to show that using ensemble-based training can help the generator to achieve this feature.
G.1 CIFAR-10
The results of CIFAR-10 in which NeuralFuse is trained at 1% B.E.R. are shown in Table 16.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1%) UNetL (1%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.5% 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 ResNet50 1% 92.6 26.1 9.4 89.5 36.0 19 9.9 85.2 38.8 19 12.7 0.5% 61.0 10.3 75.1 10 14.1 77.1 5.0 16.1 VGG11 1% 88.4 42.2 11.6 88.4 62.5 8.4 20.3 76.8 61.1 8.5 18.9 0.5% 63.6 9.3 81.0 4.6 17.4 73.7 3.0 10.1 VGG16 1% 90.3 35.7 7.9 89.6 63.3 18 27.6 85.2 59.9 16 24.2 0.5% 66.6 8.1 85.0 3.4 18.4 80.2 4.5 13.6 VGG19 1% 90.5 36.0 12.0 89.6 50.7 22 14.7 85.3 51.1 16 15.1 0.5% 64.2 12.4 80.2 8.7 16.0 76.5 7.8 12.3 VGG19 ResNet18 1% 92.6 38.9 12.4 89.8 61.0 17 22.1 87.0 69.7 11 30.8 0.5% 70.1 11.6 86.1 6.9 16.0 84.2 3.0 14.1 ResNet50 1% 92.6 26.1 9.4 89.9 34.0 19 7.9 87.0 44.2 17 18.1 0.5% 61.0 10.3 76.5 10 15.5 80.7 4.2 19.7 VGG11 1% 88.4 42.2 11.6 89.7 76.5 7.0 34.3 87.1 79.9 5.6 37.7 0.5% 63.6 9.3 88.0 2.1 24.4 85.4 0.8 21.8 VGG16 1% 90.3 35.7 7.9 89.6 75.5 12 39.8 87.2 78.9 7.8 43.2 0.5% 66.6 8.1 88.9 0.6 22.3 86.2 0.3 19.6 VGG19 0.5% 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
G.2 GTSRB
In Tables 17 and 18, we show the results on GTSRB in which NeuralFuse is trained at 1.5% and 1% B.E.R., respectively.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1.5%) UNetL (1.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 1% 95.5 36.9 16.0 95.7 93.9 1.9 57.0 94.9 94.4 0.4 57.5 0.5% 75.2 12.7 95.7 0.2 20.5 94.8 0.2 19.6 ResNet50 1% 95.0 29.5 16.9 94.4 37.0 22 7.5 94.4 47.1 23 17.6 0.5% 74.0 13.0 77.5 13 3.5 84.8 9.5 10.8 VGG11 1% 91.9 34.9 12.4 92.8 45.2 10 10.3 91.4 50.5 13 15.6 0.5% 64.9 10.8 79.4 5.8 14.5 83.9 4.2 19.0 VGG16 1% 95.2 15.1 6.8 95.4 31.1 13 15.9 94.6 36.8 12 21.7 0.5% 58.8 8.9 84.5 8.3 25.8 86.0 8.6 27.2 VGG19 1% 95.5 36.6 6.8 95.0 56.4 15 19.8 94.3 60.8 15 24.2 0.5% 69.1 11.1 86.9 3.4 17.8 87.7 3.8 18.6 VGG19 ResNet18 1% 95.5 36.9 16.0 88.4 50.3 12 13.4 92.8 63.7 16 26.8 0.5% 75.2 12.7 77.9 7.4 2.7 87.5 3.9 12.3 ResNet50 1% 95.0 29.5 16.9 87.5 29.7 17 0.2 92.5 40.4 21 10.9 0.5% 74.0 13.0 67.9 17 -6.1 77.5 15 3.5 VGG11 1% 91.9 34.9 12.4 89.7 47.1 11 12.2 93.5 60.0 12 25.1 0.5% 64.9 10.8 76.3 5.1 11.4 86.0 3.8 21.1 VGG16 1% 95.2 15.1 6.8 93.0 29.2 15 14.1 93.0 38.5 16 23.4 0.5% 58.8 8.9 75.7 12 16.9 79.9 8.3 21.1 VGG19 1% 95.5 36.6 6.8 95.1 87.4 6.0 50.8 94.6 88.7 5.0 52.1 0.5% 69.1 11.1 92.4 2.4 23.3 92.4 2.2 23.3 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1%) UNetL (1%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.5% 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 ResNet50 1% 95.0 29.5 16.9 94.5 35.6 21 6.1 95.6 42.6 23 13.1 0.5% 74.0 13.0 78.8 13 4.8 87.3 9.0 13.3 VGG11 1% 91.9 34.9 12.4 93.1 45.8 11 10.9 94.0 47.1 14 12.2 0.5% 64.9 10.8 81.8 5.0 16.9 84.2 4.8 19.3 VGG16 1% 95.2 15.1 6.8 95.5 26.5 12 11.4 95.5 32.4 11 17.3 0.5% 58.8 8.9 82.2 9.0 23.4 85.4 6.7 26.6 VGG19 1% 95.5 36.6 6.8 94.9 53.2 14 16.6 95.6 60.9 15 24.3 0.5% 69.1 11.1 85.4 4.5 16.3 87.5 3.7 18.4 VGG19 ResNet18 1% 95.5 36.9 16.0 93.7 53.1 16 16.2 95.0 63.4 18 26.5 0.5% 75.2 12.7 83.9 7.6 8.7 89.7 4.8 14.5 ResNet50 1% 95.0 29.5 16.9 92.8 30.6 18 1.1 95.4 38.9 22 9.4 0.5% 74.0 13.0 74.7 18 0.7 81.5 16 7.5 VGG11 1% 91.9 34.9 12.4 93.7 50.6 11 15.7 95.1 58.9 15 24.0 0.5% 64.9 10.8 82.3 5.1 17.4 87.5 3.7 22.6 VGG16 1% 95.2 15.1 6.8 95.2 27.8 15 12.7 95.2 33.5 14 18.4 0.5% 58.8 8.9 79.0 12 20.2 81.8 7.8 23.0 VGG19 0.5% 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
G.3 CIFAR-100
In Tables 19 and 20, we show results on CIFAR-100 with NeuralFuse trained at 1% and 0.5% B.E.R., respectively.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1%) UNetL (1%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.5% 73.7 20.9 7.4 54.8 35.8 5.2 14.9 50.6 39.3 2.8 18.4 0.35% 31.4 7.6 41.7 3.7 10.3 43.3 1.4 11.9 ResNet50 1% 73.5 3.0 1.8 44.9 2.2 2.0 -0.8 41.5 2.4 1.9 -0.6 0.5% 21.3 7.0 15.9 8.2 -5.4 17.1 7.1 -4.2 0.35% 35.7 8.6 23.7 7.1 -12.0 26.2 5.6 -9.5 VGG11 1% 64.8 8.2 5.7 41.2 9.8 5.6 1.6 37.5 10.2 5.1 2.0 0.5% 23.9 9.4 24.2 5.9 0.3 24.5 4.7 0.6 0.35% 31.3 10.0 29.0 5.4 -2.3 28.2 4.5 -3.1 VGG16 1% 67.8 7.0 3.5 44.0 7.9 3.7 0.9 39.5 10.1 4.5 3.1 0.5% 22.4 7.0 22.4 7.6 0.0 26.3 5.3 3.9 0.35% 31.1 7.2 28.1 5.9 -3.0 30.6 3.6 -0.5 VGG19 1% 67.8 10.6 4.3 44.2 13.5 6.1 2.9 40.7 15.6 6.2 5.0 0.5% 34.0 9.6 27.9 4.8 -6.1 29.3 4.6 -4.7 0.35% 42.1 9.4 33.2 48 -8.9 32.8 3.9 -9.3 VGG19 ResNet18 1% 73.7 4.6 2.9 55.5 5.8 3.7 1.2 57.3 6.8 4.4 2.2 0.5% 20.9 7.4 24.6 6.3 3.7 28.1 5.9 7.2 0.35% 31.4 7.6 31.1 5.0 -0.3 36.4 4.5 5.0 ResNet50 1% 73.5 3.0 1.8 56.1 2.8 2.1 -0.2 56.1 3.7 2.4 0.7 0.5% 21.3 7.0 18.9 8.6 -2.4 22.8 8.5 1.5 0.35% 35.7 8.6 28.7 8.2 -7.0 33.7 7.0 -2.0 VGG11 1% 64.8 8.2 5.7 52.8 12.3 8.4 4.1 53.9 15.4 9.4 7.2 0.5% 23.9 9.4 30.0 9.3 6.1 33.3 7.2 9.4 0.35% 31.3 10.0 36.5 7.7 5.2 38.8 6.5 7.5 VGG16 1% 67.8 7.0 3.5 53.6 11.2 4.4 4.2 55.2 13.6 5.2 6.6 0.5% 22.4 7.0 32.4 7.3 10.0 35.9 6.2 13.5 0.35% 31.1 7.2 39.4 6.3 8.3 42.4 4.9 11.3 VGG19 0.5% 67.8 34.0 9.6 59.4 50.2 3.1 16.2 58.7 49.1 3.5 15.1 0.35% 42.1 9.4 53.1 2.3 11.0 52.0 3.1 9.9 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
S.M. T.M. B.E.R. C.A. P.A. ConvL (0.5%) UNetL (0.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 0.35% 73.7 31.4 7.6 65.2 47.7 4.9 16.3 66.2 49.2 4.1 17.8 ResNet50 0.5% 73.5 21.3 7.0 62.5 24.0 9.9 2.8 63.5 26.4 9.1 5.1 0.35% 35.7 8.6 36.3 8.9 0.6 39.4 8.1 3.7 VGG11 0.5% 64.8 23.9 9.4 59.2 33.0 9.8 9.2 61.1 34.2 9.8 10.3 0.35% 31.3 10.0 40.4 8.7 9.1 41.4 9.0 10.1 VGG16 0.5% 67.8 22.4 7.0 59.5 34.7 8.0 12.3 61.4 37.5 6.8 15.2 0.35% 31.1 7.2 42.9 6.0 11.8 45.3 4.9 14.2 VGG19 0.5% 67.8 34.0 9.6 61.6 43.7 6.2 9.6 62.0 45.0 6.3 11.0 0.35% 42.1 9.4 49.0 5.5 6.8 50.5 5.3 8.3 VGG19 ResNet18 0.5% 73.7 20.9 7.4 66.1 24.9 6.7 4.0 67.8 27.7 6.8 6.8 0.35% 31.4 7.6 34.4 5.4 3.0 38.1 5.6 6.7 ResNet50 0.5% 73.5 21.3 7.0 66.2 22.7 7.8 1.4 66.7 25.4 8.0 4.2 0.35% 35.7 8.6 35.5 7.7 -0.2 38.8 7.5 3.2 VGG11 0.5% 64.8 23.9 9.4 59.9 29.3 10 5.4 61.0 31.2 9.8 7.4 0.35% 31.3 10.0 36.6 9.5 5.3 38.1 9.0 6.8 VGG16 0.5% 67.8 22.4 7.0 62.5 30.8 7.3 8.4 62.6 33.0 7.3 10.7 0.35% 31.1 7.2 40.0 6.5 8.9 42.5 5.9 11.3 VGG19 0.35% 67.8 42.1 9.4 65.6 52.0 6.2 9.8 65.5 52.6 6.1 10.4 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
G.4 Generator Ensembling
To improve the transferability performance on cross-architecture cases (e.g., using ResNet-based models as surrogate models to train NeuralFuse and then transfer NeuralFuse to VGG-based target models), we try to adopt ensemble surrogate models to train our NeuralFuse. The experimental results are shown in Table 21. We use the same experimental settings mentioned in Table 1 but change one source model (e.g., ResNet18 or VGG19) into two (ResNet18 with VGG19) for training. The results show that the overall performance is better than the results shown in Table 1, which means ensemble-based training can easily solve the performance degradation on cross-architecture target models.
S.M. T.M. B.E.R. C.A. P.A. ConvL (1.5%) UNetL (1.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 + VGG19 ResNet18 1% 92.6 38.9 12.4 89.4 88.1 1.0 49.2 86.3 85.4 0.5 46.5 0.5% 70.1 11.6 89.2 0.2 19.1 86.1 0.2 16.0 ResNet50 1% 92.6 26.1 9.4 89.3 44.0 22 17.9 86.1 50.9 20 24.8 0.5% 61.0 10.3 80.3 6.7 19.3 78.6 3.9 17.6 VGG11 1% 88.4 42.2 11.6 89.1 77.0 5.6 34.8 85.9 82.3 4.1 40.1 0.5% 63.6 9.3 87.5 1.6 23.9 85.0 0.6 21.4 VGG16 1% 90.3 35.7 7.9 89.1 80.5 8.6 44.8 85.7 81.4 5.5 45.7 0.5% 66.6 8.1 88.2 0.7 21.6 85.0 0.7 18.4 VGG19 1% 90.5 36.0 12.0 89.2 75.1 17 39.1 86.1 83.0 3.4 47.0 0.5% 64.2 12.4 89.0 0.2 24.8 85.9 0.4 21.7 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Appendix H NeuralFuse on Reduced-precision Quantization and Random Bit Errors
As mentioned in Sec. 4.6, we explore the robustness of NeuralFuse to low-precision quantization of model weights and consider the case of random bit errors. Here, we demonstrate that NeuralFuse can recover not only the accuracy drop due to reduced precision, but also the drop caused by low-voltage-induced bit errors (0.5% B.E.R.) under low precision. We selected two NeuralFuse generators (ConvL and UNetL) for our experiments, and these generators were trained with the corresponding base models (ResNet18 and VGG19) at 1% B.E.R. (CIFAR-10, GTSRB) and 0.5% B.E.R. (ImageNet-10). The experimental results are shown as follows: CIFAR-10 (Sec. H.1), GTSRB (Sec. H.2), and ImageNet-10 (Sec. H.3). Similarly, for ease of comparison, we visualize the experimental results in the figures below each table. Our results show that NeuralFuse can consistently perform well in low-precision regimes as well as recover the low-voltage-induced accuracy drop.
H.1 CIFAR-10
Base #Bits C.A. P.A. ConvL (1%) UNetL (1%) Model C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 8 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 7 92.5 68.8 10.4 89.8 89.5 1.7 20.7 86.5 86.0 0.5 17.2 6 92.6 68.4 11.2 89.7 89.5 0.2 21.1 86.6 85.9 0.3 17.5 5 92.4 52.7 14.1 89.7 90.0 0.7 37.3 86.5 85.5 0.8 32.8 4 91.8 26.3 12.7 89.8 58.7 24.5 32.4 86.6 64.9 22.5 38.6 3 84.8 11.3 1.8 89.8 12.8 5.8 1.5 86.0 14.8 10.0 3.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 7 90.3 66.5 8.5 89.8 89.6 0.2 23.1 87.4 86.7 0.3 20.2 6 90.1 59.8 13.2 89.9 89.4 3.8 29.6 87.4 86.4 0.7 26.6 5 90.2 37.7 14.1 89.8 78.0 15.8 40.3 87.2 79.8 0.8 42.1 4 87.5 14.7 6.0 89.8 27.8 18.9 13.1 87.2 34.4 20.5 19.7 3 78.3 10.5 1.5 89.7 10.9 2.6 0.4 86.8 11.0 2.9 0.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
H.2 GTSRB
Base #Bits C.A. P.A. ConvL (1%) UNetL (1%) Model C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 8 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 7 95.5 69.5 10.6 95.7 95.3 0.3 25.8 96.2 95.9 0.3 26.4 6 95.4 67.2 14.4 95.7 95.2 0.5 28.0 96.2 95.7 0.5 28.5 5 95.4 48.6 18.2 95.8 92.6 5.1 44.0 96.2 94.8 2.5 46.2 4 92.6 24.6 9.8 95.9 75.6 16.2 51.0 96.2 86.6 9.5 62.0 3 67.7 5.3 3.5 95.4 18.4 15.3 13.1 96.2 25.3 22.5 20.0 2 3.8 3.8 0.0 4.1 3.8 0.0 0.0 3.8 3.8 0.0 0.0 VGG19 8 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 7 95.6 66.1 14.8 96.0 92.2 5.7 26.1 95.4 92.6 3.7 26.5 6 95.3 64.2 8.4 96.0 92.2 5.7 28.0 95.4 92.3 2.3 28.1 5 95.2 48.2 14.0 96.0 92.2 5.7 44.0 95.4 86.2 8.4 38.0 4 92.0 18.2 14.3 93.0 92.2 5.7 74.0 95.0 49.6 22.8 31.4 3 60.0 2.0 0.9 87.3 92.2 5.7 90.2 87.2 1.7 0.9 -0.3 2 5.9 3.8 0.0 5.9 3.8 0.0 0.0 5.9 3.8 0.0 0.0 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
H.3 ImageNet-10
Base #Bits C.A. P.A. ConvL (0.5%) UNetL (0.5%) Model C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 8 92.2 72.3 7.0 94.0 88.0 2.0 15.7 94.0 88.1 1.4 15.8 7 92.4 70.6 13.0 94.2 86.7 4.1 16.1 93.6 87.8 3.5 17.2 6 92.4 68.9 9.9 94.2 85.1 4.8 16.2 93.6 86.4 3.7 17.5 5 91.0 60.9 13.0 94.2 82.5 6.8 21.6 94.0 83.2 5.9 22.3 4 91.4 47.4 9.8 93.8 68.6 9.8 21.2 92.6 68.7 9.2 21.3 3 85.2 28.8 11.8 89.2 44.1 14.0 15.3 89.4 42.7 14.2 13.9 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 92.4 37.2 11.0 91.4 75.5 8.8 38.3 89.4 77.9 6.1 40.7 7 92.0 27.3 6.6 91.2 59.3 13.0 32.0 89.4 65.4 10.0 38.1 6 92.4 27.9 6.4 91.0 59.7 11.8 31.8 89.4 64.9 9.9 37.0 5 92.0 15.1 4.4 91.6 23.1 0.7 8.0 89.0 27.9 8.8 12.8 4 89.4 12.2 2.7 90.8 14.0 4.3 1.8 89.6 14.6 4.9 2.4 3 46.8 9.9 0.5 83.2 10.4 0.6 0.5 84.2 9.9 0.7 0.0 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Appendix I Additional Experiments on Adversarial Training
Adversarial training is a common strategy to derive a robust neural network against certain perturbations. By training the generator using adversarial training proposed in Stutz et al. (2021), we report its performance against low voltage-induced bit errors. We use ConvL as the generator and ResNet18 as the base model, trained on CIFAR-10. Furthermore, we explore different flip bits as the perturbation on weights of the base model during adversarial training, and then for evaluation, the trained-generator will be applied against 1% of bit errors rate on the base model. The results are shown in Table 25. After careful tuning of hyperparameters, we find that we are not able to obtain satisfactory recovery when adopting adversarial training. Empirically, we argue that adversarial training may not be suitable for training generator-based methods.
K-bits C.A. P.A. C.A. (NF) P.A. (NF) R.P. 100 92.6 38.9 12.4 92.4 38.3 12.1 -0.6 500 92.1 38.7 12.5 -0.2 5,000 92.6 38.9 12.5 0 20,000 60.1 23.0 8.1 -16 100,000 71.1 23.6 6.6 -16 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Appendix J Additional Experiments on Robust Model Trained with Adversarial Weight Perturbation with NeuralFuse
Previously, Wu et al. (2020) proposed that one could obtain a more robust model via adversarial weight perturbation. To seek whether such models could also be robust to random bit errors, we conducted an experiment on CIFAR-10 with the proposed adversarially trained PreAct ResNet18. The experimental results are shown in Table 26. We find that the average perturbed accuracy is 23% and 63.2% for PreAct ResNet18 under 1% and 0.5% B.E.R., respectively. This result is lower than 38.9% and 70.1% from ResNet18 in Table 12, indicating their poor generalization ability against random bit errors. Nevertheless, when equipped NeuralFuse on the perturbed model, we could still witness a significant recover percentage under both 1% and 0.5% B.E.R. This result further demonstrates that NeuralFuse could be adapted to various models (i.e., trained in different learning algorithms).
| Base Model | B.E.R. | NF | C.A. | P.A. | C.A. (NF) | P.A. (NF) | R.P. |
| PreAct ResNet18 | 1% | ConvL | 89.7 | 23.0 9.3 | 87.6 | 53.7 26 | 30.7 |
| ConvS | 83.1 | 34.6 15 | 11.6 | ||||
| DeConvL | 87.7 | 55.4 27 | 32.4 | ||||
| DeConvS | 82.9 | 32.4 14 | 9.4 | ||||
| UNetL | 86.1 | 60.4 28 | 37.4 | ||||
| UNetS | 80.4 | 51.9 24 | 28.9 | ||||
| 0.5% | ConvL | 89.7 | 63.2 8.7 | 89.2 | 87.8 1.1 | 24.6 | |
| ConvS | 89.2 | 74.0 6.5 | 10.8 | ||||
| DeConvL | 89.0 | 87.4 1.1 | 24.2 | ||||
| DeConvS | 89.9 | 74.4 7.0 | 11.2 | ||||
| UNetL | 87.5 | 85.9 0.8 | 22.7 | ||||
| UNetS | 88.2 | 80.4 3.9 | 17.2 |
[Note] B.E.R.: the bit error rate of the base model, NF: NeuralFuse, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Appendix K Data Embeddings Visualization
To further understand how our proposed NeuralFuse works, we visualize the output distribution from the final linear layer of the base models and project the results onto the 2D space using t-SNE (van der Maaten & Hinton, 2008). Figure 15 shows the output distribution from ResNet18 (trained on CIFAR-10) under a 1% bit error rate. We chose two generators that have similar architecture: ConvL and ConvS, for this experiment. We can observe that: (a) The output distribution of the clean model without NeuralFuse can be grouped into 10 classes denoted by different colors. (b) The output distribution of the perturbed model under a 1% bit error rate without NeuralFuse shows mixed representations and therefore degraded accuracy. (c) The output distribution of the clean model with ConvL shows that applying NeuralFuse will not hurt the prediction of the clean model too much (i.e., it retains high accuracy in the regular voltage setting). (d) The output distribution of the perturbed model with ConvL shows high separability (and therefore high perturbed accuracy) as opposed to (b). (e)/(f) shows the output distribution of the clean/perturbed model with ConvS. For both (e) and (f), we can see nosier clustering when compared to (c) and (d), which means the degraded performance of ConvS compared to ConvL. The visualization validates that NeuralFuse can help retain good data representations under random bit errors and that larger generators in NeuralFuse have better performance than smaller ones.
Appendix L Qualitative Analysis of Transformed Inputs
In this section, we conduct a qualitative study to visualize the images transformed by NeuralFuse and present some properties and observations of these images. We utilize six different architectures of NeuralFuse generators trained with ResNet18 under a bit error rate.
Figure 16 (a) showcases several images from the truck class in CIFAR-10. Notably, images of the same class, when transformed by the same NeuralFuse, exhibit similar patterns, such as circles symbolizing the wheels of the trucks.
In Figures 16 (b) and 16 (c), we observe analogous phenomena in the GTSRB and CIFAR-100 datasets. Transformed images of the same class using the same generator consistently display patterns. On GTSRB, NeuralFuse-generated patterns highlight the sign’s shape with a green background, even if the original images have a dark background and are under different lighting conditions. These results further underscore the efficacy and efficiency of NeuralFuse.
Figure 17 presents more images from different classes in (a) CIFAR-10, (b) GTSRB, and (c) CIFAR-100. The transformed images exhibit distinct patterns for each class, suggesting that NeuralFuse effectively transforms images into class-specific patterns, making associated features robust to random bit errors and easily recognizable by the base model in low-voltage settings.
Clean
ConvL
ConvS
DeConvL
DeConvS
UNetL
UNetS
Clean |
ConvL |
ConvS |
DeConvL |
DeConvS |
UNetL |
UNetS |
Clean |
ConvL |
ConvS |
DeConvL |
DeConvS |
UNetL |
UNetS |

### v3

@@
 Abstract
-Deep neural networks (DNNs) have become ubiquitous in machine learning, but their energy consumption remains a notable issue.
-Lowering the supply voltage is an effective strategy for reducing energy consumption.
-However, aggressively scaling down the supply voltage can lead to accuracy degradation due to random bit flips in static random access memory (SRAM) where model parameters are stored.
-To address this challenge, we introduce NeuralFuse, a novel add-on module that addresses the accuracy-energy tradeoff in low-voltage regimes by learning input transformations to generate error-resistant data representations.
-NeuralFuse protects DNN accuracy in both nominal and low-voltage scenarios.
-Moreover, NeuralFuse is easy to implement and can be readily applied to DNNs with limited access, such as non-configurable hardware or remote access to cloud-based APIs.
-Experimental results demonstrate that, at a 1% bit error rate, NeuralFuse can reduce SRAM memory access energy by up to 24% while recovering accuracy by up to 57%.
-To the best of our knowledge, this is the first model-agnostic approach (i.e., no model retraining) to address low-voltage-induced bit errors.
-The source code is available at https://github.com/IBM/NeuralFuse.
+Deep neural networks (DNNs) have become ubiquitous in machine learning, but their energy consumption remains problematically high.
+An effective strategy for reducing such consumption is supply-voltage reduction, but if done too aggressively, it can lead to accuracy degradation.
+This is due to random bit-flips in static random access memory (SRAM), where model parameters are stored.
+To address this challenge, we have developed NeuralFuse, a novel add-on module that handles the energy-accuracy tradeoff in low-voltage regimes by learning input transformations and using them to generate error-resistant data representations, thereby protecting DNN accuracy in both nominal and low-voltage scenarios.
+As well as being easy to implement, NeuralFuse can be readily applied to DNNs with limited access, such cloud-based APIs that are accessed remotely or non-configurable hardware.
+Our experimental results demonstrate that, at a 1% bit-error rate, NeuralFuse can reduce SRAM access energy by up to 24% while recovering accuracy by up to 57%.
+To the best of our knowledge, this is the first approach to addressing low-voltage-induced bit errors that requires no model retraining.††footnotetext: Project Page: https://trustsafeai-neuralfuse.static.hf.space††footnotetext: Code: https://github.com/IBM/NeuralFuse
 1 Introduction
-Energy-efficient computing is a primary consideration for the deployment of Deep Neural Networks (DNNs), particularly on edge devices and on-chip AI systems.
-Increasing energy efficiency and lowering the carbon footprint of DNN computation involves iterative efforts from both chip designers and algorithm developers.
-Processors with specialized hardware accelerators for AI computing are now ubiquitous, capable of providing orders of magnitude more performance and energy efficiency for AI computation.
-In addition to reduced precision/quantization and architectural optimizations, low voltage operation is a powerful knob that impacts power consumption.
-There is ample evidence in computer engineering literature that studies the effects of undervolting and low-voltage operation on accelerator memories that store weights and activations during computation.
-Aggressively scaling down the SRAM (Static Random Access Memory) supply voltage below the rated value leads to an exponential increase in bit failures, but saves power on account of the quadratic dependence of dynamic power on voltage (Chandramoorthy et al., 2019; Ganapathy et al., 2017).
-Such memory bit flips in the stored weight and activation values can cause catastrophic accuracy loss.
-A recent spate of works advocates low voltage operation of DNN accelerators using numerous techniques to preserve accuracy ranging from hardware-based error mitigation techniques (Chandramoorthy et al., 2019; Reagen et al., 2016) to error-aware robust training of DNN models (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021).
-On-chip error mitigation methods have significant performance and power overheads.
-On the other hand, some have also proposed to generate models that are more robust to bit errors via a specific learning algorithm (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021), thereby eliminating the need for on-chip error mitigation.
-However, error-aware training to find the optimal set of robust parameters for each model is time and energy-intensive and may not be possible in all access-limited settings.
-In this paper, we propose a novel model-agnostic approach: NeuralFuse.
-NeuralFuse is a machine learning module that allows for mitigating bit errors caused by very low voltage operation, through a trainable input transformation parameterized by a relatively small DNN, to enhance the robustness of the original input and provide accurate inference.
+Energy-efficient computing is of primary importance to the effective deployment of deep neural networks (DNNs), particularly in edge devices and in on-chip AI systems.
+Increasing DNN computation’s energy efficiency and lowering its carbon footprint require iterative efforts from both chip designers and algorithm developers.
+Processors with specialized hardware accelerators for AI computing, capable of providing orders of magnitude better performance and energy efficiency for AI computation, are now ubiquitous.
+However, alongside reduced precision/quantization and architectural optimizations, endowing such systems with the capacity for low-voltage operation is a powerful lever for reducing their power consumption.
+The computer engineering literature contains ample evidence of the effects of undervolting and low-voltage operation on accelerator memories that store weights and activations during computation.
+Aggressive scaling-down of static random access memory’s (SRAM’s) supply voltage to below the rated value saves power, thanks to the quadratic dependence of dynamic power on voltage.
+Crucially, however, it also leads to an exponential increase in bit failures.
+Memory bit flips cause errors in the stored weight and activation values (Chandramoorthy et al., 2019; Ganapathy et al., 2017), leading to catastrophic accuracy loss.
+A recent wave of research has proposed numerous techniques for allowing low-voltage operation of DNN accelerators while preserving their accuracy.
+Most of these have been either hardware-based error-mitigation techniques or error-aware robust training of DNN models.
+On-chip error mitigation methods have significant performance and power overheads (Chandramoorthy et al., 2019; Reagen et al., 2016).
+On the other hand, some have proposed to generate models that are robust to bit errors via a specific learning algorithm (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021), thereby eliminating the need for on-chip error mitigation.
+However, error-aware robust training to find the optimal set of robust parameters for each model is time- and energy-intensive and may not be possible in all access-limited settings.
+In this paper, therefore, we propose a novel model-agnostic approach: NeuralFuse.
+This proof-of-concept machine-learning module offers trainable input transformation parameterized by a relatively small DNN; and, by enhancing input’s robustness, it mitigates bit errors caused by very low-voltage operation, thus serving the wider goal of more accurate inferencing.
 The pipeline of NeuralFuse is illustrated in Figure 1.
-NeuralFuse accepts the scenarios under access-limited neural networks (e.g., non-configurable hardware or remote access to cloud-based APIs) to protect the deployed models from making wrong predictions under low power.
-Specifically, we consider two practical access-limited scenarios: (a) Relaxed Access, where the model details are unknown but backpropagation through the black-box models is possible, and (b) Restricted Access, where models are unknown, and backpropagation is disallowed.
-For relaxed access, we train NeuralFuse via backpropagation, and for restricted access, we train NeuralFuse on a white-box surrogate model and transfer it to the restricted access models.
-To the best of our knowledge, this is the first study that leverages a learning-based method to address random bit errors for recovering accuracy in low-voltage and access-limited settings.
+To protect the deployed models from making wrong predictions under low-power conditions, NeuralFuse accepts scenarios under access-limited neural networks (e.g., non-configurable hardware or cloud-based APIs).
+Specifically, we consider two access-limited scenarios that are common in the real world: 1) relaxed access, in which ‘black box’ model details are unknown, but backpropagation through those models is possible; and 2) restricted access, in which the model details are unknown and backpropagation is disallowed.
+To enable it to deal with relaxed access, we trained NeuralFuse via backpropagation, and for restricted-access cases, we trained it on a white-box surrogate model.
+To the best of our knowledge, this is the first study that leverages a learning-based method to address random bit errors as a means of recovering accuracy in low-voltage and access-limited settings.
 We summarize our main contributions as follows:
-We propose NeuralFuse, a novel learning-based input transformation framework to enhance the accuracy of DNNs subject to random bit errors caused by undervolting.
-NeuralFuse is model-agnostic because it operates in a plug-and-play manner at the data input and does not require re-training the deployed DNN model.
-We consider two practical access-limited scenarios for neural network inference: Relaxed Access and Restricted Access.
-In the former setting, we use gradient-based methods to train the NeuralFuse module.
-In the latter setting, we use a white-box surrogate model to train NeuralFuse and show its high transferability to other types of DNN architectures.
-We conduct extensive experiments on various combinations of DNN models (ResNet18, ResNet50, VGG11, VGG16, and VGG19), datasets (CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10), and NeuralFuse implementations with different architectures and sizes.
-The results show that NeuralFuse can consistently increase the perturbed accuracy (accuracy evaluated under random bit errors in weights) by up to , while simultaneously saving the overall SRAM memory access energy by up to , based on the realistic characterization of bit cell failures for a given memory array in a low-voltage regime inducing a of bit error rate.
-We demonstrate NeuralFuse’s advantages in various scenarios.
-The experimental results show that NeuralFuse has high transferability (adaptability to unseen base models), versatility (capable of recovering low-precision quantization loss, as well as bit errors), and competitiveness (achieving state-of-the-art performance).
+We propose NeuralFuse, a novel learning-based input-transformation module aimed at enhancing the accuracy of DNNs that are subject to random bit errors caused by very low voltage operation.
+NeuralFuse is model-agnostic, i.e., operates on a plug-and-play basis at the data-input stage and does not require any re-training of deployed DNN models.
+We explore two practical limited-access scenarios for neural-network inference: relaxed access and restricted access.
+In the former setting, we use gradient-based methods for module training.
+In the latter one, we use a white-box surrogate model for training, which is highly transferable to other types of DNN architecture.
+We report the results of an extensive program of experiments with various combinations of DNN models (ResNet18, ResNet50, VGG11, VGG16, and VGG19), datasets (CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10), and NeuralFuse implementations of different architectures and sizes.
+These show that NeuralFuse can consistently increase the perturbed accuracy (accuracy evaluated under random bit errors in weights) by up to , while simultaneously saving up to of the energy normally required for SRAM access, based on our realistic characterization of bit-cell failures for a given memory array in a low-voltage regime inducing a bit-error rate.
+We demonstrate NeuralFuse’s transferability (i.e., adaptability to unseen base models), versatility (i.e., ability to recover low-precision quantization loss), and competitiveness (i.e., state-of-the-art performance) in various scenarios, establishing it as a promising proof-of-concept for energy-efficient, resilient DNN inference.
 2 Related Work and Background
-Software-based Energy Saving Strategies.
-Recent studies have proposed reducing energy consumption from a software perspective.
-For instance, quantization techniques have been proposed to reduce the precision of storing model weights and decrease total memory storage (Gong et al., 2014; Rastegari et al., 2016; Wu et al., 2016).
+Software-based Energy-saving Strategies.
+Various recent studies have proposed software-based methods of reducing computing’s energy consumption.
+For instance, quantization techniques have been reported to reduce the precision required for storing model weights, and thus to decrease total memory storage (Gong et al., 2014; Rastegari et al., 2016; Wu et al., 2016).
 On the other hand, Yang et al.
-(2017) proposed energy-aware pruning on each layer and finetunes the weights to maximize the final accuracy.
-Yang et al.
-have also proposed several ways to reduce the energy consumption of DNN.
-For example, they proposed ECC, a DNN compression framework, that compresses a DNN model to meet the given energy constraint (Yang et al., 2019a); they have also proposed to compress DNN models via joint pruning and quantization (Yang et al., 2020).
-Besides, it is also feasible to treat energy constraint as an optimization problem during DNN training to reduce energy consumption and maximize training accuracy (Yang et al., 2019b).
-However, these methods focus on changing either the model architectures or model weights to reduce energy consumption, which is orthogonal to our approach with NeuralFuse, which serves as an add-on module for any given model.
-Hardware-based Energy Saving Strategies.
-Existing works have studied improving energy efficiency by designing specific hardware.
-Several works have studied the undervolting of DNN accelerators and proposed methods to maintain accuracy in the presence of bit errors.
+(2017) - who proposed energy-aware pruning on each layer and fine-tuning of weights to maximize final accuracy - suggested several ways to reduce DNNs’ energy consumption.
+For example, they devised the ECC framework, which compresses DNN models to meet a given energy constraint (Yang et al., 2019a), and a method of compressing such models via joint pruning and quantization (Yang et al., 2020).
+It is also feasible, during DNN training, to treat energy constraints as an optimization problem and thereby reduce energy consumption while maximizing training accuracy (Yang et al., 2019b).
+However, unlike ours, all these methods imply changing either model architectures or model weights.
+Hardware-based Energy-saving Strategies.
+Prior studies have also explored ways of improving energy efficiency via specially designed hardware.
+Several of them have focused on the undervolting of DNN accelerators and proposed methods to maintain accuracy in the presence of bit errors.
 For instance, Reagen et al.
-(2016) proposed an SRAM fault mitigation technique that rounds the faulty weights into zeros to avoid the degradation of the prediction accuracy.
+(2016) proposed an SRAM fault-mitigation technique that rounds faulty weights to zero to avoid degradation of prediction accuracy.
 Srinivasan et al.
-(2016) proposed to store the sensitive MSBs (most significant bits) in robust SRAM cells to preserve accuracy.
+(2016) recommended storing sensitive MSBs (most significant bits) in robust SRAM cells to preserve accuracy.
 Chandramoorthy et al.
-(2019) proposed dynamic supply voltage boosting to a higher voltage to improve the resilience of the memory access operation.
-On the other hand, Stutz et al.
-(2021) considered a learning-based approach that tries to find models that are robust to bit errors.
-The paper discusses several techniques to improve the robustness, such as quantization methods, weight clipping, random bit error training, and adversarial bit error training.
-The authors observe that the combination of quantization, weight clipping, and adversarial bit error training achieves excellent performance in their experiments.
-However, they also admitted that the training process for that is sensitive to hyperparameter settings, and hence, it might come with a challenging training procedure.
-We argue that the methods mentioned above are not easy to implement or not suitable for real-world scenarios in access-limited settings.
-For example, the weights of DNN models packed on embedded systems may not be configurable or updatable.
-Therefore, model retraining (e.g., Stutz et al.
-(2021)) is not a viable option in this scenario.
-Moreover, the training of DNNs is already a tedious and time-consuming task.
-Adding error-aware training during training may further increase the training complexity and introduce challenges in hyperparameter search, as identified in previous literature.
-Özdenizci & Legenstein (2022) also note that this error-aware training was also found ineffective for large DNNs with millions of bits.
-Our proposed NerualFuse spares the need for model retraining by attaching a trainable input transformation function parameterized by a relatively small DNN as an add-on module to any DNN model as is.
+(2019) proposed dynamic supply-voltage boosting to improve the resilience of memory-access operations; and the learning-based approach proposed by Stutz et al.
+(2021) aims to find models that are robust to bit errors.
+The latter paper discusses several techniques for improving such robustness, notably quantization, weight-clipping, random bit-error training, and adversarial bit-error training.
+Its authors concluded from their experiments that a combination of quantization, weight-clipping, and adversarial bit-error training will yield excellent performance.
+However, they also admitted that the relevant training process was sensitive to hyperparameter settings, and hence, it might come with a challenging training procedure.
+However, we suggest that all the methods mentioned above are difficult to implement and/or unsuitable for use in real-world access-limited settings.
+For example, the weights of DNN models packed on embedded systems may not be configurable or updatable, making model retraining (e.g., Stutz et al.
+(2021)) non-viable in that scenario.
+Moreover, DNN training is already a tedious and time-consuming task, so adding error-aware training to it may further increase its complexity and, in particular, make hyperparameter searches more challenging.
+Özdenizci and Legenstein (2022) also reported that error-aware training was ineffective for large DNNs with millions of bits.
+NeuralFuse obviates the need for model retraining via an add-on trainable input-transformation function parameterized by a relatively small secondary DNN.
 SRAM Bit Errors in DNNs.
-Low-voltage-induced memory bit cell failures can cause bit flips from 0 to 1 and vice versa.
-In practice, SRAM memory bit errors increase exponentially when the supply voltage is scaled below , which is the minimum voltage required to avoid bit errors.
-This phenomenon has been well studied in previous literature, such as the work by Chandramoorthy et al.
+Low voltage-induced memory bit-cell failures can cause bit-flips from 0 to 1 and vice versa.
+In practice, SRAM bit errors increase exponentially when the supply voltage is scaled below , i.e., the minimum voltage required to avoid them.
+This phenomenon has been studied extensively in the prior literature, including work by Chandramoorthy et al.
 (2019) and Ganapathy et al.
 (2017).
-As shown in Figure 2, for an SRAM array of size 51264 bits and 14nm technology node, the number of bit errors increases as the voltage scales down.
-The corresponding dynamic energy per read access of SRAM is shown on the right, measured at each voltage at a constant frequency.
-In this example, accessing the SRAM at 0.83 leads to a bit error rate, and at the same time, the dynamic energy per access is reduced by approximately .
-This can lead to inaccurate inferences in DNNs, particularly when bit flips occur at the MSBs.
-However, improving the robustness to bit errors allows us to lower the voltage and exploit the resulting energy savings.
-It has been observed that bit cell failures for a given memory array are randomly distributed and independent of each other.
-That is, the spatial distribution of bit flips could be assumed random, as it is generally different between arrays, even within a chip and between different chips.
-In this paper, we follow the methodology in Chandramoorthy et al.
-(2019) and model bit errors in a memory array of a given size by generating a random distribution of bit errors with equal likelihood of 0-to-1 and 1-to-0 bit flip.
-Specifically, we assume that the model weights are quantized to 8-bit precision (i.e., from 32-bit floats to 8-bit integers), and randomly distributed bit errors are injected into the quantized 2’s complement representation of weights to generate perturbed models.
-Please refer to Section 4.1 for more implementation details.
+The specific increases in bit errors as voltage scales down, in the case of an SRAM array of bits with a 14nm technology node, is illustrated in Figure 2.
+The corresponding dynamic energy per SRAM read access, measured at each voltage at a constant frequency, is shown on the right-hand side of the figure.
+In this example, accessing the SRAM at 0.83 leads to a bit-error rate, and at the same time, dynamic energy per access is reduced by approximately .
+This can lead to DNNs making inaccurate inferences, particularly when bit-flips occur at the MSBs.
+However, improving robustness to bit errors can allow us to lower and exploit the resulting energy savings.
+It has been observed that bit-cell failures for a given memory array are randomly distributed and independent of each other.
+That is, the spatial distribution of bit-flips can be assumed to be random, as it generally differs from one array to another, within as well as between chips.
+Below, following Chandramoorthy et al.
+(2019), we model bit errors in a memory array of a given size by generating a random distribution of such errors with equal likelihood of 0-to-1 and 1-to-0 bit-flipping.
+More specifically, we assume that the model weights are quantized to 8-bit precision (i.e., from 32-bit floats to 8-bit integers), and generate perturbed models by injecting our randomly distributed bit errors into the two’s complement representation of weights.
+For more implementation details, please refer to Section 4.1.
 3 NeuralFuse: Framework and Algorithms
 3.1 Error-Resistant Input Transformation
-As illustrated in Figure 1, to overcome the drawback of performance degradation in low-voltage regimes, we propose a novel trainable input transformation module parametrized by a relatively small DNN, NeuralFuse, to mitigate the accuracy-energy tradeoff for model inference.
-The rationale is to use a specially designed loss function and training scheme to derive the NeuralFuse and apply it to the input data such that the transformed inputs will become robust to low-voltage-induced bit errors.
-Consider the input sampled from the data distribution and a model with random bit errors on its weights (which is called a perturbed model).
-When there are no bit errors (i.e., the normal-voltage settings), the perturbed model reduces to a nominal deterministic model denoted by .
-NeuralFuse aims to ensure the perturbed model can make correct inferences on the transformed inputs as well as retain consistent results of in regular (normal-voltage) settings.
-In order to adapt to different data characteristics, NeuralFuse () is designed to be input-aware, which can be formally defined as:
-where is a “generator” (i.e., an input transformation function) that can generate a perturbation based on the input .
-The NeuralFuse transformed data will be passed to the deployed model ( or ) for final inference.
+As illustrated in Figure 1, we propose a novel trainable input-transformation module, NeuralFuse, parametrized by a relatively small DNN, to mitigate the accuracy-energy tradeoff for model inference and thus overcome the drawback of performance degradation in low-voltage regimes.
+A specially designed loss function and training scheme are used to derive NeuralFuse and apply it to the input data such that the transformed inputs will become robust to low voltage-induced bit errors.
+Consider the input sampled from the data distribution and a set of models with random bit errors on weights (i.e., perturbed models).
+When it is not manifesting any bit errors (i.e., at normal-voltage settings), the perturbed model operates as a nominal deterministic one, denoted by .
+NeuralFuse aims to ensure that a model can make correct inferences on the transformed inputs while also delivering consistent results in its state.
+To adapt to various data characteristics, NeuralFuse – designated as in Eq.
+(1), below – is designed to be input-aware.
+This characteristic can be formally defined as
+where is a “generator” (i.e., an input-transformation function) that can generate a perturbation based on input .
+As transformed by NeuralFuse, i.e., as , that input is passed to the deployed model ( or ) for final inference.
 Without loss of generality, we assume the transformed input lies within a scaled input range , where is the (flattened) dimension of .
 3.2 Training Objective and Optimizer
-To train the generator , which should ensure the correctness of both the perturbed model and the clean model , we parameterize by a neural network and design the following training objective function:
-where is the set of trainable parameters for , is the ground-truth label of , denotes the likelihood of computed by a model given a transformed input , is the distribution of the perturbed models inherited from the clean model under a random bit error rate, and is a hyperparameter that balances the importance between the nominal and perturbed models.
-The training objective function can be readily converted to a loss function (loss) that evaluates the cross-entropy between the ground-truth label and the prediction .
-That is, the total loss function becomes
-To optimize the loss function entailing the evaluation of the loss term on randomly perturbed models, our training process is inspired by the EOT (Expectation Over Transformation) attacks (Athalye et al., 2018).
-EOT attacks aim to find a robust adversarial example against a variety of image transformations.
-Based on the idea, we propose a new optimizer for solving (3), which we call Expectation Over Perturbed Models (EOPM).
-EOPM-trained generators can generate error-resistant input transformations and mitigate the inherent bit errors.
-However, it is computationally impossible to enumerate all possible perturbed models with random bit errors, and the number of realizations for perturbed models is limited by the memory constraint of GPUs used for training.
-In practice, we only take perturbed models for each iteration to calculate the empirical average loss, i.e.,
-where is the number of simulated perturbed models under random bit errors to calculate the loss. Therefore, the gradient used to update the generator can be calculated as follows:
-In our implementation, we find that can already deliver stable performance, and there is little gain in using a larger value. The ablation study for different can be found in Appendix E.
+To train our generator , which ought to be able to ensure the correctness of both the perturbed model and the clean model , we parameterized it with a neural network and apply our training objective function
+where is the set of trainable parameters for ; is the ground-truth label of ; denotes the likelihood of as computed by a model being given a transformed input ; is the distribution of the perturbed models inherited from the clean model , under a random bit-error rate; and is a hyperparameter that balances the importance of the nominal and perturbed models.
+The training objective function can be readily converted to a loss function () that evaluates cross-entropy between the ground-truth label and the prediction .
+That is, the total loss function can be calculated as
+In particular, optimizing the loss function requires evaluation of the impact of the loss term on randomly perturbed models.
+Our training process is inspired by expectation over transformation (EOT) attacks (Athalye et al., 2018), which aim to produce robust adversarial examples that are simultaneously adversarial over the entire transformation distribution.
+Based on that idea, we propose a new optimizer for solving Eq.
+(3), which we call expectation over perturbed models (EOPM).
+EOPM-trained generators can generate error-resistant input transformations and mitigate inherent bit errors.
+However, it would be computationally impossible to enumerate all possible perturbed models with random bit errors, and the number of realizations for perturbed models is constrained by the memory size of the GPUs used for training.
+In practice, therefore, we only use perturbed models per iteration to calculate empirical average loss, i.e.,
+where is the number of perturbed models that are simulated to calculate the loss caused by random bit errors. Therefore, the gradient used to update the generator can be calculated as follows:
+Through our implementation, we found that stable performance could be delivered when , and that there was little to be gained by using a larger value.
+The results of our ablation study for different values of can be found in Appendix E.
 3.3 Training Algorithm
-Algorithm 1 in Appendix A summarizes the training steps of NeuralFuse.
-We split the training data into mini-batches for training the generator in each epoch.
-For each mini-batch, we first feed these data into to get the transformed inputs.
-Also, we simulate perturbed models using a random bit error rate, denoted by , from .
-Then, the transformed inputs are fed into these perturbed models and the clean model to calculate their losses and gradients.
+Algorithm 1 in Appendix A summarizes NeuralFuse’s training steps.
+Briefly, this involves splitting the training data into mini-batches for training the generator in each epoch.
+For each mini-batch, we first feed these data into to obtain the transformed inputs.
+Also, we simulate perturbed models using a random bit-error rate, denoted by , from .
+Then, the transformed inputs are fed into those perturbed models as well as into the clean model , and their respective losses and gradients are calculated.
 Finally, NeuralFuse parameters are updated based on the gradient obtained by EOPM.
 4 Experiments
-In this section, we present the experimental setup and results of NeuralFuse on different datasets and various architectures.
-In addition, we also provide the visualization results, detailed analysis, and ablation studies to better understand the properties of NeuralFuse.
 4.1 Experiment Setups
 Datasets.
-We evaluate NeuralFuse on four different datasets: CIFAR-10 (Krizhevsky & Hinton, 2009), CIFAR-100 (Krizhevsky & Hinton, 2009), GTSRB (Stallkamp et al., 2012), and ImageNet-10 (Deng et al., 2009).
-CIFAR-10 consists of ten classes, with 50,000 training images and 10,000 for testing.
-Similarly, CIFAR-100 consists of 100 classes, with 500 training images and 100 testing images in each class.
-GTSRB (German Traffic Sign Recognition Benchmark) is a dataset that contains 43 classes with 39,209 training images and 12,630 testing images.
+We evaluate NeuralFuse on four different datasets: CIFAR-10 (Krizhevsky and Hinton, 2009), CIFAR-100 (Krizhevsky and Hinton, 2009), the German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2012), and ImageNet-10 (Deng et al., 2009).
+CIFAR-10 consists of 10 classes, with 50,000 training images and 10,000 testing images in total.
+Similarly, CIFAR-100 consists of 100 classes, with 500 training images and 100 testing images in each.
+The GTSRB contains 43 classes with a total of 39,209 training images and 12,630 testing images.
 Similar to CIFAR-10 and CIFAR-100, we resize GTSRB into 32323 in our experiment.
 For ImageNet-10, we chose the same ten categories as Huang et al.
-(2022), and there are 13,000 training images and 500 test images, which are cropped into 2242243.
-Due to the space limit, the CIFAR-100 results are given in Appendix F and G.
+(2022), in which there are 13,000 training images and 500 test images cropped into 2242243.
+Due to space limitations, our CIFAR-100 results are presented in Appendices F and G.
 Base Models.
-We select several common architectures for our base models: ResNet18, ResNet50 (He et al., 2016), VGG11, VGG16, and VGG19 (Simonyan & Zisserman, 2015).
-In order to meet the setting of deploying models on a chip, all of our based models are trained by quantization-aware training following Stutz et al.
+We selected several common architectures for our base models: ResNet18, ResNet50 (He et al., 2016), VGG11, VGG16, and VGG19 (Simonyan and Zisserman, 2015).
+To replicate the deployment of models on chips, all our based models were given quantization-aware training that followed Stutz et al.
 (2021).
 NeuralFuse Generators.
-The architecture of the NeuralFuse generator () is based on the Encoder-Decoder structure.
-We design and compare three types of generators, namely Convolution-based, Deconvolution-based, and UNet-based.
-We also consider large(L)/small(S) network sizes for each type.
-Both Convolution-based and Deconvolution-based variants follow a similar architecture for ease of comparison.
-Also, the generators were trained with quantization-aware training.
-More details are given in Appendix B.
+The architecture of the NeuralFuse generator () ) is based on an encoder-decoder structure.
+We designed and compared three types of generators, namely convolution-based, deconvolution-based, and UNet-based.
+We also considered large(L)/small(S) network sizes for each type.
+Further details can be found below and in Appendix B.
 Convolution-based (Conv).
-Conv uses Convolution with MaxPool layers for the encoder and Convolution with UpSample layers for the decoder.
-This architecture has shown its efficiency and effectiveness for generating input-aware backdoor triggers (Nguyen & Tran, 2020).
+Conv uses convolution with MaxPool layers for its encoder and convolution with UpSample layers for its decoder.
+This architecture has previously been shown to be efficient and effective at generating input-aware backdoor triggers (Nguyen and Tran, 2020).
 Deconvolution-based (DeConv).
-DeConv uses Convolution with MaxPool layers for the encoder and Deconvolution layers for the decoder.
-This modification is expected to enhance its performance and reduce energy consumption.
+DeConv uses convolution with MaxPool layers for its encoder and deconvolution layers for its decoder.
+We expected this modification both to enhance its performance and to reduce its energy consumption.
 UNet-based (UNet).
-UNet uses Convolution with MaxPool layers for the encoder and Deconvolution layers for the decoder.
-UNet is known for its robust performance in image segmentation (Ronneberger et al., 2015), and thus, we incorporated it as one of our implementations.
-Energy Consumption Calculation.
-In this paper, we focus on the resilience to low-voltage-induced bit errors in model weights, and our reported energy consumption in Figure 1 is based on the product of the total number of SRAM memory accesses in a systolic-array-based CNN accelerator and the dynamic energy per read access at a given voltage.
+UNet uses convolution with MaxPool layers for its encoder, and deconvolution layers for its decoder.
+UNet is known for its robust performance in image segmentation (Ronneberger et al., 2015).
+Energy-consumption Calculation.
+The energy consumption reported in Figure 1 is based on the product of the total number of SRAM memory accesses in a systolic array-based convolution neural network (CNN) accelerator and the dynamic energy per read access at a given voltage.
 Research by Chen et al.
-(2016) also shows that the energy consumption of the SRAM part (buffer and array) accounts for a large part of the total system energy consumption.
-We assume that there is no bit error on NeuralFuse, as we consider it as an add-on data preprocessing module that could also be performed by the general-purpose core.
-Therefore, when implemented in the accelerator, NeuralFuse computation is performed at a higher error-free voltage with a dynamically scaled voltage supply to the memories.
-We report the reduction in overall weight memory energy consumption (i.e., NeuralFuse + Base Model under bit error rate) with respect to the unprotected base model in the regular voltage mode (i.e., bit error rate and without NeuralFuse).
-To obtain the number of memory accesses, we used the SCALE-SIM simulator (Samajdar et al., 2020), and our chosen configuration simulates an output-stationary dataflow and a 3232 systolic array with 256KB of weight memory.
-We obtained the dynamic energy per read access of the SRAM at the minimum voltage () and at the voltage corresponding to a bit error rate () from Cadence ADE Spectre simulations, both at the same clock frequency.
+(2016) previously showed that energy consumption by SRAM buffers and arrays accounts for a high proportion of total system energy consumption.
+We assume that there are no bit errors on NeuralFuse, given its status as an add-on data preprocessing module whose functions could also be performed by a general-purpose core.
+In this work we assume it is implemented on the accelerator equipped with dynamic voltage scaling and therefore NeuralFuse computation is performed at nominal error-free voltage.
+We report a reduction in overall weight-memory energy consumption (i.e., NeuralFuse + Base Model under a bit-error rate) with respect to the unprotected base model in the regular-voltage mode (i.e., bit-error rate and without NeuralFuse).
+To quantify memory accesses, we used the SCALE-SIM simulator (Samajdar et al., 2020), and our chosen configuration simulated an output-stationary dataflow and a 3232 systolic array with 256KB of weight memory.
+We collected data on the dynamic energy per read access of the SRAM both at and at the voltage corresponding to a bit-error rate () from Cadence ADE Spectre simulations, both at the same clock frequency.
 Relaxed and Restricted Access Settings.
-Our experiments consider two scenarios (relaxed/restricted access).
-For relaxed access, the information of the base model is not entirely transparent but it allows obtaining gradients from the black-box model through backpropagation.
-Therefore, this setting allows direct training of NeuralFuse with the base model using EOPM.
-On the other hand, for restricted access, only the inference function is allowed for the base model.
-Therefore, we train NeuralFuse by using a white-box surrogate base model and then transfer the generator to the access-restricted model.
+In the first of our experiments’ two scenarios, relaxed access, the base-model information was not entirely transparent, but allowed us to obtain gradients from the black-box model through backpropagation.
+Therefore, this scenario allowed direct training of NeuralFuse with the base model using EOPM.
+In the restricted-access scenario, on the other hand, only the inference function was allowed for the base model, and we therefore trained NeuralFuse using a white-box surrogate base model and then transfering the generator to the access-restricted model.
+(a) CIFAR-10, Bit-error Rate
+(b) GTSRB, Bit-error Rate
+(c) ImageNet-10, Bit-error Rate
 Computing Resources.
-Our experiments are performed using 8 Nvidia Tesla V100 GPUs and are implemented with PyTorch.
-NeuralFuse generally takes 150 epochs to converge, and its training time is similar for the same base model.
-On both CIFAR-10 and CIFAR-100 datasets, the average training times were 17 hours (ResNet18), 50 hours (ResNet50), 9 hours (VGG11), 13 hours (VGG16), and 15 hours (VGG19).
-For GTSRB, the average training times were 9 hours (ResNet18), 27 hours (ResNet50), 5 hours (VGG11), 7 hours (VGG16), and 8 hours (VGG19).
-For ImageNet-10, the average training times were 32 hours (ResNet18), 54 hours (ResNet50), 50 hours (VGG11), 90 hours (VGG16), and 102 hours (VGG19).
-(a) CIFAR-10, Bit Error Rate
-(b) GTSRB, Bit Error Rate
-(c) ImageNet-10, Bit Error Rate
-4.2 Performance Evaluation on Relaxed Access Setting
-The experimental results of the Relaxed Access setting are shown in Figure 3.
-We train and test NeuralFuse with various base models (ResNet18, ResNet50, VGG11, VGG16, and VGG19).
-Two power settings have been considered: nominal voltage (no bit error) and low voltage (random bit errors), and the corresponding bit error rate (B.E.R.) due to low voltage is (CIFAR-10, GTSRB) and (ImageNet-10).
-The B.E.R.
-of ImageNet-10 is lower because the pre-trained models have more parameters than CIFAR-10 and GTSRB.
-For each experiment, we sample perturbed models (independent from training) for evaluation and report the mean and standard deviation of the test accuracy.
-In the following, clean accuracy (C.A.) means that the model is measured at nominal voltage, and perturbed accuracy (P.A.) means that it is measured at low voltage.
-For CIFAR-10 and GTSRB, we observe that large generators like ConvL and UNetL can significantly recover the perturbed accuracy in the range of to on ResNet18, VGG11, VGG16, and VGG19.
-For ResNet50, the recover percentage is slightly worse than other base models, but it can attain up to recover percentage on GTSRB.
-On the other hand, the recover percentage based on small generators like DeConvS are worse than larger generators.
-This can be explained by the better ability to learn error-resistant generators for larger-sized networks (though they may consume more energy).
-For ImageNet-10, using larger generators can also gain better performance recovery on perturbed accuracy.
-This also demonstrates that NeuralFuse can still work well even with large input sizes and is applicable to different datasets.
-4.3 Transferability for Restricted Access Setting
-In the Restricted Access scenario, we train NeuralFuse generators on a white-box surrogate base model and transfer it to other black-box base models.
-The experimental results are shown in Table 1.
-We adopt ResNet18 and VGG19 as the white-box surrogate (source) models for training the generators under a bit error rate.
-For the generators, we choose ConvL and UNetL as they obtain better performance in Figure 3.
-From Table 1, we can find that transferring from a larger B.E.R.
-like can give strong resilience to a smaller B.E.R.
-like or .
-We also find that using VGG19 as a surrogate model with UNet-based generators like UNetL can give better recovery than other combinations.
-On the other hand, in some cases, we observe that if we transfer between the same source and target models (but with different B.E.R.
-for training and testing), the performance may outperform the original relaxed-access results.
-For instance, when transferring VGG19 with UNetL under a B.E.R.
-to VGG19 or VGG11 under a B.E.R., the accuracy is v.s.
-for VGG19 (original), and v.s.
-for VGG11 (original), respectively.
-We conjecture that the generators trained on a larger B.E.R.
-can actually cover the error patterns of a smaller B.E.R., and even help improve generalization under a smaller B.E.R.
-These findings show great promise for recovering the accuracy of access-limited base models in low-voltage settings.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (1.5%) UNetL (1.5%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ResNet18 1% 92.6 38.9 12.4 89.8 89.0 0.5 50.1 85.8 85.2 0.5 46.3 0.5% 70.1 11.6 89.6 0.2 19.5 85.7 0.2 15.6 ResNet50 1% 92.6 26.1 9.4 89.2 36.1 18 10.0 84.4 38.9 16 12.8 0.5% 61.0 10.3 74.1 10 13.1 72.7 4.6 11.7 VGG11 1% 88.4 42.2 11.6 86.3 59.2 10 17.0 82.3 69.8 7.5 27.6 0.5% 63.6 9.3 78.9 4.9 15.3 77.0 4.0 13.4 VGG16 1% 90.3 35.7 7.9 89.4 62.2 18 26.5 84.7 68.9 14 33.2 0.5% 66.6 8.1 83.4 5.5 16.8 80.5 5.9 13.9 VGG19 1% 90.5 36.0 12.0 89.8 49.9 23 13.9 85.0 55.1 17 19.1 0.5% 64.2 12.4 81.8 8.5 17.6 78.5 6.8 14.3 VGG19 ResNet18 1% 92.6 38.9 12.4 88.9 62.6 13 23.7 85.0 72.3 11 33.4 0.5% 70.1 11.6 84.2 7.2 14.1 82.1 2.2 12.0 ResNet50 1% 92.6 26.1 9.4 88.8 37.9 18 11.8 85.2 46.7 17 20.6 0.5% 61.0 10.3 76.6 7.8 15.6 78.3 3.7 17.3 VGG11 1% 88.4 42.2 11.6 88.9 76.0 6.1 33.8 85.5 81.9 3.9 39.7 0.5% 63.6 9.3 85.9 2.6 22.3 84.8 0.5 21.2 VGG16 1% 90.3 35.7 7.9 89.0 76.5 9.0 40.8 85.9 79.2 7.5 43.5 0.5% 66.6 8.1 87.7 0.7 21.1 84.7 0.9 18.1 VGG19 1% 90.5 36.0 12.0 89.1 80.2 12 44.2 86.3 84.3 1.2 48.3 0.5% 64.2 12.4 88.8 0.4 24.6 85.9 0.3 21.7 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(): clean accuracy, P.A.
-(): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
-Base Model ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 19.0 29.1 21.2 27.5 24.0 28.9 ResNet50 25.4 29.9 26.4 29.2 27.7 29.9 VGG11 6.6 27.5 11.2 24.1 17.1 27.2 VGG16 17.1 28.9 19.7 27.0 23.0 28.7 VGG19 20.3 29.7 22.3 27.8 24.8 29.1
-4.4 Energy-Accuracy Tradeoff
-We report the total dynamic energy consumption as the total number of SRAM accesses times the dynamic energy of a single SRAM access.
-Specifically, we used SCALE-SIM to calculate the total weight memory access (T.W.M.A.), which can be found in Table 6 in Appendix C.
-In Table 2, we report the percentage of energy savings at voltages that would yield a bit error rate for various base model and generator combinations.
-The formula of the energy saving percentage (ES, ) is defined as:
-where NV denotes nominal voltage, and LV denotes low voltage regime.
-The results demonstrate that when utilizing ResNet18 as a base model, NeuralFuse can recover model accuracy by while saving energy.
-In Appendix C, we provided more results on the tradeoff between energy and accuracy of different NeuralFuse and Base Model combinations.
-Overall, our results show promising results that using NeuralFuse can effectively restore model accuracy when the SRAM encounters low-voltage-induced random bit errors.
+Our experiments were performed using eight Nvidia Tesla V100 GPUs and implemented with PyTorch.
+NeuralFuse was found to generally take 150 epochs to converge, and its training time was similar to that of the base model it incorporated.
+On both the CIFAR-10 and CIFAR-100 datasets, average training times were 17 hours (ResNet18), 50 hours (ResNet50), 9 hours (VGG11), 13 hours (VGG16), and 15 hours (VGG19).
+For GTSRB, the average training times were 9 hours (ResNet18), 27 hours (ResNet50), 5 hours (VGG11), 7 hours (VGG16), and 8 hours (VGG19); and for ImageNet-10, the average training times were 32 hours (ResNet18), 54 hours (ResNet50), 50 hours (VGG11), 90 hours (VGG16), and 102 hours (VGG19).
+4.2 Performance Evaluation, Relaxed-access Scenario
+Our experimental results pertaining to the relaxed-access scenario are shown in Figure 3.
+The bit-error rate (BER) due to low voltage was in the cases of CIFAR-10 and GTSRB, and for ImageNet-10.
+The BER of ImageNet-10 was lower than that of the other two because, being pre-trained, it has more parameters than either of them.
+For each experiment, we sampled and evaluated perturbed models (independent from training), and below, we report the means and standard deviations of their respective accuracies.
+Below, clean accuracy (CA) refers to a model’s accuracy measured at nominal voltage, and perturbed accuracy (PA) to its accuracy measured at low voltage.
+In the cases of CIFAR-10 and the GTSRB, we observed that large generators like ConvL and UNetL recovered PA considerably, i.e., in the range of to on ResNet18, VGG11, VGG16, and VGG19.
+ResNet50’s recovery percentage was slightly worse than those of the other base models, but it nevertheless attained up to recovery on the GTSRB.
+On the other hand, the recovery percentages achieved when we used small generators like DeConvS were worse than those of their larger counterparts.
+This could be explained by larger-sized networks’ better ability to learn error-resistant generators (though perhaps at the cost of higher energy consumption).
+In the case of ImageNet-10, using larger generators also yielded better PA performance recovery, further demonstrating NeuralFuse’s ability to work well with large input sizes and varied datasets.
+4.3 Performance Evaluation, Restricted-access Scenario (Transferability)
+The experimental results of our restricted-access scenario are shown in Table 1.
+We adopted ResNet18 and VGG19 as our white-box surrogate source models for training the generators under a bit-error rate.
+We chose ConvL and UNetL as our generators because they performed best out of the six we tested (see Figure 3).
+SM TM BER CA PA ConvL (1.5%) UNetL (1.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 1% 92.6 38.9 12.4 89.8 89.0 0.5 50.1 85.8 85.2 0.5 46.3 0.5% 70.1 11.6 89.6 0.2 19.5 85.7 0.2 15.6 ResNet50 1% 92.6 26.1 9.4 89.2 36.1 18 10.0 84.4 38.9 16 12.8 0.5% 61.0 10.3 74.1 10 13.1 72.7 4.6 11.7 VGG11 1% 88.4 42.2 11.6 86.3 59.2 10 17.0 82.3 69.8 7.5 27.6 0.5% 63.6 9.3 78.9 4.9 15.3 77.0 4.0 13.4 VGG16 1% 90.3 35.7 7.9 89.4 62.2 18 26.5 84.7 68.9 14 33.2 0.5% 66.6 8.1 83.4 5.5 16.8 80.5 5.9 13.9 VGG19 1% 90.5 36.0 12.0 89.8 49.9 23 13.9 85.0 55.1 17 19.1 0.5% 64.2 12.4 81.8 8.5 17.6 78.5 6.8 14.3 VGG19 ResNet18 1% 92.6 38.9 12.4 88.9 62.6 13 23.7 85.0 72.3 11 33.4 0.5% 70.1 11.6 84.2 7.2 14.1 82.1 2.2 12.0 ResNet50 1% 92.6 26.1 9.4 88.8 37.9 18 11.8 85.2 46.7 17 20.6 0.5% 61.0 10.3 76.6 7.8 15.6 78.3 3.7 17.3 VGG11 1% 88.4 42.2 11.6 88.9 76.0 6.1 33.8 85.5 81.9 3.9 39.7 0.5% 63.6 9.3 85.9 2.6 22.3 84.8 0.5 21.2 VGG16 1% 90.3 35.7 7.9 89.0 76.5 9.0 40.8 85.9 79.2 7.5 43.5 0.5% 66.6 8.1 87.7 0.7 21.1 84.7 0.9 18.1 VGG19 1% 90.5 36.0 12.0 89.1 80.2 12 44.2 86.3 84.3 1.2 48.3 0.5% 64.2 12.4 88.8 0.4 24.6 85.9 0.3 21.7 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
+From Table 1, we can see that transferring from a larger BER such as can endow a smaller one (e.g., ) with strong resilience.
+The table also makes it clear that using VGG19 as a surrogate model with UNet-based generators like UNetL can yield better recovery performance than other model/generator combinations.
+On the other hand, we observed in some cases that if we transferred between source and target models of the same type (but with different BERs for training and testing), performance results could exceed those we had obtained during the original relaxed-access scenario.
+For instance, when we transferred VGG19 with UNetL under a BER to VGG19 or VGG11 under a BER, the resulting accuracies were (as against for original VGG19) and (as against for original VGG11).
+We conjecture that generators trained on relatively large BERs can cover the error patterns of smaller BERs, and even help improve the latter’s generalization.
+These findings indicate the considerable promise of access-limited base models in low-voltage settings to recover accuracy.
+ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 19.0 29.1 21.2 27.5 24.0 28.9 ResNet50 25.4 29.9 26.4 29.2 27.7 29.9 VGG11 6.6 27.5 11.2 24.1 17.1 27.2 VGG16 17.1 28.9 19.7 27.0 23.0 28.7 VGG19 20.3 29.7 22.3 27.8 24.8 29.1
+4.4 Energy/Accuracy Tradeoff
+We report total dynamic energy consumption as the total number of SRAM-access events multiplied by the dynamic energy of a single such event.
+Specifically, we used SCALE-SIM to calculate total weight-memory access (TWMA), specifics of which can be found in Appendix C’s Table 6.
+In Table 2, below, we report the percentages of energy saved (ES) at voltages that yield a bit-error rate for various base-model and generator combinations.
+The formula for computing ES is
+where NV denotes nominal voltage regime, and LV, a low-voltage one.
+Our results indicate that when ResNet18 is utilized as a base model, NeuralFuse can recover model accuracy by while reducing energy use by .
+In Appendix C, we provide more results on the tradeoff between energy and accuracy of different NeuralFuse and base-model combinations.
+Overall, it would appear that using NeuralFuse can effectively restore model accuracy when SRAM encounters low-voltage-induced random bit errors.
 Runtime and Latency.
-On the other hand, runtime and its corresponding energy consumption may also affect the overall energy saving.
-For instance, previous research has shown that Multiply-and-accumulate (MAC) operations account for over 99% of total operations in state-of-the-art CNNs, dominating both processing runtime and energy consumption (Yang et al., 2017).
-Therefore, we also report the results of MACs-based energy consumption estimation in Appendix C, and latency report in Appendix D.
-We want to note that the additional latency overhead is an inevitable tradeoff for reducing energy consumption in our setting.
-Although runtime and latency are not the major focus of this paper, we envision future research designing a more lightweight version of the NeuralFuse module or using model compression techniques on NeuralFuse to reduce latency and runtime.
-4.5 Model Size and Efficiency of NeuralFuse
-To provide a full performance characterization of NeuralFuse, we analyze the relationship between the final recovery of each base model and generators of varying parameter counts.
-The efficiency ratio is defined as the recover percentage in perturbed accuracy divided by the parameter count of NeuralFuse.
-We compare the efficiency ratio for all NeuralFuse generators (trained on CIFAR-10) in Table 3.
-The results show that UNet-based generators have better efficiency than both Convolution-based and Deconvolution-based ones per million parameters.
-![[Uncaptioned image]](x19.png)
-![[Uncaptioned image]](x20.png)
-Figure 4: Reduced-precision accuracy
-4.6 NeuralFuse on Reduced-precision Quantization
-Here we explore the robustness of NeuralFuse to low-precision quantization on model weights.
-Uniform quantization is the de-facto method for quantizing model weights (Gholami et al., 2022).
-However, it is possible to cause an accuracy drop due to precision loss.
-As a bit-error-oriented protector, we seek to understand whether NeuralFuse could also make a recovery to mitigate the model’s accuracy drop in this scope.
-We conducted an experiment that uniformly quantized the model weights to a lower bit precision and reported the resulting accuracy.
-Specifically, we apply symmetric uniform quantization on our base models with different numbers of bits to induce precision loss.
-Therefore, the quantized weight (integer) is defined as , where denotes the original model weight (full precision), denotes the quantization scale parameter, and is the precision (number of bits) that used to quantize the models.
-In addition, we considered the possibility of random bit errors (the low-voltage setup) in low-precision regimes.
-We use GTSRB pre-trained ResNet18 as an example to evaluate two NeuralFuse generators (ConvL and UNetL, trained with B.E.R.), and vary the precision from 8 bits to 2 bits (integer).
-The result is shown in Figure 4.
-We find that when bits, NeuralFuse can take effect to recover the accuracy in both scenarios.
-When , while NeuralFuse can still handle the bit-error-free model (top panel), it exhibits a limited ability to recover the random bit-error case (bottom panel).
-We find the result encouraging because the observed robustness to reduced-precision inference is an emergent ability of NeuralFuse.
-That is, NeuralFuse was only trained with random bit errors, but it demonstrated high accuracy to unseen bit quantization errors.
-This experiment showcases NeuralFuse’s potential in protecting against accuracy drops caused by different sources of bit errors.
-More experimental results of different base models and datasets can be found in Appendix H.
+On the other hand, runtime and its corresponding energy consumption may also affect overall energy savings.
+For instance, previous research has shown that multiply-and-accumulate (MAC) operations account for more than 99% of all operations in state-of-the-art CNNs, dominating processing runtime and energy consumption alike Yang et al.
+(2017).
+Therefore, we also report the results of our MAC-based energy-consumption estimation in Appendix C, and of our latency analysis in Appendix D.
+Here, it should also be noted that an additional latency overhead is an inevitable tradeoff for reducing energy consumption in our scenarios.
+Although neither runtime nor latency is a major focus of this paper, future researchers could usefully design a lighter-weight version of the NeuralFuse module, or apply model-compression techniques to it, to reduce these two factors.
+Base Model BER NeuralFuse ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 1% 67.5 182 76.6 190.7 94.5 245.9 0.5% 24.7 73.3 30.7 62.5 33.6 88.3 ResNet50 1% 37.4 75.1 57.4 102.7 102.3 248.4 0.5% 35.2 108.7 40.4 92.5 47.4 124.6 VGG11 1% 62.3 212.9 69.5 165.8 92.0 251.7 0.5% 32.3 96.3 35.8 77.2 38.9 100.7 VGG16 1% 69.6 211.2 76.9 196.5 98.8 292.9 0.5% 30.3 98.1 33.0 75.3 40.6 113 VGG19 1% 57.6 147.5 65.5 141.6 95.4 250.8 0.5% 33.0 91.0 37.5 70.2 43.1 106.4
+4.5 Model Size and NeuralFuse Efficiency
+To arrive at a full performance characterization of NeuralFuse, we analyzed the relationship between the final recovery achieved by each base model in combination with generators of varying parameter counts.
+For this purpose, we defined efficiency ratio as the recovery percentage in PA divided by NeuralFuse’s parameter count.
+Table 3 compares the efficiency ratios of all NeuralFuse generators trained on CIFAR-10.
+Those results show that UNet-based generators had better efficiency per million parameters than either convolution-based or deconvolution-based ones.
+4.6 NeuralFuse’s Robustness to Reduced-precision Quantization
+Lastly, we also explored NeuralFuse’s robustness to low-precision quantization on model weights.
+Uniform quantization is the usual method for quantizing model weights (Gholami et al., 2022).
+However, it is possible for it to cause an accuracy drop due to lack of precision.
+Given our aim of offering protection from bit errors, we hoped to understand whether NeuralFuse could also recover model-accuracy drops caused by this phenomenon.
+We therefore uniformly quantized the model weights to a lower bit precision and measured the resulting accuracy.
+Specifically, we applied symmetric uniform quantization to our base models with various numbers of bits to induce precision loss, and defined the quantized weight (integer) as , where denotes the original model weight (full precision), is the quantization scale parameter, and is the precision (number of bits) used to quantize the models.
+Bit errors induced by low voltage operation as previously described, are also applied to low precision weights.
+We used the GTSRB pre-trained ResNet18 as our example in an evaluation of two NeuralFuse generators, i.e., ConvL and UNetL trained with BER, and varied precision from 8 bits to 2 bits (integer).
+The results, shown in Figure 4, indicated that when bits, NeuralFuse could effectively recover accuracy in both the low-voltage and low-precision scenarios.
+When , while NeuralFuse could still handle the bit-error-free model (Fig.
+4 top), it exhibited a limited ability to recover the random bit-error case (Fig.
+4 bottom).
+We find these results encouraging, insofar as NeuralFuse was only trained on random bit errors, yet demonstrated high accuracy in dealing with unseen bit-quantization errors.
+Further experimental results derived from other base models and datasets can be found in Appendix H.
 4.7 Extended Analysis
-We highlight some key findings from the additional results in Appendix.
-In Appendix E, we compare NeuralFuse to the simple baseline of learning a universal input perturbation.
-We find that the baseline is much worse than NeuralFuse, which validates the necessity of adopting input-aware transformation for learning error-resistant data representations in low-voltage scenarios.
-In Appendix G, we find that ensemble training of white-box surrogate base models can further improve the transferability of NeuralFuse in the restricted-access setting.
-In Appendix K and Appendix L, we present visualization results of data embeddings and transformed inputs via NeuralFuse.
-In Appendix J, we show that NeuralFuse can further recover the accuracy of a base model trained with adversarial weight perturbation in the low-voltage setting.
+Here, we would like to highlight some key findings from the additional results in the Appendices.
+In Appendix E, we compare NeuralFuse against a simple baseline of learning a universal input perturbation.
+We found that such baseline performed much worse than NeuralFuse at that task, validating the necessity of adopting input-aware transformation if the goal is to learn error-resistant data representations in low-voltage scenarios.
+In Appendix G, we report that ensemble training of white-box surrogate base models could further improve the transferability of NeuralFuse in restricted-access scenarios.
+Appendices K and L present visualization results of NeuralFuse’s data embeddings and transformed inputs.
+In Appendix J, we show that NeuralFuse can further recover the accuracy of a base model trained with adversarial weight perturbation in a low-voltage setting.
 5 Conclusion
-In this paper, we propose NeuralFuse, the first non-intrusive post-hoc protection module for model inference against bit errors induced by low voltage.
-NeuralFuse is particularly suited for practical machine deployment settings where access to the base model is limited or relaxed.
-The design of NeuralFuse includes a novel loss function and a new optimizer named EOPM to handle simulated randomness in perturbed models.
-Our comprehensive experimental results and analysis show that NeuralFuse can significantly recover test accuracy (by up to ) while simultaneously enjoying up to reduction in memory access energy.
-Furthermore, NeuralFuse demonstrates high transferability (to access-constrained models), and versatility (e.g., robustness to low-precision quantization).
-Our results show that NeuralFuse provides significant improvements in mitigating the energy-accuracy tradeoff of neural network inference in low-voltage regimes and sheds new insights on green AI technology.
-Our future work includes extending our study to other neural network architectures and modalities, such as transformer-based language models.
-Impact Statements
-We see no immediate ethical or negative societal consequences of our work, and it holds the potential for positive social impacts.
-Previous works have highlighted scenarios affected by energy concerns, necessitating strategies to reduce energy consumption (Yang et al., 2017, 2019a).
-By enhancing the accuracy of models in low-voltage regimes, NeuralFuse can function as an effective protector and an efficient energy saver.
-This technology should be particularly valuable for extending the battery life of devices reliant on batteries.
+In this paper, we have proposed NeuralFuse, the first non-intrusive post hoc module that protects model inference against bit errors induced by low voltage.
+NeuralFuse is particularly well suited to practical machine-deployment cases in which access to the base model is limited or relaxed.
+The design of NeuralFuse includes a novel loss function and a new optimizer, EOPM, that enable it to handle simulated randomness in perturbed models.
+Our comprehensive experimental results and analysis show that NeuralFuse can recover test accuracy by up to while simultaneously enjoying an up to reduction in memory-access energy requirements.
+Moreover, NeuralFuse demonstrates high transferability to access-constrained models and high versatility, e.g., robustness to low-precision quantization.
+In short, NeuralFuse is a notable advancement in mitigating neural-network inference’s energy/accuracy tradeoff in low-voltage regimes, and points the way to greener future AI technology.
+Our future work will include extending this study to other neural-network architectures and modalities, such as transformer-based language models.
+Limitations.
+We acknowledge the challenge of achieving significant power savings without accuracy loss and view NeuralFuse as a foundational, proof-of-concept step toward this goal.
+Future research could enhance this approach by optimizing the pre-processing module to adapt to specific error characteristics of low-voltage SRAM or by integrating lightweight hardware modifications to further improve the energy-accuracy trade-off.
+Broader Impacts.
+We see no ethical or immediate negative societal consequence of our work, and it holds the potential for positive social impacts, from environmental benefits to improved access to technology and enhanced safety in critical applications.
+Acknowledgments and Disclosure of Funding
+We thank the anonymous reviewers for their insightful comments and valuable suggestions.
+The research described in this paper was conducted in the JC STEM Lab of Intelligent Design Automation, which is funded by The Hong Kong Jockey Club Charities Trust in support of Tsung-Yi Ho.
 Appendix
@@
 Accordingly, we choose to utilize an encode-decoder architecture in implementing the generator.
-The design of ConvL is inspired by Nguyen & Tran (2020), in which the authors utilize a similar architecture to design an input-aware trigger generator, and have demonstrated its efficiency and effectiveness.
+The design of ConvL is inspired by Nguyen and Tran (2020), in which the authors utilize a similar architecture to design an input-aware trigger generator, and have demonstrated its efficiency and effectiveness.
 Furthermore, we attempted to enhance it by replacing the Upsampling layer with a Deconvolution layer, leading to the creation of DeConvL.
@@
 UNetL UNetS Layers #Channels Layers #Channels L1: (ConvBlock)2 16 L1: (ConvBlock)2 8 L2: Maxpool, (ConvBlock)2 32 L2: Maxpool, (ConvBlock)2 16 L3: Maxpool, (ConvBlock)2 64 L3: Maxpool, (ConvBlock)2 32 L4: Maxpool, (ConvBlock)2 128 L4: Maxpool, (ConvBlock)2 64 L5: DeConv 64 L5: DeConv 32 L6: Concat[L3, L5] 128 L6: Concat[L3, L5] 64 L7: (ConvBlock)2 64 L7: (ConvBlock)2 32 L8: DeConv 32 L8: DeConv 16 L9: Concat[L2, L8] 64 L9: Concat[L2, L8] 32 L10: (ConvBlock)2 32 L10: (ConvBlock)2 16 L11: DeConv 16 L11: DeConv 8 L12: Concat[L1, L11] 32 L12: Concat[L1, L11] 16 L13: (ConvBlock)2 16 L13: (ConvBlock)2 8 L14: Conv 3 L14: Conv 3
-Appendix C NeuralFuse’s Energy-Accuracy Tradeoff
+Appendix C NeuralFuse’s Energy/Accuracy Tradeoff
 SCALE-SIM.
@@
 In Table 6, we report the total weight memory access (T.W.M.A.) using SCALE-SIM.
-We then showed the energy-accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 5.
+We then showed the energy/accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 5.
 | Base Model | ResNet18 | ResNet50 | VGG11 | VGG16 | VGG19 | - |
@@
 These results are consistent with the results reported in Table 2.
-In addition, we also showed the MACs-based energy-accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 6.
-Base Model ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 16.2 28.7 19.0 26.6 23.2 28.7 ResNet50 24.5 29.8 25.7 28.9 27.4 29.8 VGG11 -21.8 23.9 -11.5 16.0 3.6 23.7 VGG16 5.0 27.3 10.0 23.5 17.4 27.2 VGG19 10.4 28.0 14.4 25.0 20.2 28.0
+In addition, we also showed the MACs-based energy/accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 6.
+ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 16.2 28.7 19.0 26.6 23.2 28.7 ResNet50 24.5 29.8 25.7 28.9 27.4 29.8 VGG11 -21.8 23.9 -11.5 16.0 3.6 23.7 VGG16 5.0 27.3 10.0 23.5 17.4 27.2 VGG19 10.4 28.0 14.4 25.0 20.2 28.0
 Although using ConvL or DeConvL along with base model VGG11 for CIFAR-10 implies an increase in energy consumption, using other smaller-scale generators, we can still save the overall energy and recover the base model’s accuracy.
@@
 In Table 9, we report the latency (batch_size=1, CIFAR-10/ImageNet-10 testing dataset) of utilizing the different NeuralFuse generators with two different base models, ResNet18 and VGG19.
-We can see that although NeuralFuse brings a certain degree of extra latency, we deemed it an inevitable tradeoff for reducing energy consumption in our setting.
-Although latency is not the main focus of this paper, we envision future research designing a more lightweight version of the NeuralFuse module or using model compression techniques on NeuralFuse to reduce latency.
+While NeuralFuse contributes some additional latency, we consider this an unavoidable tradeoff necessary to achieve reduced energy consumption within our framework.
+Although the primary focus of this paper is not on latency, we acknowledge its importance.
+Future research could explore designing a more lightweight version of the NeuralFuse module or applying model compression techniques to minimize latency.
+Additionally, we recognize that running NeuralFuse on a general-purpose CPU could lead to different latency and energy consumption figures due to various influencing factors like CPU architecture and manufacturing processes.
 ResNet18 (CIFAR-10) VGG19 (CIFAR-10) ResNet18 (ImageNet-10) VGG19 (ImageNet-10) Base Model 5.84 ms 5.32 ms 6.21 ms 14.34 ms + ConvL 9.37 ms (+3.53) 8.96 ms (+3.64) 10.51 ms (+4.3) 17.66 ms (+3.32) + ConvS 7.86 ms (+2.02) 7.40 ms (+2.08) 8.28 ms (+2.07) 16.72 ms (+2.38) + DeConvL 9.18 ms (+3.34) 8.59 ms (+3.27) 10.07 ms (+3.86) 17.24 ms (+2.90) + DeConvS 7.49 ms (+1.65) 7.04 ms (+1.72) 7.79 ms (+1.58) 15.67 ms (+1.33) + UNetL 10.69 ms (+4.85) 10.06 ms (+4.74) 11.14 ms (+4.93) 18.54 ms (+4.20) + UNetS 10.63 ms (+4.79) 10.13 ms (+4.81) 11.36 ms (+5.15) 18.60 ms (+4.26)
@@
 5).
-In Figure 7, we report the results for ConvL and ConvS on CIFAR-10 pre-trained ResNet18, under a 1% bit error rate (B.E.R.).
+In Figure 7, we report the results for ConvL and ConvS on CIFAR-10 pre-trained ResNet18, under a 1% bit error rate (BER).
 The results demonstrate that if we apply larger , the performance increases until convergence.
@@
 In our implementation, choosing or is a good balance.
-Tradeoff Between Clean Accuracy (C.A.) and Perturbed Accuracy (P.A.).
+Tradeoff Between Clean Accuracy (CA) and Perturbed Accuracy (PA).
 We conducted an experiment to study the effect of different values, which balance the ratio of clean accuracy and perturbed accuracy.
@@
 This phenomenon has also been observed in adversarial training (Zhang et al., 2019).
-| Base Model | C.A. | P.A. | ConvL | |||
-| C.A. (NF) | P.A. (NF) | R.P. | ||||
+| Base Model | CA | PA | ConvL | |||
+| CA (NF) | PA (NF) | RP | ||||
 | ResNet18 | 10 | 92.6 | 38.9 12.4 | 90.1 | 88.0 1.7 | 49.1 |
@@
 | 0.01 | 90.2 | 36.8 12 | 0.8 | |||
-[Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
+Note. CA (%): clean accuracy; PA (%): perturbed accuracy; NF: NeuralFuse; and RP: total recover percentage of PA (NF) vs. PA
 Comparison to Universal Input Perturbation (UIP).
@@
 This result validates the necessity of adopting input-aware transformation for learning error-resistant data representations in low-voltage scenarios.
-| Base Model | B.E.R. | C.A. | P.A. | C.A. (UIP) | P.A. (UIP) | R.P. |
+| Base Model | BER | CA | PA | CA (UIP) | PA (UIP) | RP |
 | ResNet18 | 1% | 92.6 | 38.9 12.4 | 91.8 | 37.9 11 | -1.0 |
@@
 | 0.5% | 64.2 12.4 | 90.1 | 64.4 12 | 0.2 |
-[Note] B.E.R.: the bit error rate of the base model, C.A.
-(%): clean accuracy, UIP: universal input transformation parameter, P.A.(%): perturbed accuracy, and R.P.: total recover percentage of P.A.
-(UIP) v.s.
-P.A.
+Note. BER: the bit-error rate of the base model; CA (): clean accuracy; PA (): perturbed accuracy; UIP: universal input transformation parameter; and RP: total recovery percentage of PA (UIP) vs. PA
 Appendix F Additional Experimental Results on Relaxed Access Settings
-We conducted more experiments on Relaxed Access settings to show that our NeuralFuse can protect the models under different B.E.R.
-The results can be found in Sec.
+We conducted more experiments on Relaxed Access settings to show that our NeuralFuse can protect the models under different BER The results can be found in Sec.
 F.1 (CIFAR-10), Sec.
@@
 F.1 CIFAR-10
-Base Model NF C.A.
-1% B.E.R.
-0.5% B.E.R.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ConvL 92.6 89.8 87.8 1.7 48.8 90.4 87.9 2.2 17.8 ConvS 88.2 59.5 11 20.6 91.7 78.4 8.3 8.3 DeConvL 38.9 89.6 88.5 0.8 49.6 70.1 90.2 90.0 0.2 19.9 DeConvS 12.4 82.9 68.8 6.4 29.9 11.6 84.1 79.9 3.6 9.8 UNetL 86.6 84.6 0.8 45.6 89.7 86.3 2.4 16.2 UNetS 84.4 68.8 6.0 29.8 90.9 80.7 5.8 10.7 ResNet50 ConvL 92.6 85.5 53.2 22 27.1 90.3 86.5 3.2 25.5 ConvS 85.2 34.6 14 8.5 90.8 73.3 8.7 12.3 DeConvL 26.1 87.4 63.3 21 37.2 61.0 89.5 87.2 2.5 26.2 DeConvS 9.4 82.4 42.2 17 16.1 10.3 90.3 75.5 8.1 14.5 UNetL 86.2 75.5 12 49.4 89.9 83.9 3.6 22.9 UNetS 77.3 56.2 19 30.1 89.7 76.1 7.2 15.1 VGG11 ConvL 88.4 89.6 87.2 2.9 45.1 89.8 87.0 1.3 23.3 ConvS 84.9 66.3 7.5 24.1 88.2 74.5 5.7 10.9 DeConvL 42.2 89.3 87.2 2.6 45.0 63.6 89.6 86.9 1.1 23.2 DeConvS 11.6 85.6 68.2 7.1 26.0 9.3 88.3 75.7 4.6 12.1 UNetL 87.1 83.6 1.3 41.4 88.0 82.4 1.8 18.8 UNetS 85.5 72.7 4.6 30.5 88.1 75.8 4.3 12.2 VGG16 ConvL 90.3 90.1 86.0 6.2 50.3 90.2 88.5 0.9 21.9 ConvS 87.4 59.6 12 23.9 89.9 77.8 4.8 11.1 DeConvL 35.7 89.7 85.5 6.8 49.8 66.6 89.7 88.2 1.0 21.4 DeConvS 7.9 86.8 66.5 11 30.8 8.1 90.0 78.4 4.7 11.8 UNetL 87.4 83.4 4.4 47.7 89.0 86.2 1.5 19.6 UNetS 87.4 71.2 8.2 35.5 89.0 80.2 3.5 13.7 VGG19 ConvL 90.5 89.8 77.7 19 41.7 90.4 88.1 1.8 23.9 ConvS 87.3 52.7 17 16.7 89.6 74.5 9.0 10.3 DeConvL 36.0 86.3 78.4 18 42.4 64.2 90.4 88.5 1.4 24.3 DeConvS 12.0 86.5 58.2 18 22.2 12.4 89.7 75.2 8.6 11.0 UNetL 86.3 82.1 4.8 46.0 89.1 85.0 2.7 20.8 UNetS 86.3 66.4 13 30.4 89.2 77.1 7.3 12.9 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+Base Model NF CA 1% BER 0.5% BER PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP ResNet18 ConvL 92.6 89.8 87.8 1.7 48.8 90.4 87.9 2.2 17.8 ConvS 88.2 59.5 11 20.6 91.7 78.4 8.3 8.3 DeConvL 38.9 89.6 88.5 0.8 49.6 70.1 90.2 90.0 0.2 19.9 DeConvS 12.4 82.9 68.8 6.4 29.9 11.6 84.1 79.9 3.6 9.8 UNetL 86.6 84.6 0.8 45.6 89.7 86.3 2.4 16.2 UNetS 84.4 68.8 6.0 29.8 90.9 80.7 5.8 10.7 ResNet50 ConvL 92.6 85.5 53.2 22 27.1 90.3 86.5 3.2 25.5 ConvS 85.2 34.6 14 8.5 90.8 73.3 8.7 12.3 DeConvL 26.1 87.4 63.3 21 37.2 61.0 89.5 87.2 2.5 26.2 DeConvS 9.4 82.4 42.2 17 16.1 10.3 90.3 75.5 8.1 14.5 UNetL 86.2 75.5 12 49.4 89.9 83.9 3.6 22.9 UNetS 77.3 56.2 19 30.1 89.7 76.1 7.2 15.1 VGG11 ConvL 88.4 89.6 87.2 2.9 45.1 89.8 87.0 1.3 23.3 ConvS 84.9 66.3 7.5 24.1 88.2 74.5 5.7 10.9 DeConvL 42.2 89.3 87.2 2.6 45.0 63.6 89.6 86.9 1.1 23.2 DeConvS 11.6 85.6 68.2 7.1 26.0 9.3 88.3 75.7 4.6 12.1 UNetL 87.1 83.6 1.3 41.4 88.0 82.4 1.8 18.8 UNetS 85.5 72.7 4.6 30.5 88.1 75.8 4.3 12.2 VGG16 ConvL 90.3 90.1 86.0 6.2 50.3 90.2 88.5 0.9 21.9 ConvS 87.4 59.6 12 23.9 89.9 77.8 4.8 11.1 DeConvL 35.7 89.7 85.5 6.8 49.8 66.6 89.7 88.2 1.0 21.4 DeConvS 7.9 86.8 66.5 11 30.8 8.1 90.0 78.4 4.7 11.8 UNetL 87.4 83.4 4.4 47.7 89.0 86.2 1.5 19.6 UNetS 87.4 71.2 8.2 35.5 89.0 80.2 3.5 13.7 VGG19 ConvL 90.5 89.8 77.7 19 41.7 90.4 88.1 1.8 23.9 ConvS 87.3 52.7 17 16.7 89.6 74.5 9.0 10.3 DeConvL 36.0 86.3 78.4 18 42.4 64.2 90.4 88.5 1.4 24.3 DeConvS 12.0 86.5 58.2 18 22.2 12.4 89.7 75.2 8.6 11.0 UNetL 86.3 82.1 4.8 46.0 89.1 85.0 2.7 20.8 UNetS 86.3 66.4 13 30.4 89.2 77.1 7.3 12.9 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 (a) CIFAR-10, Bit Error Rate
@@
 F.2 GTSRB
-Base Model NF C.A.
-1% B.E.R.
-0.5% B.E.R.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ConvL 95.5 95.7 91.1 4.7 54.2 93.4 89.5 1.9 14.3 ConvS 94.4 68.6 12 31.7 94.8 87.7 4.2 12.4 DeConvL 36.9 95.6 91.3 4.3 54.4 75.2 95.4 93.4 1.1 18.1 DeConvS 16.0 95.7 78.1 9.1 41.2 12.7 95.8 90.1 3.3 14.9 UNetL 96.2 93.8 1.0 56.9 96.2 93.5 1.6 18.3 UNetS 95.9 85.1 6.9 48.2 95.5 91.4 2.8 16.2 ResNet50 ConvL 95.0 95.6 71.6 20 42.1 94.6 90.6 3.7 16.6 ConvS 94.8 50.5 22 21.0 95.4 84.5 8.5 10.5 DeConvL 29.5 94.9 71.6 21 42.0 74.0 94.7 91.6 2.9 17.6 DeConvS 16.9 93.0 56.4 17 26.9 13.0 94.6 87.4 5.9 13.5 UNetL 94.5 80.6 15 51.1 96.5 93.7 2.3 19.7 UNetS 94.7 64.7 22 35.2 95.9 90.6 4.8 16.7 VGG11 ConvL 91.9 94.8 85.7 7.2 50.9 93.9 92.6 0.7 27.7 ConvS 91.1 62.2 11 27.3 90.9 80.5 3.5 15.7 DeConvL 34.9 95.0 84.6 7.6 49.7 64.9 93.6 91.9 0.6 27.1 DeConvS 12.4 92.4 67.5 11 32.6 10.8 92.3 83.1 3.7 18.2 UNetL 92.2 83.2 6.0 48.3 94.8 90.6 1.7 25.7 UNetS 94.7 73.4 10 38.5 94.6 88.9 2.2 24.1 VGG16 ConvL 95.2 96.3 72.4 12 57.3 95.6 93.2 1.8 34.4 ConvS 94.1 39.8 13 24.6 94.3 82.2 6.2 23.4 DeConvL 15.1 96.4 72.0 12 56.9 58.8 95.6 93.1 2.0 34.3 DeConvS 6.8 93.8 50.9 13 35.8 8.9 95.1 84.0 5.3 25.2 UNetL 95.8 78.6 11 63.5 96.0 92.8 2.0 34.0 UNetS 94.3 63.3 14 48.1 95.4 87.8 3.6 29.0 VGG19 ConvL 95.5 96.0 88.3 7.2 51.7 95.6 93.4 2.1 24.2 ConvS 93.8 69.0 14 32.4 94.9 87.0 4.4 17.8 DeConvL 36.6 95.4 87.2 7.5 50.6 69.1 95.5 92.4 2.2 23.3 DeConvS 6.8 94.5 73.1 12 36.5 11.1 95.5 88.8 3.7 19.7 UNetL 95.4 88.2 6.7 51.7 94.9 91.7 2.5 22.6 UNetS 94.6 80.6 9.0 44.1 96.5 90.8 3.4 21.6 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+Base Model NF CA 1% BER 0.5% BER PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP ResNet18 ConvL 95.5 95.7 91.1 4.7 54.2 93.4 89.5 1.9 14.3 ConvS 94.4 68.6 12 31.7 94.8 87.7 4.2 12.4 DeConvL 36.9 95.6 91.3 4.3 54.4 75.2 95.4 93.4 1.1 18.1 DeConvS 16.0 95.7 78.1 9.1 41.2 12.7 95.8 90.1 3.3 14.9 UNetL 96.2 93.8 1.0 56.9 96.2 93.5 1.6 18.3 UNetS 95.9 85.1 6.9 48.2 95.5 91.4 2.8 16.2 ResNet50 ConvL 95.0 95.6 71.6 20 42.1 94.6 90.6 3.7 16.6 ConvS 94.8 50.5 22 21.0 95.4 84.5 8.5 10.5 DeConvL 29.5 94.9 71.6 21 42.0 74.0 94.7 91.6 2.9 17.6 DeConvS 16.9 93.0 56.4 17 26.9 13.0 94.6 87.4 5.9 13.5 UNetL 94.5 80.6 15 51.1 96.5 93.7 2.3 19.7 UNetS 94.7 64.7 22 35.2 95.9 90.6 4.8 16.7 VGG11 ConvL 91.9 94.8 85.7 7.2 50.9 93.9 92.6 0.7 27.7 ConvS 91.1 62.2 11 27.3 90.9 80.5 3.5 15.7 DeConvL 34.9 95.0 84.6 7.6 49.7 64.9 93.6 91.9 0.6 27.1 DeConvS 12.4 92.4 67.5 11 32.6 10.8 92.3 83.1 3.7 18.2 UNetL 92.2 83.2 6.0 48.3 94.8 90.6 1.7 25.7 UNetS 94.7 73.4 10 38.5 94.6 88.9 2.2 24.1 VGG16 ConvL 95.2 96.3 72.4 12 57.3 95.6 93.2 1.8 34.4 ConvS 94.1 39.8 13 24.6 94.3 82.2 6.2 23.4 DeConvL 15.1 96.4 72.0 12 56.9 58.8 95.6 93.1 2.0 34.3 DeConvS 6.8 93.8 50.9 13 35.8 8.9 95.1 84.0 5.3 25.2 UNetL 95.8 78.6 11 63.5 96.0 92.8 2.0 34.0 UNetS 94.3 63.3 14 48.1 95.4 87.8 3.6 29.0 VGG19 ConvL 95.5 96.0 88.3 7.2 51.7 95.6 93.4 2.1 24.2 ConvS 93.8 69.0 14 32.4 94.9 87.0 4.4 17.8 DeConvL 36.6 95.4 87.2 7.5 50.6 69.1 95.5 92.4 2.2 23.3 DeConvS 6.8 94.5 73.1 12 36.5 11.1 95.5 88.8 3.7 19.7 UNetL 95.4 88.2 6.7 51.7 94.9 91.7 2.5 22.6 UNetS 94.6 80.6 9.0 44.1 96.5 90.8 3.4 21.6 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 (a) GTSRB, Bit Error Rate
@@
 F.3 ImageNet-10
-| Base Model | NF | C.A. | 0.5% B.E.R. | |||
-| P.A. | C.A. (NF) | P.A. (NF) | R.P. | |||
+| Base Model | NF | CA | 0.5% BER | |||
+| PA | CA (NF) | PA (NF) | RP | |||
 | ResNet18 | ConvL | 92.2 | 72.3 7.0 | 94.0 | 88.0 2.0 | 15.7 |
@@
 | UNetS | 87.6 | 65.9 10 | 28.7 | |||
-[Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
+Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
 F.4 CIFAR-100
@@
 Base Model NF C.A.
-1% B.E.R.
-0.5% B.E.R.
-0.35% B.E.R.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ConvL 73.7 54.8 11.0 7.7 6.4 65.2 39.0 7.1 18.1 69.4 42.9 6.2 11.4 ConvS 49.7 4.2 2.2 -0.4 70.0 24.5 7.6 3.6 72.1 35.1 7.3 3.7 DeConvL 4.6 55.2 11.9 8.2 7.3 20.9 66.3 38.2 6.9 17.3 31.4 69.2 42.9 5.5 11.4 DeConvS 2.9 32.7 4.0 2.2 -0.6 7.4 68.2 25.9 6.8 5 7.6 71.6 35.8 5.5 4.4 UNetL 50.6 14.5 8.9 10.0 66.2 40.1 6.4 19.2 70.3 46.3 5.5 14.9 UNetS 26.8 4.6 2.5 -0.0 67.1 28.8 6.8 7.9 70.9 38.3 6.4 6.9 ResNet50 ConvL 73.5 63.5 3.2 1.7 0.1 68.4 28.8 6.7 7.6 72.0 40.8 7.5 5.1 ConvS 65.5 3.2 1.6 0.1 71.9 23.1 6.9 1.9 73.0 37.4 8.0 1.7 DeConvL 3.0 59.6 3.2 1.7 0.2 21.3 68.1 28.6 7.0 7.4 35.7 71.7 41.7 7.7 6.1 DeConvS 1.8 61.1 3.2 1.7 0.1 7.0 70.3 25.0 6.7 3.7 8.6 72.8 38.9 7.9 3.3 UNetL 39.0 5.0 1.7 1.9 66.6 36.5 6.2 15.3 70.8 45.3 6.7 9.6 UNetS 47.7 3.4 1.8 0.3 69.1 26.1 6.6 4.8 72.6 39.6 7.8 3.9 VGG11 ConvL 64.8 58.3 19.7 11 11.5 63.1 38.8 9.3 15.0 63.9 42.4 9.0 11.1 ConvS 56.6 10.4 7.4 2.2 62.7 27.9 10 4.0 63.9 41.8 8.3 10.5 DeConvL 8.2 60.3 21.2 11 13.0 23.9 63.9 40.0 9.0 16.2 31.3 64.0 42.8 9.1 11.5 DeConvS 5.7 58.3 11.8 7.9 3.5 9.4 61.9 29.8 9.9 5.9 10 63.5 36.1 10 4.8 UNetL 51.1 22.1 8.2 13.9 61.8 37.8 9.0 13.9 63.5 40.9 9.3 9.6 UNetS 51.9 13.1 7.9 4.9 61.7 29.8 9.7 6.0 63.8 35.7 9.9 4.5 VGG16 ConvL 67.8 51.4 19.2 6.0 12.6 61.8 41.1 5.6 18.7 64.9 44.9 5.3 13.8 ConvS 44.3 6.7 2.3 0.1 63.8 27.5 6.8 5.1 66.0 36.3 6.1 5.1 DeConvL 7.0 53.1 20.8 6.2 14.2 22.4 62.8 42.1 5.5 19.8 31.1 65.0 46.6 5.2 15.5 DeConvS 3.5 23.5 4.8 1.7 -1.8 7.0 62.1 29.9 6.7 7.5 7.2 64.9 38.1 6.3 7.0 UNetL 50.2 25.3 1.7 18.7 61.7 41.3 5.0 18.9 64.8 46.8 4.6 15.7 UNetS 27.7 9.9 2.1 3.3 61.6 31.3 6.3 8.9 65.0 39.8 5.9 8.7 VGG19 ConvL 67.8 59.4 29.2 8.1 18.6 65.6 46.5 6.8 12.5 66.9 49.2 7.4 7.0 ConvS 63.7 14.4 5.1 3.8 66.6 38.3 6.8 4.2 67.7 45.3 8.5 3.2 DeConvL 10.6 60.1 29.6 8.5 19.0 34.0 65.7 46.9 7.1 12.9 42.1 67.3 49.8 7.6 7.6 DeConvS 4.3 60.9 16.1 6.0 5.6 9.6 66.5 39.0 3.7 5.0 9.4 67.7 45.7 8.4 3.6 UNetL 58.7 30.2 8.2 19.6 65.5 46.9 6.5 12.9 67.4 50.0 7.5 7.9 UNetS 59.1 18.0 6.2 7.4 66.3 40.1 8.0 6.1 67.5 46.6 8.4 4.5 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+1% BER 0.5% BER 0.35% BER PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP ResNet18 ConvL 73.7 54.8 11.0 7.7 6.4 65.2 39.0 7.1 18.1 69.4 42.9 6.2 11.4 ConvS 49.7 4.2 2.2 -0.4 70.0 24.5 7.6 3.6 72.1 35.1 7.3 3.7 DeConvL 4.6 55.2 11.9 8.2 7.3 20.9 66.3 38.2 6.9 17.3 31.4 69.2 42.9 5.5 11.4 DeConvS 2.9 32.7 4.0 2.2 -0.6 7.4 68.2 25.9 6.8 5 7.6 71.6 35.8 5.5 4.4 UNetL 50.6 14.5 8.9 10.0 66.2 40.1 6.4 19.2 70.3 46.3 5.5 14.9 UNetS 26.8 4.6 2.5 -0.0 67.1 28.8 6.8 7.9 70.9 38.3 6.4 6.9 ResNet50 ConvL 73.5 63.5 3.2 1.7 0.1 68.4 28.8 6.7 7.6 72.0 40.8 7.5 5.1 ConvS 65.5 3.2 1.6 0.1 71.9 23.1 6.9 1.9 73.0 37.4 8.0 1.7 DeConvL 3.0 59.6 3.2 1.7 0.2 21.3 68.1 28.6 7.0 7.4 35.7 71.7 41.7 7.7 6.1 DeConvS 1.8 61.1 3.2 1.7 0.1 7.0 70.3 25.0 6.7 3.7 8.6 72.8 38.9 7.9 3.3 UNetL 39.0 5.0 1.7 1.9 66.6 36.5 6.2 15.3 70.8 45.3 6.7 9.6 UNetS 47.7 3.4 1.8 0.3 69.1 26.1 6.6 4.8 72.6 39.6 7.8 3.9 VGG11 ConvL 64.8 58.3 19.7 11 11.5 63.1 38.8 9.3 15.0 63.9 42.4 9.0 11.1 ConvS 56.6 10.4 7.4 2.2 62.7 27.9 10 4.0 63.9 41.8 8.3 10.5 DeConvL 8.2 60.3 21.2 11 13.0 23.9 63.9 40.0 9.0 16.2 31.3 64.0 42.8 9.1 11.5 DeConvS 5.7 58.3 11.8 7.9 3.5 9.4 61.9 29.8 9.9 5.9 10 63.5 36.1 10 4.8 UNetL 51.1 22.1 8.2 13.9 61.8 37.8 9.0 13.9 63.5 40.9 9.3 9.6 UNetS 51.9 13.1 7.9 4.9 61.7 29.8 9.7 6.0 63.8 35.7 9.9 4.5 VGG16 ConvL 67.8 51.4 19.2 6.0 12.6 61.8 41.1 5.6 18.7 64.9 44.9 5.3 13.8 ConvS 44.3 6.7 2.3 0.1 63.8 27.5 6.8 5.1 66.0 36.3 6.1 5.1 DeConvL 7.0 53.1 20.8 6.2 14.2 22.4 62.8 42.1 5.5 19.8 31.1 65.0 46.6 5.2 15.5 DeConvS 3.5 23.5 4.8 1.7 -1.8 7.0 62.1 29.9 6.7 7.5 7.2 64.9 38.1 6.3 7.0 UNetL 50.2 25.3 1.7 18.7 61.7 41.3 5.0 18.9 64.8 46.8 4.6 15.7 UNetS 27.7 9.9 2.1 3.3 61.6 31.3 6.3 8.9 65.0 39.8 5.9 8.7 VGG19 ConvL 67.8 59.4 29.2 8.1 18.6 65.6 46.5 6.8 12.5 66.9 49.2 7.4 7.0 ConvS 63.7 14.4 5.1 3.8 66.6 38.3 6.8 4.2 67.7 45.3 8.5 3.2 DeConvL 10.6 60.1 29.6 8.5 19.0 34.0 65.7 46.9 7.1 12.9 42.1 67.3 49.8 7.6 7.6 DeConvS 4.3 60.9 16.1 6.0 5.6 9.6 66.5 39.0 3.7 5.0 9.4 67.7 45.7 8.4 3.6 UNetL 58.7 30.2 8.2 19.6 65.5 46.9 6.5 12.9 67.4 50.0 7.5 7.9 UNetS 59.1 18.0 6.2 7.4 66.3 40.1 8.0 6.1 67.5 46.6 8.4 4.5 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 (a) CIFAR-100, Bit Error Rate
@@
 In addition, we can observe that some NeuralFuse generators have downward applicability if base models have a similar architecture.
-In other words, if we try to transfer a generator trained on a large B.E.R.
-(e.g., 1%) to a model with a small B.E.R.
-(e.g., 0.5%), the performance will be better than that of a generator trained with the original B.E.R.
-(e.g., 0.5%).
-For example, in Table 16, we could find that if we use VGG19 as a source model to train the generator ConvL (1%), the generator could deliver better performance (in terms of P.A.
-(NF)) when applied to similar base models (e.g., VGG11, VGG16, or VGG19) under a 0.5% B.E.R., compared to using itself as a source model (shown in Table 12).
-We conjecture that this is because the generators trained on a larger B.E.R.
-can also cover the error patterns of a smaller B.E.R., and thus they have better generalizability across smaller B.E.Rs.
+In other words, if we try to transfer a generator trained on a large BER (e.g., 1%) to a model with a small BER (e.g., 0.5%), the performance will be better than that of a generator trained with the original BER (e.g., 0.5%).
+For example, in Table 16, we could find that if we use VGG19 as a source model to train the generator ConvL (1%), the generator could deliver better performance (in terms of PA (NF)) when applied to similar base models (e.g., VGG11, VGG16, or VGG19) under a 0.5% BER, compared to using itself as a source model (shown in Table 12).
+We conjecture that this is because the generators trained on a larger BER can also cover the error patterns of a smaller BER, and thus they have better generalizability across smaller B.E.Rs.
 To further improve the transferability to cross-architecture target models, we also conduct an experiment in Sec.
@@
 G.1 CIFAR-10
-The results of CIFAR-10 in which NeuralFuse is trained at 1% B.E.R. are shown in Table 16.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (1%) UNetL (1%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ResNet18 0.5% 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 ResNet50 1% 92.6 26.1 9.4 89.5 36.0 19 9.9 85.2 38.8 19 12.7 0.5% 61.0 10.3 75.1 10 14.1 77.1 5.0 16.1 VGG11 1% 88.4 42.2 11.6 88.4 62.5 8.4 20.3 76.8 61.1 8.5 18.9 0.5% 63.6 9.3 81.0 4.6 17.4 73.7 3.0 10.1 VGG16 1% 90.3 35.7 7.9 89.6 63.3 18 27.6 85.2 59.9 16 24.2 0.5% 66.6 8.1 85.0 3.4 18.4 80.2 4.5 13.6 VGG19 1% 90.5 36.0 12.0 89.6 50.7 22 14.7 85.3 51.1 16 15.1 0.5% 64.2 12.4 80.2 8.7 16.0 76.5 7.8 12.3 VGG19 ResNet18 1% 92.6 38.9 12.4 89.8 61.0 17 22.1 87.0 69.7 11 30.8 0.5% 70.1 11.6 86.1 6.9 16.0 84.2 3.0 14.1 ResNet50 1% 92.6 26.1 9.4 89.9 34.0 19 7.9 87.0 44.2 17 18.1 0.5% 61.0 10.3 76.5 10 15.5 80.7 4.2 19.7 VGG11 1% 88.4 42.2 11.6 89.7 76.5 7.0 34.3 87.1 79.9 5.6 37.7 0.5% 63.6 9.3 88.0 2.1 24.4 85.4 0.8 21.8 VGG16 1% 90.3 35.7 7.9 89.6 75.5 12 39.8 87.2 78.9 7.8 43.2 0.5% 66.6 8.1 88.9 0.6 22.3 86.2 0.3 19.6 VGG19 0.5% 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+The results of CIFAR-10 in which NeuralFuse is trained at 1% BER are shown in Table 16.
+SM TM BER CA PA ConvL (1%) UNetL (1%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.5% 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 ResNet50 1% 92.6 26.1 9.4 89.5 36.0 19 9.9 85.2 38.8 19 12.7 0.5% 61.0 10.3 75.1 10 14.1 77.1 5.0 16.1 VGG11 1% 88.4 42.2 11.6 88.4 62.5 8.4 20.3 76.8 61.1 8.5 18.9 0.5% 63.6 9.3 81.0 4.6 17.4 73.7 3.0 10.1 VGG16 1% 90.3 35.7 7.9 89.6 63.3 18 27.6 85.2 59.9 16 24.2 0.5% 66.6 8.1 85.0 3.4 18.4 80.2 4.5 13.6 VGG19 1% 90.5 36.0 12.0 89.6 50.7 22 14.7 85.3 51.1 16 15.1 0.5% 64.2 12.4 80.2 8.7 16.0 76.5 7.8 12.3 VGG19 ResNet18 1% 92.6 38.9 12.4 89.8 61.0 17 22.1 87.0 69.7 11 30.8 0.5% 70.1 11.6 86.1 6.9 16.0 84.2 3.0 14.1 ResNet50 1% 92.6 26.1 9.4 89.9 34.0 19 7.9 87.0 44.2 17 18.1 0.5% 61.0 10.3 76.5 10 15.5 80.7 4.2 19.7 VGG11 1% 88.4 42.2 11.6 89.7 76.5 7.0 34.3 87.1 79.9 5.6 37.7 0.5% 63.6 9.3 88.0 2.1 24.4 85.4 0.8 21.8 VGG16 1% 90.3 35.7 7.9 89.6 75.5 12 39.8 87.2 78.9 7.8 43.2 0.5% 66.6 8.1 88.9 0.6 22.3 86.2 0.3 19.6 VGG19 0.5% 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 G.2 GTSRB
-In Tables 17 and 18, we show the results on GTSRB in which NeuralFuse is trained at 1.5% and 1% B.E.R., respectively.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (1.5%) UNetL (1.5%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ResNet18 1% 95.5 36.9 16.0 95.7 93.9 1.9 57.0 94.9 94.4 0.4 57.5 0.5% 75.2 12.7 95.7 0.2 20.5 94.8 0.2 19.6 ResNet50 1% 95.0 29.5 16.9 94.4 37.0 22 7.5 94.4 47.1 23 17.6 0.5% 74.0 13.0 77.5 13 3.5 84.8 9.5 10.8 VGG11 1% 91.9 34.9 12.4 92.8 45.2 10 10.3 91.4 50.5 13 15.6 0.5% 64.9 10.8 79.4 5.8 14.5 83.9 4.2 19.0 VGG16 1% 95.2 15.1 6.8 95.4 31.1 13 15.9 94.6 36.8 12 21.7 0.5% 58.8 8.9 84.5 8.3 25.8 86.0 8.6 27.2 VGG19 1% 95.5 36.6 6.8 95.0 56.4 15 19.8 94.3 60.8 15 24.2 0.5% 69.1 11.1 86.9 3.4 17.8 87.7 3.8 18.6 VGG19 ResNet18 1% 95.5 36.9 16.0 88.4 50.3 12 13.4 92.8 63.7 16 26.8 0.5% 75.2 12.7 77.9 7.4 2.7 87.5 3.9 12.3 ResNet50 1% 95.0 29.5 16.9 87.5 29.7 17 0.2 92.5 40.4 21 10.9 0.5% 74.0 13.0 67.9 17 -6.1 77.5 15 3.5 VGG11 1% 91.9 34.9 12.4 89.7 47.1 11 12.2 93.5 60.0 12 25.1 0.5% 64.9 10.8 76.3 5.1 11.4 86.0 3.8 21.1 VGG16 1% 95.2 15.1 6.8 93.0 29.2 15 14.1 93.0 38.5 16 23.4 0.5% 58.8 8.9 75.7 12 16.9 79.9 8.3 21.1 VGG19 1% 95.5 36.6 6.8 95.1 87.4 6.0 50.8 94.6 88.7 5.0 52.1 0.5% 69.1 11.1 92.4 2.4 23.3 92.4 2.2 23.3 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (1%) UNetL (1%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ResNet18 0.5% 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 ResNet50 1% 95.0 29.5 16.9 94.5 35.6 21 6.1 95.6 42.6 23 13.1 0.5% 74.0 13.0 78.8 13 4.8 87.3 9.0 13.3 VGG11 1% 91.9 34.9 12.4 93.1 45.8 11 10.9 94.0 47.1 14 12.2 0.5% 64.9 10.8 81.8 5.0 16.9 84.2 4.8 19.3 VGG16 1% 95.2 15.1 6.8 95.5 26.5 12 11.4 95.5 32.4 11 17.3 0.5% 58.8 8.9 82.2 9.0 23.4 85.4 6.7 26.6 VGG19 1% 95.5 36.6 6.8 94.9 53.2 14 16.6 95.6 60.9 15 24.3 0.5% 69.1 11.1 85.4 4.5 16.3 87.5 3.7 18.4 VGG19 ResNet18 1% 95.5 36.9 16.0 93.7 53.1 16 16.2 95.0 63.4 18 26.5 0.5% 75.2 12.7 83.9 7.6 8.7 89.7 4.8 14.5 ResNet50 1% 95.0 29.5 16.9 92.8 30.6 18 1.1 95.4 38.9 22 9.4 0.5% 74.0 13.0 74.7 18 0.7 81.5 16 7.5 VGG11 1% 91.9 34.9 12.4 93.7 50.6 11 15.7 95.1 58.9 15 24.0 0.5% 64.9 10.8 82.3 5.1 17.4 87.5 3.7 22.6 VGG16 1% 95.2 15.1 6.8 95.2 27.8 15 12.7 95.2 33.5 14 18.4 0.5% 58.8 8.9 79.0 12 20.2 81.8 7.8 23.0 VGG19 0.5% 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+In Tables 17 and 18, we show the results on GTSRB in which NeuralFuse is trained at 1.5% and 1% BER, respectively.
+SM TM BER CA PA ConvL (1.5%) UNetL (1.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 1% 95.5 36.9 16.0 95.7 93.9 1.9 57.0 94.9 94.4 0.4 57.5 0.5% 75.2 12.7 95.7 0.2 20.5 94.8 0.2 19.6 ResNet50 1% 95.0 29.5 16.9 94.4 37.0 22 7.5 94.4 47.1 23 17.6 0.5% 74.0 13.0 77.5 13 3.5 84.8 9.5 10.8 VGG11 1% 91.9 34.9 12.4 92.8 45.2 10 10.3 91.4 50.5 13 15.6 0.5% 64.9 10.8 79.4 5.8 14.5 83.9 4.2 19.0 VGG16 1% 95.2 15.1 6.8 95.4 31.1 13 15.9 94.6 36.8 12 21.7 0.5% 58.8 8.9 84.5 8.3 25.8 86.0 8.6 27.2 VGG19 1% 95.5 36.6 6.8 95.0 56.4 15 19.8 94.3 60.8 15 24.2 0.5% 69.1 11.1 86.9 3.4 17.8 87.7 3.8 18.6 VGG19 ResNet18 1% 95.5 36.9 16.0 88.4 50.3 12 13.4 92.8 63.7 16 26.8 0.5% 75.2 12.7 77.9 7.4 2.7 87.5 3.9 12.3 ResNet50 1% 95.0 29.5 16.9 87.5 29.7 17 0.2 92.5 40.4 21 10.9 0.5% 74.0 13.0 67.9 17 -6.1 77.5 15 3.5 VGG11 1% 91.9 34.9 12.4 89.7 47.1 11 12.2 93.5 60.0 12 25.1 0.5% 64.9 10.8 76.3 5.1 11.4 86.0 3.8 21.1 VGG16 1% 95.2 15.1 6.8 93.0 29.2 15 14.1 93.0 38.5 16 23.4 0.5% 58.8 8.9 75.7 12 16.9 79.9 8.3 21.1 VGG19 1% 95.5 36.6 6.8 95.1 87.4 6.0 50.8 94.6 88.7 5.0 52.1 0.5% 69.1 11.1 92.4 2.4 23.3 92.4 2.2 23.3 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
+SM TM BER CA PA ConvL (1%) UNetL (1%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.5% 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 ResNet50 1% 95.0 29.5 16.9 94.5 35.6 21 6.1 95.6 42.6 23 13.1 0.5% 74.0 13.0 78.8 13 4.8 87.3 9.0 13.3 VGG11 1% 91.9 34.9 12.4 93.1 45.8 11 10.9 94.0 47.1 14 12.2 0.5% 64.9 10.8 81.8 5.0 16.9 84.2 4.8 19.3 VGG16 1% 95.2 15.1 6.8 95.5 26.5 12 11.4 95.5 32.4 11 17.3 0.5% 58.8 8.9 82.2 9.0 23.4 85.4 6.7 26.6 VGG19 1% 95.5 36.6 6.8 94.9 53.2 14 16.6 95.6 60.9 15 24.3 0.5% 69.1 11.1 85.4 4.5 16.3 87.5 3.7 18.4 VGG19 ResNet18 1% 95.5 36.9 16.0 93.7 53.1 16 16.2 95.0 63.4 18 26.5 0.5% 75.2 12.7 83.9 7.6 8.7 89.7 4.8 14.5 ResNet50 1% 95.0 29.5 16.9 92.8 30.6 18 1.1 95.4 38.9 22 9.4 0.5% 74.0 13.0 74.7 18 0.7 81.5 16 7.5 VGG11 1% 91.9 34.9 12.4 93.7 50.6 11 15.7 95.1 58.9 15 24.0 0.5% 64.9 10.8 82.3 5.1 17.4 87.5 3.7 22.6 VGG16 1% 95.2 15.1 6.8 95.2 27.8 15 12.7 95.2 33.5 14 18.4 0.5% 58.8 8.9 79.0 12 20.2 81.8 7.8 23.0 VGG19 0.5% 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 G.3 CIFAR-100
-In Tables 19 and 20, we show results on CIFAR-100 with NeuralFuse trained at 1% and 0.5% B.E.R., respectively.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (1%) UNetL (1%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ResNet18 0.5% 73.7 20.9 7.4 54.8 35.8 5.2 14.9 50.6 39.3 2.8 18.4 0.35% 31.4 7.6 41.7 3.7 10.3 43.3 1.4 11.9 ResNet50 1% 73.5 3.0 1.8 44.9 2.2 2.0 -0.8 41.5 2.4 1.9 -0.6 0.5% 21.3 7.0 15.9 8.2 -5.4 17.1 7.1 -4.2 0.35% 35.7 8.6 23.7 7.1 -12.0 26.2 5.6 -9.5 VGG11 1% 64.8 8.2 5.7 41.2 9.8 5.6 1.6 37.5 10.2 5.1 2.0 0.5% 23.9 9.4 24.2 5.9 0.3 24.5 4.7 0.6 0.35% 31.3 10.0 29.0 5.4 -2.3 28.2 4.5 -3.1 VGG16 1% 67.8 7.0 3.5 44.0 7.9 3.7 0.9 39.5 10.1 4.5 3.1 0.5% 22.4 7.0 22.4 7.6 0.0 26.3 5.3 3.9 0.35% 31.1 7.2 28.1 5.9 -3.0 30.6 3.6 -0.5 VGG19 1% 67.8 10.6 4.3 44.2 13.5 6.1 2.9 40.7 15.6 6.2 5.0 0.5% 34.0 9.6 27.9 4.8 -6.1 29.3 4.6 -4.7 0.35% 42.1 9.4 33.2 48 -8.9 32.8 3.9 -9.3 VGG19 ResNet18 1% 73.7 4.6 2.9 55.5 5.8 3.7 1.2 57.3 6.8 4.4 2.2 0.5% 20.9 7.4 24.6 6.3 3.7 28.1 5.9 7.2 0.35% 31.4 7.6 31.1 5.0 -0.3 36.4 4.5 5.0 ResNet50 1% 73.5 3.0 1.8 56.1 2.8 2.1 -0.2 56.1 3.7 2.4 0.7 0.5% 21.3 7.0 18.9 8.6 -2.4 22.8 8.5 1.5 0.35% 35.7 8.6 28.7 8.2 -7.0 33.7 7.0 -2.0 VGG11 1% 64.8 8.2 5.7 52.8 12.3 8.4 4.1 53.9 15.4 9.4 7.2 0.5% 23.9 9.4 30.0 9.3 6.1 33.3 7.2 9.4 0.35% 31.3 10.0 36.5 7.7 5.2 38.8 6.5 7.5 VGG16 1% 67.8 7.0 3.5 53.6 11.2 4.4 4.2 55.2 13.6 5.2 6.6 0.5% 22.4 7.0 32.4 7.3 10.0 35.9 6.2 13.5 0.35% 31.1 7.2 39.4 6.3 8.3 42.4 4.9 11.3 VGG19 0.5% 67.8 34.0 9.6 59.4 50.2 3.1 16.2 58.7 49.1 3.5 15.1 0.35% 42.1 9.4 53.1 2.3 11.0 52.0 3.1 9.9 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (0.5%) UNetL (0.5%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 ResNet18 0.35% 73.7 31.4 7.6 65.2 47.7 4.9 16.3 66.2 49.2 4.1 17.8 ResNet50 0.5% 73.5 21.3 7.0 62.5 24.0 9.9 2.8 63.5 26.4 9.1 5.1 0.35% 35.7 8.6 36.3 8.9 0.6 39.4 8.1 3.7 VGG11 0.5% 64.8 23.9 9.4 59.2 33.0 9.8 9.2 61.1 34.2 9.8 10.3 0.35% 31.3 10.0 40.4 8.7 9.1 41.4 9.0 10.1 VGG16 0.5% 67.8 22.4 7.0 59.5 34.7 8.0 12.3 61.4 37.5 6.8 15.2 0.35% 31.1 7.2 42.9 6.0 11.8 45.3 4.9 14.2 VGG19 0.5% 67.8 34.0 9.6 61.6 43.7 6.2 9.6 62.0 45.0 6.3 11.0 0.35% 42.1 9.4 49.0 5.5 6.8 50.5 5.3 8.3 VGG19 ResNet18 0.5% 73.7 20.9 7.4 66.1 24.9 6.7 4.0 67.8 27.7 6.8 6.8 0.35% 31.4 7.6 34.4 5.4 3.0 38.1 5.6 6.7 ResNet50 0.5% 73.5 21.3 7.0 66.2 22.7 7.8 1.4 66.7 25.4 8.0 4.2 0.35% 35.7 8.6 35.5 7.7 -0.2 38.8 7.5 3.2 VGG11 0.5% 64.8 23.9 9.4 59.9 29.3 10 5.4 61.0 31.2 9.8 7.4 0.35% 31.3 10.0 36.6 9.5 5.3 38.1 9.0 6.8 VGG16 0.5% 67.8 22.4 7.0 62.5 30.8 7.3 8.4 62.6 33.0 7.3 10.7 0.35% 31.1 7.2 40.0 6.5 8.9 42.5 5.9 11.3 VGG19 0.35% 67.8 42.1 9.4 65.6 52.0 6.2 9.8 65.5 52.6 6.1 10.4 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+In Tables 19 and 20, we show results on CIFAR-100 with NeuralFuse trained at 1% and 0.5% BER, respectively.
+SM TM BER CA PA ConvL (1%) UNetL (1%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.5% 73.7 20.9 7.4 54.8 35.8 5.2 14.9 50.6 39.3 2.8 18.4 0.35% 31.4 7.6 41.7 3.7 10.3 43.3 1.4 11.9 ResNet50 1% 73.5 3.0 1.8 44.9 2.2 2.0 -0.8 41.5 2.4 1.9 -0.6 0.5% 21.3 7.0 15.9 8.2 -5.4 17.1 7.1 -4.2 0.35% 35.7 8.6 23.7 7.1 -12.0 26.2 5.6 -9.5 VGG11 1% 64.8 8.2 5.7 41.2 9.8 5.6 1.6 37.5 10.2 5.1 2.0 0.5% 23.9 9.4 24.2 5.9 0.3 24.5 4.7 0.6 0.35% 31.3 10.0 29.0 5.4 -2.3 28.2 4.5 -3.1 VGG16 1% 67.8 7.0 3.5 44.0 7.9 3.7 0.9 39.5 10.1 4.5 3.1 0.5% 22.4 7.0 22.4 7.6 0.0 26.3 5.3 3.9 0.35% 31.1 7.2 28.1 5.9 -3.0 30.6 3.6 -0.5 VGG19 1% 67.8 10.6 4.3 44.2 13.5 6.1 2.9 40.7 15.6 6.2 5.0 0.5% 34.0 9.6 27.9 4.8 -6.1 29.3 4.6 -4.7 0.35% 42.1 9.4 33.2 48 -8.9 32.8 3.9 -9.3 VGG19 ResNet18 1% 73.7 4.6 2.9 55.5 5.8 3.7 1.2 57.3 6.8 4.4 2.2 0.5% 20.9 7.4 24.6 6.3 3.7 28.1 5.9 7.2 0.35% 31.4 7.6 31.1 5.0 -0.3 36.4 4.5 5.0 ResNet50 1% 73.5 3.0 1.8 56.1 2.8 2.1 -0.2 56.1 3.7 2.4 0.7 0.5% 21.3 7.0 18.9 8.6 -2.4 22.8 8.5 1.5 0.35% 35.7 8.6 28.7 8.2 -7.0 33.7 7.0 -2.0 VGG11 1% 64.8 8.2 5.7 52.8 12.3 8.4 4.1 53.9 15.4 9.4 7.2 0.5% 23.9 9.4 30.0 9.3 6.1 33.3 7.2 9.4 0.35% 31.3 10.0 36.5 7.7 5.2 38.8 6.5 7.5 VGG16 1% 67.8 7.0 3.5 53.6 11.2 4.4 4.2 55.2 13.6 5.2 6.6 0.5% 22.4 7.0 32.4 7.3 10.0 35.9 6.2 13.5 0.35% 31.1 7.2 39.4 6.3 8.3 42.4 4.9 11.3 VGG19 0.5% 67.8 34.0 9.6 59.4 50.2 3.1 16.2 58.7 49.1 3.5 15.1 0.35% 42.1 9.4 53.1 2.3 11.0 52.0 3.1 9.9 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
+SM TM BER CA PA ConvL (0.5%) UNetL (0.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 0.35% 73.7 31.4 7.6 65.2 47.7 4.9 16.3 66.2 49.2 4.1 17.8 ResNet50 0.5% 73.5 21.3 7.0 62.5 24.0 9.9 2.8 63.5 26.4 9.1 5.1 0.35% 35.7 8.6 36.3 8.9 0.6 39.4 8.1 3.7 VGG11 0.5% 64.8 23.9 9.4 59.2 33.0 9.8 9.2 61.1 34.2 9.8 10.3 0.35% 31.3 10.0 40.4 8.7 9.1 41.4 9.0 10.1 VGG16 0.5% 67.8 22.4 7.0 59.5 34.7 8.0 12.3 61.4 37.5 6.8 15.2 0.35% 31.1 7.2 42.9 6.0 11.8 45.3 4.9 14.2 VGG19 0.5% 67.8 34.0 9.6 61.6 43.7 6.2 9.6 62.0 45.0 6.3 11.0 0.35% 42.1 9.4 49.0 5.5 6.8 50.5 5.3 8.3 VGG19 ResNet18 0.5% 73.7 20.9 7.4 66.1 24.9 6.7 4.0 67.8 27.7 6.8 6.8 0.35% 31.4 7.6 34.4 5.4 3.0 38.1 5.6 6.7 ResNet50 0.5% 73.5 21.3 7.0 66.2 22.7 7.8 1.4 66.7 25.4 8.0 4.2 0.35% 35.7 8.6 35.5 7.7 -0.2 38.8 7.5 3.2 VGG11 0.5% 64.8 23.9 9.4 59.9 29.3 10 5.4 61.0 31.2 9.8 7.4 0.35% 31.3 10.0 36.6 9.5 5.3 38.1 9.0 6.8 VGG16 0.5% 67.8 22.4 7.0 62.5 30.8 7.3 8.4 62.6 33.0 7.3 10.7 0.35% 31.1 7.2 40.0 6.5 8.9 42.5 5.9 11.3 VGG19 0.35% 67.8 42.1 9.4 65.6 52.0 6.2 9.8 65.5 52.6 6.1 10.4 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 G.4 Generator Ensembling
@@
 The results show that the overall performance is better than the results shown in Table 1, which means ensemble-based training can easily solve the performance degradation on cross-architecture target models.
-S.M.
-T.M.
-B.E.R.
-C.A.
-P.A.
-ConvL (1.5%) UNetL (1.5%) C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 + VGG19 ResNet18 1% 92.6 38.9 12.4 89.4 88.1 1.0 49.2 86.3 85.4 0.5 46.5 0.5% 70.1 11.6 89.2 0.2 19.1 86.1 0.2 16.0 ResNet50 1% 92.6 26.1 9.4 89.3 44.0 22 17.9 86.1 50.9 20 24.8 0.5% 61.0 10.3 80.3 6.7 19.3 78.6 3.9 17.6 VGG11 1% 88.4 42.2 11.6 89.1 77.0 5.6 34.8 85.9 82.3 4.1 40.1 0.5% 63.6 9.3 87.5 1.6 23.9 85.0 0.6 21.4 VGG16 1% 90.3 35.7 7.9 89.1 80.5 8.6 44.8 85.7 81.4 5.5 45.7 0.5% 66.6 8.1 88.2 0.7 21.6 85.0 0.7 18.4 VGG19 1% 90.5 36.0 12.0 89.2 75.1 17 39.1 86.1 83.0 3.4 47.0 0.5% 64.2 12.4 89.0 0.2 24.8 85.9 0.4 21.7 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+SM TM BER CA PA ConvL (1.5%) UNetL (1.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 + VGG19 ResNet18 1% 92.6 38.9 12.4 89.4 88.1 1.0 49.2 86.3 85.4 0.5 46.5 0.5% 70.1 11.6 89.2 0.2 19.1 86.1 0.2 16.0 ResNet50 1% 92.6 26.1 9.4 89.3 44.0 22 17.9 86.1 50.9 20 24.8 0.5% 61.0 10.3 80.3 6.7 19.3 78.6 3.9 17.6 VGG11 1% 88.4 42.2 11.6 89.1 77.0 5.6 34.8 85.9 82.3 4.1 40.1 0.5% 63.6 9.3 87.5 1.6 23.9 85.0 0.6 21.4 VGG16 1% 90.3 35.7 7.9 89.1 80.5 8.6 44.8 85.7 81.4 5.5 45.7 0.5% 66.6 8.1 88.2 0.7 21.6 85.0 0.7 18.4 VGG19 1% 90.5 36.0 12.0 89.2 75.1 17 39.1 86.1 83.0 3.4 47.0 0.5% 64.2 12.4 89.0 0.2 24.8 85.9 0.4 21.7 • Note.
+SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 Appendix H NeuralFuse on Reduced-precision Quantization and Random Bit Errors
@@
 4.6, we explore the robustness of NeuralFuse to low-precision quantization of model weights and consider the case of random bit errors.
-Here, we demonstrate that NeuralFuse can recover not only the accuracy drop due to reduced precision, but also the drop caused by low-voltage-induced bit errors (0.5% B.E.R.) under low precision.
-We selected two NeuralFuse generators (ConvL and UNetL) for our experiments, and these generators were trained with the corresponding base models (ResNet18 and VGG19) at 1% B.E.R.
-(CIFAR-10, GTSRB) and 0.5% B.E.R.
-(ImageNet-10).
+Here, we demonstrate that NeuralFuse can recover not only the accuracy drop due to reduced precision, but also the drop caused by low-voltage-induced bit errors (0.5% BER) under low precision.
+We selected two NeuralFuse generators (ConvL and UNetL) for our experiments, and these generators were trained with the corresponding base models (ResNet18 and VGG19) at 1% BER (CIFAR-10, GTSRB) and 0.5% BER (ImageNet-10).
 The experimental results are shown as follows: CIFAR-10 (Sec.
@@
 H.1 CIFAR-10
-Base #Bits C.A.
-P.A.
-ConvL (1%) UNetL (1%) Model C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 8 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 7 92.5 68.8 10.4 89.8 89.5 1.7 20.7 86.5 86.0 0.5 17.2 6 92.6 68.4 11.2 89.7 89.5 0.2 21.1 86.6 85.9 0.3 17.5 5 92.4 52.7 14.1 89.7 90.0 0.7 37.3 86.5 85.5 0.8 32.8 4 91.8 26.3 12.7 89.8 58.7 24.5 32.4 86.6 64.9 22.5 38.6 3 84.8 11.3 1.8 89.8 12.8 5.8 1.5 86.0 14.8 10.0 3.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 7 90.3 66.5 8.5 89.8 89.6 0.2 23.1 87.4 86.7 0.3 20.2 6 90.1 59.8 13.2 89.9 89.4 3.8 29.6 87.4 86.4 0.7 26.6 5 90.2 37.7 14.1 89.8 78.0 15.8 40.3 87.2 79.8 0.8 42.1 4 87.5 14.7 6.0 89.8 27.8 18.9 13.1 87.2 34.4 20.5 19.7 3 78.3 10.5 1.5 89.7 10.9 2.6 0.4 86.8 11.0 2.9 0.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+Base #Bits CA PA ConvL (1%) UNetL (1%) Model CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 8 92.6 70.1 11.6 89.8 89.5 0.2 19.4 86.6 86.2 0.3 16.1 7 92.5 68.8 10.4 89.8 89.5 1.7 20.7 86.5 86.0 0.5 17.2 6 92.6 68.4 11.2 89.7 89.5 0.2 21.1 86.6 85.9 0.3 17.5 5 92.4 52.7 14.1 89.7 90.0 0.7 37.3 86.5 85.5 0.8 32.8 4 91.8 26.3 12.7 89.8 58.7 24.5 32.4 86.6 64.9 22.5 38.6 3 84.8 11.3 1.8 89.8 12.8 5.8 1.5 86.0 14.8 10.0 3.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 90.5 64.2 12.4 89.8 89.6 8.7 25.4 87.4 86.8 0.4 22.6 7 90.3 66.5 8.5 89.8 89.6 0.2 23.1 87.4 86.7 0.3 20.2 6 90.1 59.8 13.2 89.9 89.4 3.8 29.6 87.4 86.4 0.7 26.6 5 90.2 37.7 14.1 89.8 78.0 15.8 40.3 87.2 79.8 0.8 42.1 4 87.5 14.7 6.0 89.8 27.8 18.9 13.1 87.2 34.4 20.5 19.7 3 78.3 10.5 1.5 89.7 10.9 2.6 0.4 86.8 11.0 2.9 0.5 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 H.2 GTSRB
-Base #Bits C.A.
-P.A.
-ConvL (1%) UNetL (1%) Model C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 8 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 7 95.5 69.5 10.6 95.7 95.3 0.3 25.8 96.2 95.9 0.3 26.4 6 95.4 67.2 14.4 95.7 95.2 0.5 28.0 96.2 95.7 0.5 28.5 5 95.4 48.6 18.2 95.8 92.6 5.1 44.0 96.2 94.8 2.5 46.2 4 92.6 24.6 9.8 95.9 75.6 16.2 51.0 96.2 86.6 9.5 62.0 3 67.7 5.3 3.5 95.4 18.4 15.3 13.1 96.2 25.3 22.5 20.0 2 3.8 3.8 0.0 4.1 3.8 0.0 0.0 3.8 3.8 0.0 0.0 VGG19 8 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 7 95.6 66.1 14.8 96.0 92.2 5.7 26.1 95.4 92.6 3.7 26.5 6 95.3 64.2 8.4 96.0 92.2 5.7 28.0 95.4 92.3 2.3 28.1 5 95.2 48.2 14.0 96.0 92.2 5.7 44.0 95.4 86.2 8.4 38.0 4 92.0 18.2 14.3 93.0 92.2 5.7 74.0 95.0 49.6 22.8 31.4 3 60.0 2.0 0.9 87.3 92.2 5.7 90.2 87.2 1.7 0.9 -0.3 2 5.9 3.8 0.0 5.9 3.8 0.0 0.0 5.9 3.8 0.0 0.0 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+Base #Bits CA PA ConvL (1%) UNetL (1%) Model CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 8 95.5 75.2 12.7 95.7 95.3 0.5 20.1 96.2 95.7 0.3 20.5 7 95.5 69.5 10.6 95.7 95.3 0.3 25.8 96.2 95.9 0.3 26.4 6 95.4 67.2 14.4 95.7 95.2 0.5 28.0 96.2 95.7 0.5 28.5 5 95.4 48.6 18.2 95.8 92.6 5.1 44.0 96.2 94.8 2.5 46.2 4 92.6 24.6 9.8 95.9 75.6 16.2 51.0 96.2 86.6 9.5 62.0 3 67.7 5.3 3.5 95.4 18.4 15.3 13.1 96.2 25.3 22.5 20.0 2 3.8 3.8 0.0 4.1 3.8 0.0 0.0 3.8 3.8 0.0 0.0 VGG19 8 95.5 69.1 11.1 96.0 94.0 2.2 24.9 95.4 93.9 2.1 24.8 7 95.6 66.1 14.8 96.0 92.2 5.7 26.1 95.4 92.6 3.7 26.5 6 95.3 64.2 8.4 96.0 92.2 5.7 28.0 95.4 92.3 2.3 28.1 5 95.2 48.2 14.0 96.0 92.2 5.7 44.0 95.4 86.2 8.4 38.0 4 92.0 18.2 14.3 93.0 92.2 5.7 74.0 95.0 49.6 22.8 31.4 3 60.0 2.0 0.9 87.3 92.2 5.7 90.2 87.2 1.7 0.9 -0.3 2 5.9 3.8 0.0 5.9 3.8 0.0 0.0 5.9 3.8 0.0 0.0 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 H.3 ImageNet-10
-Base #Bits C.A.
-P.A.
-ConvL (0.5%) UNetL (0.5%) Model C.A.
-(NF) P.A.
-(NF) R.P.
-C.A.
-(NF) P.A.
-(NF) R.P.
-ResNet18 8 92.2 72.3 7.0 94.0 88.0 2.0 15.7 94.0 88.1 1.4 15.8 7 92.4 70.6 13.0 94.2 86.7 4.1 16.1 93.6 87.8 3.5 17.2 6 92.4 68.9 9.9 94.2 85.1 4.8 16.2 93.6 86.4 3.7 17.5 5 91.0 60.9 13.0 94.2 82.5 6.8 21.6 94.0 83.2 5.9 22.3 4 91.4 47.4 9.8 93.8 68.6 9.8 21.2 92.6 68.7 9.2 21.3 3 85.2 28.8 11.8 89.2 44.1 14.0 15.3 89.4 42.7 14.2 13.9 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 92.4 37.2 11.0 91.4 75.5 8.8 38.3 89.4 77.9 6.1 40.7 7 92.0 27.3 6.6 91.2 59.3 13.0 32.0 89.4 65.4 10.0 38.1 6 92.4 27.9 6.4 91.0 59.7 11.8 31.8 89.4 64.9 9.9 37.0 5 92.0 15.1 4.4 91.6 23.1 0.7 8.0 89.0 27.9 8.8 12.8 4 89.4 12.2 2.7 90.8 14.0 4.3 1.8 89.6 14.6 4.9 2.4 3 46.8 9.9 0.5 83.2 10.4 0.6 0.5 84.2 9.9 0.7 0.0 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+Base #Bits CA PA ConvL (0.5%) UNetL (0.5%) Model CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 8 92.2 72.3 7.0 94.0 88.0 2.0 15.7 94.0 88.1 1.4 15.8 7 92.4 70.6 13.0 94.2 86.7 4.1 16.1 93.6 87.8 3.5 17.2 6 92.4 68.9 9.9 94.2 85.1 4.8 16.2 93.6 86.4 3.7 17.5 5 91.0 60.9 13.0 94.2 82.5 6.8 21.6 94.0 83.2 5.9 22.3 4 91.4 47.4 9.8 93.8 68.6 9.8 21.2 92.6 68.7 9.2 21.3 3 85.2 28.8 11.8 89.2 44.1 14.0 15.3 89.4 42.7 14.2 13.9 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 VGG19 8 92.4 37.2 11.0 91.4 75.5 8.8 38.3 89.4 77.9 6.1 40.7 7 92.0 27.3 6.6 91.2 59.3 13.0 32.0 89.4 65.4 10.0 38.1 6 92.4 27.9 6.4 91.0 59.7 11.8 31.8 89.4 64.9 9.9 37.0 5 92.0 15.1 4.4 91.6 23.1 0.7 8.0 89.0 27.9 8.8 12.8 4 89.4 12.2 2.7 90.8 14.0 4.3 1.8 89.6 14.6 4.9 2.4 3 46.8 9.9 0.5 83.2 10.4 0.6 0.5 84.2 9.9 0.7 0.0 2 10.0 10.0 0.0 10.0 10.0 0.0 0.0 10.0 10.0 0.0 0.0 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 Appendix I Additional Experiments on Adversarial Training
@@
 Empirically, we argue that adversarial training may not be suitable for training generator-based methods.
-K-bits C.A.
-P.A.
-C.A.
-(NF) P.A.
-(NF) R.P.
-100 92.6 38.9 12.4 92.4 38.3 12.1 -0.6 500 92.1 38.7 12.5 -0.2 5,000 92.6 38.9 12.5 0 20,000 60.1 23.0 8.1 -16 100,000 71.1 23.6 6.6 -16 • [Note] C.A.
-(%): clean accuracy, P.A.
-(%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A.
-(NF) v.s.
-P.A.
+K-bits CA PA CA (NF) PA (NF) RP 100 92.6 38.9 12.4 92.4 38.3 12.1 -0.6 500 92.1 38.7 12.5 -0.2 5,000 92.6 38.9 12.5 0 20,000 60.1 23.0 8.1 -16 100,000 71.1 23.6 6.6 -16 • Note.
+CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs.
+PA
 Appendix J Additional Experiments on Robust Model Trained with Adversarial Weight Perturbation with NeuralFuse
@@
 The experimental results are shown in Table 26.
-We find that the average perturbed accuracy is 23% and 63.2% for PreAct ResNet18 under 1% and 0.5% B.E.R., respectively.
+We find that the average perturbed accuracy is 23% and 63.2% for PreAct ResNet18 under 1% and 0.5% BER, respectively.
 This result is lower than 38.9% and 70.1% from ResNet18 in Table 12, indicating their poor generalization ability against random bit errors.
-Nevertheless, when equipped NeuralFuse on the perturbed model, we could still witness a significant recover percentage under both 1% and 0.5% B.E.R.
-This result further demonstrates that NeuralFuse could be adapted to various models (i.e., trained in different learning algorithms).
-| Base Model | B.E.R. | NF | C.A. | P.A. | C.A. (NF) | P.A. (NF) | R.P. |
+Nevertheless, when equipped NeuralFuse on the perturbed model, we could still witness a significant recover percentage under both 1% and 0.5% BER This result further demonstrates that NeuralFuse could be adapted to various models (i.e., trained in different learning algorithms).
+| Base Model | BER | NF | CA | PA | CA (NF) | PA (NF) | RP |
 | PreAct ResNet18 | 1% | ConvL | 89.7 | 23.0 9.3 | 87.6 | 53.7 26 | 30.7 |
@@
 | UNetS | 88.2 | 80.4 3.9 | 17.2 |
-[Note] B.E.R.: the bit error rate of the base model, NF: NeuralFuse, C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
+Note. BER: the bit-error rate of the base model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
 Appendix K Data Embeddings Visualization
-To further understand how our proposed NeuralFuse works, we visualize the output distribution from the final linear layer of the base models and project the results onto the 2D space using t-SNE (van der Maaten & Hinton, 2008).
+To further understand how our proposed NeuralFuse works, we visualize the output distribution from the final linear layer of the base models and project the results onto the 2D space using t-SNE (van der Maaten and Hinton, 2008).
 Figure 15 shows the output distribution from ResNet18 (trained on CIFAR-10) under a 1% bit error rate.
@@
 UNetS
-Clean |
-ConvL |
-ConvS |
-DeConvL |
-DeConvS |
-UNetL |
+| Clean | ||
+| ConvL | ||
+| ConvS | ||
+| DeConvL | ||
+| DeConvS | ||
+| UNetL | ||
 | UNetS |
-Clean |
-ConvL |
-ConvS |
-DeConvL |
-DeConvS |
-UNetL |
+| Clean | |
+| ConvL | |
+| ConvS | |
+| DeConvL | |
+| DeConvS | |
+| UNetL | |
 | UNetS |